### v1

# Equivariant Neural Networks for General Linear Symmetries on Lie Algebras

Abstract
Encoding symmetries is a powerful inductive bias for improving the generalization of deep neural networks. However, most existing equivariant models are limited to simple symmetries like rotations, failing to address the broader class of general linear transformations, $\mathrm{GL}(n)$, that appear in many scientific domains. We introduce Reductive Lie Neurons (ReLNs), a novel neural network architecture exactly equivariant to these general linear symmetries. ReLNs are designed to operate directly on a wide range of structured inputs, including general $n$-by-$n$ matrices. ReLNs introduce a novel adjoint-invariant bilinear layer to achieve stable equivariance for both Lie-algebraic features and matrix-valued inputs, without requiring redesign for each subgroup. This architecture overcomes the limitations of prior equivariant networks that only apply to compact groups or simple vector data. We validate ReLNs’ versatility across a spectrum of tasks: they outperform existing methods on algebraic benchmarks with $\mathfrak{sl}(3)$ and $\mathfrak{sp}(4)$ symmetries and achieve competitive results on a Lorentz-equivariant particle physics task. In 3D drone state estimation with geometric uncertainty, ReLNs jointly process velocities and covariances, yielding significant improvements in trajectory accuracy. ReLNs provide a practical and general framework for learning with broad linear group symmetries on Lie algebras and matrix-valued data. Project page: https://reductive-lie-neuron.github.io/
Chankyo Kim1,*Sicheng Zhao1,*Minghan Zhu1,2Tzu-Yuan Lin3Maani Ghaffari1
1University of Michigan
2University of Pennsylvania
3Massachusetts Institute of Technology
{chankyo,sichengz,minghanz,maanigj}@umich.edu, tzuyuan@mit.edu footnotetext: * Equal contribution.
1 Introduction
Leveraging the symmetries present in real-world data improves the generalization of neural networks. When a model is equivariant, i.e., its outputs transform predictably with inputs under certain group actions, it can learn more efficiently and generalize across scenarios with less data.
Most advances in equivariant deep learning target compact symmetry groups, such as rotations, $\mathrm{SO}(3)$, or isometries, $\mathrm{E}(n)$, where strong geometric and algebraic tools support robust network design (Bronstein et al., 2021; Cohen & Welling, 2016; Thomas et al., 2018; Satorras et al., 2021; Finzi et al., 2020; Geiger & Smidt, 2022). These methods have led to practical gains. However, many real-world phenomena, such as affine distortions, sensor anisotropy, and general changes of basis, are governed by the broader general linear group $\mathrm{GL}(n)=\{A\in\mathbb{R}^{n\times n}:\det(A)\neq 0\}$, for which existing methods fail to provide exact equivariance. These more general symmetries are crucial in a wide array of scientific domains, including robotics, particle physics, and computer vision, a landscape we survey in Table 1.
Naive solutions, such as flattening matrices into vectors, discard essential geometric structure, while methods based on spectral decompositions are brittle and numerically unstable  (Magnus, 1985). This leaves a gap for a numerically stable neural architecture capable of processing diverse geometric data types equivariantly under arbitrary invertible linear transformations (Bronstein et al., 2021).
Figure 1: Examples of Lie groups and related manifolds in scientific applications. From left: the special linear group $\mathrm{SL}(3)$ (image homography), the Lorentz group $\mathrm{SO}(1,3)$ (spacetime symmetry), symplectic groups $\mathrm{Sp}(n)$ (Hamiltonian mechanics), the $\mathrm{SPD}(3)\oplus\mathbb{R}^{3}$ state space (probabilistic estimation), and the general linear group $\mathrm{GL}(3)$ (modeling stress-strain in continuum mechanics). See Table 1 for a more detailed survey.
Table 1: A survey of common Lie groups and their applications in equivariant deep learning.
Group / Data Structure Application Reference
$\mathrm{SO}(3)$ 3D Point Clouds, State Estimation Lin et al. (2024a); Deng et al. (2021); Son et al. (2024)
$\mathrm{SO}^{+}(1,3)$ Particle Physics, Jet Tagging Bogatskiy et al. (2020); Finzi et al. (2021); Batatia et al. (2023)
$\mathrm{SU}(3)$ Quantum Chromodynamics (QCD) Favoni et al. (2022)
$\mathrm{SL}(3)$ Homography Classification, 3D Vision Lin et al. (2024a); Finzi et al. (2021)
$\mathrm{Sp}(4,\mathbb{R})$ Hamiltonian dynamics Lin et al. (2024a); Finzi et al. (2021)
$\mathrm{SPD}(n)$ Geometric Uncertainty Processing -
$\mathrm{GL}(n)$ General Linear Transformations Basu et al. (2025); Finzi et al. (2021)
Not a group. $\mathrm{SPD}(n)$ is the manifold of symmetric positive-definite matrices, representable as the quotient space $\mathrm{GL}(n)/\mathrm{O}(n)$.
This work introduces Reductive Lie Neurons (ReLNs), a unified neural architecture that provides exact equivariance to the adjoint action of the general linear group, $\mathrm{GL}(n)$, and its reductive subgroups, enabling robust learning on both Lie-algebraic features and structured geometric data like covariances. We resolve the central technical obstacle for non-semisimple algebras—the degeneracy of the canonical Killing form—by introducing a learnable, provably non-degenerate, and $\mathrm{Ad}$-invariant bilinear form. This provides a mathematically stable mechanism for defining symmetry-preserving nonlinearities, pooling, and invariant layers applicable to the reductive Lie algebra $\mathfrak{gl}(n)=\mathbb{R}^{n\times n}$.
Our main contributions are:
  1. 1.
    We propose Reductive Lie Neurons (ReLNs), a novel, general-purpose, and numerically stable network architecture for exact $\mathrm{GL}(n)$ adjoint equivariance.
  2. 2.
    We establish a connection between classical left-action equivariance and our adjoint-action framework for orthogonal groups. Through a provably-equivariant embedding map, we show that problems defined on standard vector actions—such as Lorentz transformations in particle physics or 3D point cloud processing—can be solved within our unified architecture, obviating the need for specialized model designs.
  3. 3.
    We establish the framework for geometric uncertainty-aware equivariant learning, enabling models to treat matrix-valued data that transforms under congruence (e.g., covariance tensors) as geometric objects.
  4. 4.
    We demonstrate the effectiveness of ReLNs through extensive experiments, showing that they outperform prior methods on Lie-algebraic benchmarks and achieve significant improvements in accuracy and robustness on a challenging 3D drone state estimation task.
2 Related Work
Encoding symmetry into neural architectures is a powerful inductive bias that improves data efficiency and generalization (Bronstein et al., 2021). In geometric deep learning, the most mature area of research focuses on enforcing equivariance to Euclidean isometries, transformations such as rotations and rigid motions formalized by Lie groups like $\mathrm{SO}(n)$ and $\mathrm{SE}(n)$. For grid-like data, foundational works include Group-Equivariant CNNs (Cohen & Welling, 2016) and Steerable CNNs (Weiler et al., 2018; Weiler & Cesa, 2021). For unstructured data, a dominant paradigm uses features associated with irreducible group representations (tensorial methods, E(n)-GNNs, and transformers (Thomas et al., 2018; Fuchs et al., 2020; Satorras et al., 2021; Batatia et al., 2022; Battiloro et al., 2025; Liao & Smidt, 2023; Assaad et al., 2023; Hutchinson et al., 2021)), with lightweight vector-based alternatives also available (Deng et al., 2021; Son et al., 2024). Theoretical analyses have also explored the universal approximation capabilities of such invariant networks (Maron et al., 2019). Complementing these specialized layers are general, model-agnostic strategies like frame averaging and canonicalization (Puny et al., 2022; Lin et al., 2024b; Kaba et al., 2023; Panigrahi & Mondal, 2024). Despite their success, most practical implementations are centered on scalar and vector features. While frameworks such as Tensor Field Networks (Thomas et al., 2018) or EMLP (Finzi et al., 2021) can handle higher-order tensor representations, scalable architectural design for general matrix-valued quantities (e.g., covariances or inertia tensors transforming as $\Sigma\mapsto R\Sigma R^{\top}$) is still limited in practice.
A major frontier is extending equivariance to non-compact groups like the general linear group $\mathrm{GL}(n)$ and the Lorentz group $\mathrm{SO}^{+}(1,3)$. Much of the work on the Lorentz group, which is critical in particle physics, has focused on designing specialized networks for the standard left action on vectors (Bogatskiy et al., 2020; Finzi et al., 2021; Batatia et al., 2023; Zhdanov et al., 2024). In contrast, we demonstrate that this left-action problem can be addressed within our universal, adjoint-equivariant framework via an embedding map.
One line of work generalizes group convolution and kernel design, often leveraging Fourier analysis (Xu et al., 2022; Helwig et al., 2023). This includes general frameworks for constructing equivariant networks on arbitrary matrix groups (Basu et al., 2025), through matrix functions (Batatia et al., 2024), and on reductive Lie groups (Batatia et al., 2023). Other approaches define kernels in the Lie algebra (Finzi et al., 2020), use Lie group decompositions for integration (Mironenco & Forré, 2024), or adapt canonicalization using infinitesimal generators of the Lie algebra (Shumaylov et al., 2025).
For these non-compact groups in general, the theoretical machinery underpinning many equivariant models does not readily apply. One line of work generalizes group convolution and kernel design, leveraging tools from Fourier analysis (Xu et al., 2022; Helwig et al., 2023), matrix functions (Batatia et al., 2024), or operating directly on reductive Lie groups (Batatia et al., 2023). Other methods focus on the Lie algebra, either by defining kernels within the algebra itself (Finzi et al., 2020) or using Lie group decompositions for integration (Mironenco & Forré, 2024; Shumaylov et al., 2025). Another approach leverages differential geometry, using tools such as partial differential operators (He et al., 2022; Shen et al., 2020; Jenner & Weiler, 2022) and the algebra of differential invariants (Sangalli et al., 2022; Li et al., 2024) to construct equivariant layers. A different generalist approach, taken by methods like the Equivariant MLP and G-RepsNet (Finzi et al., 2021; Basu et al., 2025), is to solve the equivariance constraint algebraically, though this often lacks specialized inductive biases like locality.
In contrast to these manifold-focused approaches, a parallel line of work operates directly on the Lie algebra. Lin et al. (2024a) introduced Lie Neurons, establishing a framework for adjoint-equivariance, but their method is restricted to semisimple Lie algebras where the Killing form is non-degenerate. This limitation precludes direct application to the reductive but non-semisimple algebra $\mathfrak{gl}(n)$, whose degenerate Killing form poses a central challenge. The high computational cost of group convolution methods, combined with the semisimple constraints of existing Lie-algebraic techniques, limits their direct applicability to robotics tasks that involve real-time processing, noisy measurements, or uncertainty estimation (Eschmann et al., 2024; Yu & Sun, 2024). While traditional, model-based algorithms like Kalman filters are designed to respect this geometry (Barrau & Bonnabel, 2016; Hartley et al., 2020), they lack the flexibility.
Our work confronts this challenge with ReLNs, a practical, numerically well-conditioned architecture for the full reductive group $\mathrm{GL}(n)$ and its subgroups. By introducing a learnable, non-degenerate, and $\mathrm{Ad}$-invariant bilinear form, our framework overcomes the key obstacle of the degenerate Killing form in $\mathfrak{gl}(n)$. This unified algebraic approach extends beyond strict Lie algebra elements to other geometric features, such as covariance matrices, that transform under congruence, providing a practical tool that sidesteps the complexities of prior methods requiring group integration or degenerate invariants. Figure 2 situates ReLNs as a general framework for $\mathrm{GL}(n)$ transformations, distinguishing our approach from architectures tailored to specialized subgroups.
3 Preliminaries
Our work builds equivariant networks on the Lie algebra $\mathfrak{g}=\mathrm{Lie}(G)$, the tangent space of a Lie group $G$ at the identity $e$. We focus on the general linear group $\mathrm{GL}(n)$ and its Lie algebra $\mathfrak{gl}(n)$, aiming for equivariance under the adjoint action $\mathrm{Ad}_{g}:X\mapsto gXg^{-1},\;g\in\mathrm{GL}(n),\;X\in\mathfrak{gl}(n)$. A central challenge arises from the structure of $\mathfrak{gl}(n)$: it is a reductive Lie algebra, and the canonical invariant inner product, the Killing form, is degenerate on $\mathfrak{gl}(n)$. This degeneracy poses a critical problem, as it prevents the construction of expressive nonlinear layers, causing them to collapse into linear maps and severely limiting the model’s expressive power. We address this by introducing a learnable, non-degenerate bilinear form, enabling fully nonlinear equivariant operations. For details of Lie theory and background, see Appendix A.
Figure 2: A taxonomy of selected representative equivariant neural architectures, categorized by the symmetries to which they are equivariant. This diagram situates our work, ReLNs, among other notable methods that are often specialized for subgroups such as $\mathrm{SL}(n)$, $\mathrm{SO}(n)$, or the Euclidean group $\mathrm{E}(n)$. An asterisk () denotes methods equivariant to the group’s adjoint action.
4 Reductive Lie Neurons: Architecture
We present ReLNs, a framework for building deep networks equivariant to the adjoint action of the general linear group $\mathrm{GL}(n)$. The design centers on a learnable, non-degenerate, $\mathrm{Ad}$-invariant bilinear form on the reductive Lie algebra $\mathfrak{gl}(n)$ and a complete toolbox of equivariant linear maps, nonlinearities, pooling, and invariant readouts.
4.1 A General $\mathrm{Ad}$-invariant Bilinear Form for Reductive Lie Algebras
The primary obstacle to applying Lie-algebraic methods such as Lie Neurons (Lin et al., 2024a) to $\mathfrak{gl}(n)$ is the degeneracy of its Killing form. We resolve this by constructing a modified bilinear form $\widetilde{B}$ that restores non-degeneracy while preserving the crucial $\mathrm{Ad}$-invariance property.
Definition 4.1 (Modified Bilinear Form on a Reductive Lie Algebra).
If $\mathfrak{g}$ is reductive, then $\mathfrak{g}=\mathfrak{z}(\mathfrak{g})\oplus[\mathfrak{g},\mathfrak{g}]$, where $\mathfrak{z}(\mathfrak{g})$ is the center. Choose any $\mathrm{Ad}$-invariant inner product $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ on $\mathfrak{z}(\mathfrak{g})$ (for connected $G$ this is automatic since $\mathrm{Ad}|_{\mathfrak{z}(\mathfrak{g})}:G\to\mathrm{GL}(\mathfrak{z}(\mathfrak{g}))$ is locally constant. See Appendix A.2 for the formal definition of $\mathrm{Ad}$) . For $Z_{i}\in\mathfrak{z}(\mathfrak{g})$ and $X_{i}\in[\mathfrak{g},\mathfrak{g}]$ define
$\widetilde{B}(Z_{1}{+}X_{1},\,Z_{2}{+}X_{2}):=\langle Z_{1},Z_{2}\rangle_{\mathfrak{z}}+B(X_{1},X_{2}),\qquad$ (1)
where $B$ denotes the Killing form on $\mathfrak{g},\mathfrak{g}]$.
Proposition 4.1.
The bilinear form $\widetilde{B}$ is symmetric, $\mathrm{Ad}$-invariant, and nondegenerate. Moreover, $\mathfrak{z}(\mathfrak{g})$ and $\mathfrak{g},\mathfrak{g}]$ are $\widetilde{B}-$orthogonal, with $\widetilde{B}|_{[\mathfrak{g},\mathfrak{g}]}=B$ and $\widetilde{B}|_{\mathfrak{z}(\mathfrak{g})}=\langle\cdot,\cdot\rangle_{\mathfrak{z}}$.
Proof sketch.
$B$ vanishes on $\mathfrak{z}(\mathfrak{g})$ and is $\mathrm{Ad}$-invariant; by construction $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ is $\mathrm{Ad}$-invariant. Symmetry is immediate. Nondegeneracy follows since $B$ is nondegenerate on the semisimple ideal and $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ is nondegenerate on the center; orthogonality holds because $B(\mathfrak{z}(\mathfrak{g}),[\mathfrak{g},\mathfrak{g}])=0$. ∎
For our primary case $\mathfrak{g}=\mathfrak{gl}(n,\mathbb{R})=\mathbb{R}I\oplus\mathfrak{sl}(n,\mathbb{R})$ we choose the canonical trace-based form
$\widetilde{B}(X,Y)\;=\;2n\cdot\mathrm{tr}(XY)\;-\;\mathrm{tr}(X)\,\mathrm{tr}(Y),$ (2)
This form $\widetilde{B}$ is the fundamental tool used throughout the ReLN architecture.
Verification and Relation to Prior Bilinear Forms.
Our concrete form in Eq. 2 satisfies the conditions of Proposition 4.1. Decomposing a matrix $X=X_{0}+\frac{1}{n}\mathrm{tr}(X)I$ (where $X_{0}\in\mathfrak{sl}(n)$) reveals that our form separates orthogonally:
$\widetilde{B}(X,Y)=\underbrace{2n\cdot\mathrm{tr}(X_{0}Y_{0})}_{B_{\mathfrak{sl}(n)}(X_{0},Y_{0})}+\underbrace{\mathrm{tr}(X)\mathrm{tr}(Y)}_{\text{Inner product on }\mathbb{R}I}.$ (3)
This decomposition directly shows how $\widetilde{B}$ serves as a generalization of prior work. The first term is the Killing form on the semisimple part, which is the tool used in Lie Neurons (Lin et al., 2024a). The second term is a standard inner product on the center, which, under the isomorphism $\mathfrak{so}(3)\simeq\mathbb{R}^{3}$, recovers the dot product used in Vector Neurons (Deng et al., 2021). Our single form thus unifies these approaches, extending to the full reductive algebra $\mathfrak{gl}(n)$ (details can be found in Appendix C).
Figure 3: Adjoint equivariance using a unified representation for diverse geometric inputs. Our framework embeds inputs with different transformation rules, such as velocity ($v\mapsto Rv$) and covariance ($\Sigma\mapsto R\Sigma R^{T}$), into a common Lie algebra. Therefore, they transform under the same adjoint action $\mathrm{Ad}_{g}$, with which our network $f$ commutes as shown in the diagram.
4.2 The ReLN Layer Toolbox
We represent multi-channel input as $x\in\mathbb{R}^{K\times C}$, where each column $x_{c}\in\mathbb{R}^{K}$ corresponds to a matrix $X_{c}\in\mathfrak{g}$ (via the vee/hat isomorphism, Appendix A) .
ReLN-Linear.
A linear map applied to the channel dimension $f(x;W)=xW$ with $W\in\mathbb{R}^{C\times C^{\prime}}$ is strictly equivariant: the group acts on the left (geometric dimension) while $W$ acts on the right (channel dimension), and thus these operations commute (formal proof in Appendix D).
Equivariant Nonlinearities.
We introduce two complementary nonlinear primitives; full definitions, parameterizations, and stability prescriptions are deferred to Appendix D.
  • ReLN-ReLU: To overcome the non-equivariance of elementwise activations, we use our form $\widetilde{B}$ to define a directional nonlinearity. Each vector feature $x_{c}$ is rectified along a learnable direction $d_{c}$ via update rule:
  • ReLN-ReLU Using $\widetilde{B}$ to build an invariant gate, each channel feature is updated by
    $x_{c}^{\prime}\;=\;x_{c}\;+\;\max\big(0,\;\widetilde{B}(x_{c}^{\wedge},d_{c}^{\wedge})\big)\,d_{c},$ (4)
[…]
  • ReLN-Bracket. Following prior work (Lin et al., 2024a), we include a layer that leverages the Lie bracket (matrix commutator). This operation is an $\mathrm{Ad}$-equivariant primitive on the Lie algebra that creates nonlinear interactions by measuring the non-commutativity of features. The layer applies two independent linear maps, parameterized by weights $W_{a},W_{b}\in\mathbb{R}^{C\times C}$, to the input channels $x_{\mathrm{in}}$ to produce two intermediate features, computes their commutator, and injects the vectorized result as a shared residual:
    $x_{\mathrm{out}}=x_{\mathrm{in}}+\big(\left[(x_{\mathrm{in}}W_{a})^{\wedge},(x_{\mathrm{in}}W_{b})^{\wedge}\right]\big)^{\vee}.$ (5)
Equivariant Pooling and Invariant Layers.
The final components of the ReLN toolbox enable feature aggregation and the production of invariant outputs.
  • Max-Killing Pooling: To aggregate a set of features $\{X_{n}\}_{n=1}^{N}$, where each $X_{n}$ is a multi-channel feature tensor, this layer selects the representative feature with the maximal projection onto a learnable direction. For each channel $c$, the index is found via $n^{*}(c)=\arg\max_{n}\widetilde{B}(X_{n,c},D_{n,c})$, and the pooled feature is $X^{\max}_{c}=X_{n^{*}(c),c}$.
  • Invariant Layer: To produce a group-invariant output, this layer contracts feature $X_{c}$ using the form $\widetilde{B}$. The resulting scalar, $y_{c}=\widetilde{B}(X_{c},X_{c})$, is invariant by construction.
Unifying geometric representations.
By operating directly on $n\times n$ matrix representations, ReLNs provide a unified primitive for vectors, matrices, and higher-order geometric objects (e.g., covariances). This allows ReLNs to handle a broader class of geometric inputs without resorting to separate specialized architectures; empirical validation is presented in Section 5.3.
5 Experiments
We evaluate ReLNs on a suite of tasks designed to highlight two complementary strengths: algebraic generality on benchmarks and practical efficacy on a challenging, uncertainty-aware robotics task. We compare against standard non-equivariant baselines (MLP, ResNet), the original Lie Neurons, and a Vector Neurons–style baseline adapted for covariance inputs by eigendecomposition. For the Top-Tagging benchmark we also report results versus established physics models used in prior work.
5.1 Algebraic Benchmarks on Semisimple Lie Algebras
To verify that our general $\mathfrak{gl}(n)$ framework correctly generalizes to semisimple subalgebras, we evaluate ReLN on two Lie-algebraic benchmarks first introduced by Lie Neurons (Lin et al., 2024a).
Table 2: Platonic solid classification (mean $\pm$ std over 5 runs). ID = in-distribution; RC = rotated-camera (10 random $\mathrm{SO}(3)$ test rotations). “+Aug” denotes training with random $\mathrm{SO}(3)$ augmentation of input homographies. Higher is better ($\uparrow$).
Model # Params ID Acc (mean $\pm$ std) RC Acc (mean $\pm$ std)
MLP 206,339 95.76% $\pm$ 0.65% 36.54% $\pm$ 0.99%
MLP + Aug 206,339 81.47% $\pm$ 0.77% 81.20% $\pm$ 2.34%
MLP (wider) 411,479 96.82% $\pm$ 0.53% 36.55% $\pm$ 0.34%
MLP (wider) + Aug 411,479 85.22% $\pm$ 1.46% 83.43% $\pm$ 0.51%
Lie Neurons 331,272 99.62% $\pm$ 0.25% 99.61% $\pm$ 0.14%
ReLN (Ours) 331,272 99.78% $\pm$ 0.04% 99.78% $\pm$ 0.04%
5.1.1 Platonic Solid Classification
We first validate our model on the Platonic solid classification benchmark from (Lin et al., 2024a), testing the adjoint-equivariance where camera rotations induce a conjugation action on inter-face homographies in $\mathrm{SL}(3)$. Full experimental details are provided in Appendix E.2.
The results, summarized in Table 2, confirm that non-equivariant baselines fail to generalize to rotated camera views. This fundamental weakness is not resolved by data augmentation or increased model capacity, as our wider MLP variant with approximately double the parameters shows negligible improvement on the out-of-distribution test set. In contrast, the ReLN model achieves near-perfect accuracy with robustness, matching the performance of the Lie Neurons while demonstrating improved results. Importantly, this result validates that our general $\mathfrak{gl}(n)$ framework operates effectively on common semisimple subalgebras, as its built-in adjoint-equivariance on the parent group yields robust behavior when restricted to subgroups like $\mathrm{SO}(3)$ and $\mathrm{SL}(3)$.
5.1.2 Invariant Function Regression on $\mathfrak{sp}(4)$.
To further probe our framework’s algebraic generality, our second benchmark involves regressing a highly nonlinear invariant function on the real symplectic Lie algebra $\mathfrak{sp}(4,\mathbb{R})$. The symplectic algebra $\mathfrak{sp}(2n)$ is the mathematical foundation of Hamiltonian mechanics, which describes any physical system where energy is conserved. Our target is a scalar invariant defined for pairs $X,Y\in\mathfrak{sp}(4,\mathbb{R})$:
$g(X,Y)=\sin\!\big(\operatorname{Tr}(XY)\big)+\cos\!\big(\operatorname{Tr}(YY)\big)-\tfrac{1}{2}\operatorname{Tr}(YY)^{3}+\det(XY)+\exp\!\big(\operatorname{Tr}(XX)\big).$ (6)
We generate a dataset of 10k training and 10k test pairs by sampling from $\mathfrak{sp}(4,\mathbb{R})$. We compare ReLN against MLP baselines (trained with and without $\mathrm{Sp}(4)$ data augmentation) and the original Lie Neurons. At test time, we report the standard MSE, MSE averaged over 500 random adjoint actions, and the invariance error.
As shown in Table 3, non-equivariant MLPs are orders of magnitude less accurate and exhibit high invariance error, failing to learn the group structure even with data augmentation. Our ReLN model not only achieves the lowest MSE and near-zero invariance error, but also shows a modest but consistent improvement over Lie Neurons. This suggests that our non-degenerate bilinear form provides not only theoretical generality but also superior numerical conditioning in practice.
Table 3: Regression performance and invariance error on $\mathfrak{sp}(4)$. “Training Aug.” indicates whether $\mathrm{Sp}(4)$ conjugation was applied during training.
Model Training Aug. # Params Test Aug. Invariance Error
ID SP(4)
MLP 256 Id 137,217 0.126 1.360 0.722
SP(4) 137,217 0.192 0.587 0.476
MLP 512 Id 536,577 0.107 0.906 0.585
SP(4) 536,577 0.123 0.446 0.374
Lie Neurons Id 263,170 $5.83\times 10^{-4}$ $5.84\times 10^{-4}$ $3.84\times 10^{-7}$}
ReLN (ours) Id 263,170 $5.14\times 10^{-4}$} $5.14\times 10^{-4}$} $4.73\times 10^{-7}$
5.2 Particle Physics with Lorentz Group $\mathrm{SO}(1,3)$ Equivariance
We test our framework on the Top-Tagging benchmark (Kasieczka et al., 2019), a task to distinguish particle jets originating from top quarks against a large background from Quantum Chromodynamics. Because these relativistic collisions are subject to the symmetries of spacetime, the task requires equivariance under the Lorentz group $\mathrm{SO}(1,3)$. We solve this left-action equivariant learning problem by introducing a map that embeds the four-momentum $p\in\mathbb{R}^{1,3}$ of each constituent particle, which combines its energy and 3D momentum, into the Lie algebra $\mathfrak{gl}(5)$ as $\varphi(p)=\begin{pmatrix}0_{4\times 4}&p\\ p^{\top}\eta&0\end{pmatrix}$, where $\eta=\mathrm{diag}(-1,1,1,1)$ is the Minkowski metric. As proven in Appendix F.1, this embedding unifies left- and adjoint-action equivariance within a single Lie-algebraic framework. We adapt the LorentzNet architecture by replacing its invariant feature computation with our proposed bilinear form. To ensure a fair comparison, we created a parameter-matched version of LorentzNet. As shown in Table 4, our model achieves competitive overall performance while demonstrating an advantage on the background rejection metric. This result shows that our general Lie-algebraic approach can effectively unify adjoint- and left-action equivariance in a parameter-efficient manner.
Table 4: Comparison of performance on the Top-Tagging dataset. Rej@30% denotes the background rejection at 30% signal efficiency (higher is better). Benchmark scores are as reported in the original publications.
Architecture #Params Accuracy AUC Rej@30% Reference
PELICAN 45k 0.943 0.987 $2289\pm 204$ Bogatskiy et al. (2022)
LorentzNet (original) 224k 0.942 0.987 $2195\pm 173$ Gong et al. (2022)
LorentzNet (param-matched) 84k 0.942 0.987 $1821\pm 94$ Our reproduction
LGN 4.5k 0.929 0.964 $435\pm 95$ Bogatskiy et al. (2020)
BIP 4k 0.931 0.981 $853\pm 68$ Munoz et al. (2022)
partT 2.14M 0.940 0.986 $1602\pm 81$ Qu et al. (2022)
ParticleNet 498k 0.938 0.985 $1298\pm 46$ Qu & Gouskos (2020)
EFN 82k 0.927 0.979 $633\pm 31$ Komiske et al. (2019)
TopoDNN 59k 0.916 0.972 $295\pm 5$ Pearkes et al. (2017)
LorentzMACE 228k 0.942 0.987 $1935\pm 85$ Batatia et al. (2023)
ReLN (Ours) 84k 0.942 0.987 $1979\pm 87$
5.3 Drone State Estimation with Geometric Uncertainty
We test our framework on a challenging drone state estimation task using a large-scale dataset of aggressive, highly dynamic flights. The objective is to recover a 3D trajectory from a sequence of noisy velocity measurements and their corresponding time-varying covariances, where each covariance matrix quantifies the uncertainty of its associated velocity measurement. This setup tests a model’s ability to jointly process vector ($\mathbf{v}$) and matrix ($C$) data in a geometrically consistent and uncertainty-aware manner.
Experimental Setup.
We created a large-scale synthetic dataset of 200 aggressive drone trajectories, over 13 hours of challenging, high-speed flight (details in Appendix G). The network is trained to regress the 3D position from a sequence of velocity and covariance measurements within a time window. Our model is compared against non-equivariant ResNets and an $\mathrm{SO}(3)$-equivariant baseline using Vector Neurons (VN). Since VNs cannot directly process matrix inputs, we implement a method that handles covariance matrices via an eigendecomposition-based strategy (details in Appendix G.3).
In contrast, our ReLN architecture treats both velocity and covariance as unified geometric objects within a single algebraic space, $\mathfrak{gl}(3)$. Specifically, the velocity vector $\mathbf{v}\in\mathbb{R}^{3}$ is lifted to its matrix representation in $\mathfrak{so}(3)\subset\mathfrak{gl}(3)$. For the time-varying covariance $C$, we explore two representations that also reside in $\mathfrak{gl}(3)$: (1) the matrix $C$ directly, and (2) its matrix logarithm $\log C$, which respects the geometry of the $\mathrm{SPD}(3)$ manifold. This unified representation ensures a measurement and its uncertainty transform consistently under the adjoint action of $\mathrm{SO}(3)$. The final velocity estimate is then equivariantly extracted by projecting the network’s matrix output onto its skew-symmetric component. We test three ReLN variants: (1) velocity only, (2) velocity + covariance, and (3) velocity + log-covariance. Full implementation details are provided in Appendix G.
5.3.1 Results and Analysis
Table 5: Performance on the drone trajectory dataset. Best result in each column is shown in bold.
Model ID $\mathrm{SO}(3)$
ATE ATE% RPE ATE ATE% RPE
Non-Equivariant Baselines
   ResNet (Velocity only) $208.07$ $95.06$ $107.60$ $217.02$ $100.39$ $111.29$
   ResNet (Velocity + Covariance) $205.11$ $94.94$ $106.07$ $213.26$ $98.90$ $109.37$
Equivariant Baselines
   VN (Velocity only) $17.36$ $7.52$ $13.51$ $17.36$ $7.52$ $13.51$
   VN (Velocity + Covariance) $191.78$ $88.66$ $98.39$ $190.22$ $88.47$ $98.26$
Our Equivariant Models
   ReLN (Velocity only) $16.85$ $7.31$ $12.7$ $16.85$ $7.31$ $12.7$
   ReLN (Velocity + Covariance) $16.49$ $7.21$ $13.02$ $16.49$ $7.21$ $13.02$
   ReLN (Velocity + $\log$-Covariance) 13.92 5.99 11.04 13.92 5.99 11.04
Abbreviations and units: ID = in-distribution test (no rotation), $\mathrm{SO}(3)$ = test-time random $\mathrm{SO}(3)$ rotations. ATE = absolute trajectory error (meters), ATE% = ATE relative to trajectory length (%), RPE = relative pose error (meters).
Figure 4: Qualitative comparison of the best-case (middle), average (left), and challenging (right) test sequences. ReLN models consistently track the ground truth (black) with high fidelity, especially when leveraging covariance. Insets provide a magnified view of the two best-performing variants (ours) to highlight their accuracy.
Our experiments, summarized in Table 5 and Figure 4, reveal a clear performance hierarchy where the geometric representation of features is the critical factor.
Non-equivariant models fail to generalize. Non-equivariant models like ResNets fail to generalize to rotated trajectories, and their performance does not improve when provided with covariance data.
Equivariance is necessary, but implementation design critically affects performance. Both VNs and ReLNs establish strong velocity-only equivariant baselines. ReLNs show improved performance due to its richer feature learning from both the Lie bracket and our bilinear nonlinearity form. However, a method for incorporating covariance is crucial. The VN baseline, which decomposes covariance, degrades the performance. This result confirms that separating the principle axes and its corresponding scales (eigenvalues) prevents the network from learning the structure between a measurement and its uncertainty.
Joint processing of velocity and covariance further improves performance. Integrating covariance as a geometric object in our framework further reduces the ATE, showcasing the performance gains from unified geometric representation on symmetry-preserving uncertainty handling.
Log-covariance ablation highlights mathematical and practical contributions. Finally, the ReLN (Velocity + $\log$-Covariance) model achieves the best performance. By processing covariance matrices via the matrix logarithm, we demonstrate the power of extending Lie-algebraic architectures to operate on manifold-valued data, such as the $\mathrm{SPD}(3)$ matrices representing geometric uncertainty. This success establishes ReLNs as a practical and high-performance framework for uncertainty-aware learning in dynamical systems.
Discussion Our results indicate that ReLNs operate effectively as a geometry- and uncertainty-aware estimator that generalizes across random measurement-frame changes (i.e., arbitrary 3D rotations). While the network is not a classical recursive Markovian filter, it learns to integrate velocity measurements with their associated covariances using uncertainty-dependent weighting, producing robust and accurate trajectory estimates. This behavior suggests ReLNs are suitable as modular components in downstream systems that require handling of matrix-valued uncertainty. Future work will investigate theoretical guarantees for the learned weighting, extensions to other geometric matrix representations, and scalability to higher-dimensional structured geometric inputs.
6 Conclusion
This work introduces ReLNs, a unified neural architecture that provides exact equivariance to the adjoint action of the general $n\times n$ matrix algebra $\mathfrak{gl}(n)$ and its subgroups. ReLNs enable efficient learning on Lie-algebraic features and structured geometric data, such as covariance matrices. Furthermore, our work establishes a unified Lie-algebraic framework that handles both classical left-action symmetries on vectors and native adjoint-actions on matrices within a single architecture. ReLNs achieve state-of-the-art results on benchmarks and deliver large gains in a challenging drone state estimation task by integrating uncertainty. We’ll apply our equivariant matrix processing capability to a wider array of physical systems, including the dynamics of articulated robots and large-scale sensor fusion, to further expand the boundaries of geometric deep learning.
Acknowledgments
This work was supported by AFOSR MURI FA9550-23-1-0400 and AFOSR YIP FA9550-25-1-0224. The authors thank Dr. Frederick A. Leve for his encouragement and support.
Appendix A Lie-theoretic preliminaries
This appendix provides an overview of key concepts and derivations from Lie group theory relevant to our construction of $\mathrm{GL}(n)$-equivariant neural networks referenced in the main text.
A.1 Lie groups, Lie algebras, hat/vee
A matrix Lie group $G\subset\mathrm{GL}(n)$ is a smooth subgroup of invertible matrices. Its Lie algebra $\mathfrak{g}=\operatorname{Lie}(G)$ is the tangent space at the identity and is identified with a subspace of $\mathfrak{gl}(n)$. Fix a basis $\{E_{i}\}_{i=1}^{m}$ of $\mathfrak{g}$. The coordinate maps are:
$\wedge:\mathbb{R}^{m}\to\mathfrak{g},\quad x=(x_{i})\mapsto x^{\wedge}=\sum_{i}x_{i}E_{i},\qquad\vee:\mathfrak{g}\to\mathbb{R}^{m},\quad X\mapsto X^{\vee}.$ (7)
These maps let us implement algebra-valued features as Euclidean vectors in code.
The associated Lie algebra $\mathfrak{g}=\mathrm{Lie}(G)$ is the tangent space at the identity element $e\in G$. It carries a bilinear, antisymmetric product called the Lie bracket, given by
[A,B]=ABBA,[A,B]=AB-BA, (8)
in the case of $\mathrm{GL}(n)$ which captures the infinitesimal structure of the group near the identity. The bracket quantifies non-commutativity of generators: $[A,B]=0$ implies commutativity, whereas $A,B]\neq 0$ indicates a non-trivial interaction.
A.2 Representations and the adjoint
A representation $\Phi:G\to\mathrm{GL}(V)$ differentiates to $\phi:\mathfrak{g}\to\mathfrak{gl}(V)$ by
$\phi(X)=\left.\frac{d}{dt}\right|_{t=0}\Phi(\exp(tX)).$ (9)
The adjoint representation $\mathrm{Ad}:G\to\mathrm{GL}(\mathfrak{g})$ is defined to be the differential of group conjugation at the identity
$\mathrm{Ad}_{g}(X)=\left.\frac{d}{dt}\right|_{t=0}g(\exp(tX))g^{-1}.$ (10)
Therefore we get a map $\mathrm{Ad}:G\to\mathrm{GL}(\mathfrak{g})$. For matrix groups, this is given
$\mathrm{Ad}_{g}(X)=gXg^{-1},\qquad g\in G,\;X\in\mathfrak{g},$ (11)
and differentiating yields the Lie-algebra adjoint $\mathrm{ad}_{X}(Y)=[X,Y]$. One checks
$\mathrm{Ad}_{g}([X,Y])=[\mathrm{Ad}_{g}X,\mathrm{Ad}_{g}Y],\qquad\mathrm{ad}_{X}([Y,Z])=[\mathrm{ad}_{X}Y,Z]+[Y,\mathrm{ad}_{X}Z].$ (12)
A.3 Vectorized adjoint
Using the hat/vee maps, the adjoint action on the Lie algebra induces a corresponding action on the vector coordinates. This vectorized action is a linear map represented by a matrix:
$\mathrm{Ad}^{m}_{g}:\mathbb{R}^{m}\to\mathbb{R}^{m},\qquad\mathrm{Ad}^{m}_{g}(x)=(\mathrm{Ad}_{g}(x^{\wedge}))^{\vee}.$ (13)
In practice we precompute or assemble the $m\times m$ matrix representing $\mathrm{Ad}^{m}_{g}$ (or apply it implicitly) to implement left-multiplicative equivariant layers that act on vector features.
A.4 Structure of Lie Algebra: Semisimplicity and Reductivity
Definition A.1 (Semisimple and Reductive Lie Algebras).
A Lie algebra $\mathfrak{g}$ is:
  • Semisimple if it is a direct sum of simple Lie algebras (i.e., non-abelian and having no nontrivial ideals).
  • Reductive if it decomposes as $\mathfrak{g}=\mathfrak{s}\oplus\mathfrak{z}$, where $\mathfrak{s}$ is semisimple and $\mathfrak{z}$ is the center (an abelian Lie subalgebra).
Example A.1.
The Lie algebra $\mathfrak{gl}(n)$ decomposes as:
$\mathfrak{gl}(n)=\mathfrak{sl}(n)\oplus\mathbb{R}I,$ (14)
where $\mathfrak{sl}(n)$ (traceless matrices) is semisimple, and $\mathbb{R}I$ (scalar matrices) forms the center.
This decomposition highlights the non-semisimple nature of $\mathfrak{gl}(n)$, which plays a critical role in understanding the degeneracy of certain invariant forms such as the Killing form. This degeneracy hinders the application of standard tools in Lie-theoretic deep learning. Our work addresses this issue in the context of $\mathrm{GL}(n)$-equivariant architectures in lie algebra $\mathfrak{gl}(n)$.
Theorem A.1 (Cartan criterion; standard).
The Killing form $B(X,Y)=\operatorname{tr}(\mathrm{ad}_{X}\circ\mathrm{ad}_{Y})$ is non-degenerate iff $\mathfrak{g}$ is semisimple.
In particular, because $\mathfrak{gl}(n)$ contains the central scalar direction $\mathbb{R}I$ with $\mathrm{ad}_{I}=0$, the Killing form is degenerate on $\mathfrak{gl}(n)$.
A.5 Invariant bilinear forms; trace form
Definition A.2.
A bilinear form $B:\mathfrak{g}\times\mathfrak{g}\to\mathbb{R}$ is $\mathrm{Ad}$-invariant if
$B(\mathrm{Ad}_{g}X,\mathrm{Ad}_{g}Y)=B(X,Y)\quad\forall g\in G.$ (15)
On semisimple algebras the Killing form provides such an invariant, non-degenerate form. On $\mathfrak{gl}(n)$ we instead use the trace form:
$\langle X,Y\rangle_{\mathrm{tr}}=\operatorname{tr}(XY).$ (16)
Proposition A.1.
The trace form is $\mathrm{Ad}$-invariant on $\mathfrak{gl}(n)$:
$\operatorname{tr}((gXg^{-1})(gYg^{-1}))=\operatorname{tr}(gXYg^{-1})=\operatorname{tr}(XY).$ (17)
Proof.
$\operatorname{tr}((gXg^{-1})(gYg^{-1}))=\operatorname{tr}(gXYg^{-1})=\operatorname{tr}(XY)$ by cyclicity of trace. ∎
The trace form is non-degenerate as a bilinear form on the vector space $\mathfrak{gl}(n)$ and therefore provides a practical substitute for the Killing form when designing $\mathrm{Ad}$-invariant bilinear layers on $\mathfrak{gl}(n)$.
A.6 Killing form on $\mathfrak{sl}(n)$
Restricted to $\mathfrak{sl}(n)$ the Killing form simplifies and is non-degenerate; one frequently uses the proportionality $B(X,Y)\propto\operatorname{tr}(XY)$ on $\mathfrak{sl}(n)$.
Appendix B Proofs of Key Theorems
B.1 Proof of Non-degeneracy and Ad-invariance of Modified Killing Form $B_{e}$
Let $\mathfrak{g}$ be a real reductive Lie algebra.
Definition B.1 (Reductive decomposition).
A Lie algebra $\mathfrak{g}$ is reductive if $\mathfrak{g}=\mathfrak{z}(\mathfrak{g})\oplus[\mathfrak{g},\mathfrak{g}]$, where $\mathfrak{z}(\mathfrak{g})$ is the center and $\mathfrak{g},\mathfrak{g}]$ is semisimple. This decomposition is canonical (both summands are ideals).
Definition B.2 (Modified Killing form on a reductive Lie algebra).
Fix any symmetric, positive–definite inner product $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ on $\mathfrak{z}(\mathfrak{g})$, and let $B$ denote the Killing form on the semisimple ideal $\mathfrak{g},\mathfrak{g}]$. For $Z_{i}\in\mathfrak{z}(\mathfrak{g})$ and $X_{i}\in[\mathfrak{g},\mathfrak{g}]$ define
$\widetilde{B}(Z_{1}{+}X_{1},\,Z_{2}{+}X_{2})\;:=\;\langle Z_{1},Z_{2}\rangle_{\mathfrak{z}}\;+\;B(X_{1},X_{2}).$ (18)
Remark 1 (Canonicity).
On $\mathfrak{g},\mathfrak{g}]$ the restriction (Killing form) is canonical. On $\mathfrak{z}(\mathfrak{g})$ there is no canonical choice; any $\mathrm{Ad}$-invariant positive-definite inner product works. The choice we make in the case of $\mathfrak{gl}(n)$ ensures that it agrees with the Killing form on the semisimple part $\mathfrak{sl}(n)$, and the center $\mathbb{R}I$ is normalized by a natural trace scale.
Proposition B.1 (Block–orthogonality and restrictions).
With notation as above,
$\widetilde{B}\big(\mathfrak{z}(\mathfrak{g}),[\mathfrak{g},\mathfrak{g}]\big)=0,\qquad\widetilde{B}|_{\mathfrak{z}(\mathfrak{g})}=\langle\cdot,\cdot\rangle_{\mathfrak{z}},\qquad\widetilde{B}|_{[\mathfrak{g},\mathfrak{g}]}=B.$ (19)
Proposition B.2 (Non–degeneracy).
$\widetilde{B}$ is nondegenerate on $\mathfrak{g}$.
Proof.
Let $X=Z+W$ with $Z\in\mathfrak{z}(\mathfrak{g})$ and $W\in[\mathfrak{g},\mathfrak{g}]$. If $\widetilde{B}(X,\cdot)\equiv 0$, then testing against $Y\in\mathfrak{z}(\mathfrak{g})$ yields $\langle Z,Y\rangle_{\mathfrak{z}}=0$ for all $Y$, hence $Z=0$; testing against $Y\in[\mathfrak{g},\mathfrak{g}]$ yields $B(W,Y)=0$ for all $Y$, hence $W=0$ by the non–degeneracy of $B$ on the semisimple ideal. Thus $X=0$. ∎
Proposition B.3 ($\mathrm{Ad}$–invariance on the identity component).
$\widetilde{B}$ is $\mathrm{ad}$–invariant:
$\widetilde{B}([X,Y],Z)+\widetilde{B}(Y,[X,Z])=0\qquad\text{for all }X,Y,Z\in\mathfrak{g},$ (20)
and hence $\widetilde{B}(\mathrm{Ad}_{g}Y,\mathrm{Ad}_{g}Z)=\widetilde{B}(Y,Z)$ for all $g$ in the identity component $G^{\circ}$.
Proof.
The restriction to $\mathfrak{g},\mathfrak{g}]$ equals $B$, which is $\mathrm{ad}$–invariant. If $Z\in\mathfrak{z}(\mathfrak{g})$ then $[X,Z]=0$ for all $X$, so any bilinear form on $\mathfrak{z}(\mathfrak{g})$ is automatically $\mathrm{ad}$–invariant. Using Proposition B.1 and bilinearity gives the displayed identity. Equivalence with $\mathrm{Ad}$–invariance on $G^{\circ}$ follows by integrating the infinitesimal relation along paths in $G^{\circ}$. ∎
Remark 2 (Invariance for nonconnected groups).
In case the group is nonconnected, and one desires invariance under the full group $G$ (not just $G^{\circ}$). The component group $\Gamma=G/G^{\circ}$ acts linearly on $\mathfrak{z}(\mathfrak{g})$. In all practical cases, $\Gamma$ will be a finite group. Then averaging any positive–definite $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ over $\Gamma$ yields an $\mathrm{Ad}(G)$–invariant inner product on the center:
$\langle Z_{1},Z_{2}\rangle_{\mathfrak{z}}^{\mathrm{avg}}=\frac{1}{|\Gamma|}\sum_{\gamma\in\Gamma}\big\langle\mathrm{Ad}_{\gamma}Z_{1},\,\mathrm{Ad}_{\gamma}Z_{2}\big\rangle_{\mathfrak{z}}.$ (21)
Replacing $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ by $\langle\cdot,\cdot\rangle_{\mathfrak{z}}^{\mathrm{avg}}$ in 18 makes $\widetilde{B}$ invariant under all of $G$.
Appendix C Connections to existing bilinear forms
We demonstrate how the trace-based form $\widetilde{B}$ (Eq. 2) unifies and recovers prior bilinear constructions in the regimes used by Lie Neurons (Lin et al., 2024a) and Vector Neurons (Deng et al., 2021) . The discussion below states precise conditions under which $\widetilde{B}$ (i) equals the Killing-form contractions on semisimple inputs and (ii) is proportional to the Vector Neuron inner product under the $\mathfrak{so}(3)\simeq\mathbb{R}^{3}$ isomorphism. Our single form applies on the full reductive algebra $\mathfrak{gl}(n)$ including these specialized approaches, which encompasses general $n\times n$ matrix-valued inputs.
C.1 Reduction to the Killing form on the semisimple ideal
Write $X=X_{0}+\tfrac{1}{n}\mathrm{tr}(X)I$ with $X_{0}\in\mathfrak{sl}(n)$. Using Eq. 2 we obtain
$\widetilde{B}(X,Y)=2n\,\mathrm{tr}(X_{0}Y_{0})+n^{2}\!\left(\tfrac{1}{n}\mathrm{tr}X\right)\!\left(\tfrac{1}{n}\mathrm{tr}Y\right).$ (22)
Hence, when inputs are restricted to the semisimple ideal $\mathfrak{g},\mathfrak{g}]=\mathfrak{sl}(n)$ (so $\mathrm{tr}(X)=\mathrm{tr}(Y)=0$), the center contribution vanishes and
$\widetilde{B}|_{\mathfrak{sl}(n)}(X,Y)=2n\cdot\mathrm{tr}(XY)=B_{\mathfrak{sl}(n)}(X,Y),$ (23)
i.e. $\widetilde{B}$ coincides with the (scaled) Killing form used in Lie Neurons. More generally, for any semisimple subalgebra $\mathfrak{h}\subset[\mathfrak{g},\mathfrak{g}]$, the restriction $\widetilde{B}|_{\mathfrak{h}}$ matches the Killing-form contraction on $\mathfrak{h}$ up to global scaling.
C.2 $\mathfrak{so}(3)$ example: recovery of the Vector Neuron inner product
Recall the hat map $\widehat{(\cdot)}:\mathbb{R}^{3}\to\mathfrak{so}(3)$ with
$\widehat{v}=\begin{bmatrix}0&-v_{3}&v_{2}\\ v_{3}&0&-v_{1}\\ -v_{2}&v_{1}&0\end{bmatrix},\qquad\mathrm{tr}(\widehat{v}\,\widehat{w})=-2\,v^{\top}w.$ (24)
For $X=\widehat{v}$ and $Y=\widehat{w}$, Eq. 2 gives
$\widetilde{B}(X,Y)=2n\cdot\mathrm{tr}(XY)=2n\cdot(-2\,v^{\top}w)=(-4n)\,v^{\top}w,$ (25)
i.e. $\widetilde{B}$ is proportional to the Euclidean inner product on $\mathbb{R}^{3}$. The proportionality constant depends only on $n$; in practice this constant is absorbed by adjacent learnable linear layers or normalization, yielding behaviour identical to the inner product used in Vector Neurons to harmless scaling.
These remarks justify using the single, nondegenerate $\widetilde{B}$ across heterogeneous input types while preserving compatibility with prior architectures.
Appendix D Detailed Layer Formulations
This section provides the precise mathematical definitions and equivariance proofs for the core components of the ReLN architecture. We consider the input to a layer as a tensor $x\in\mathbb{R}^{K\times C}$, representing $C$ feature channels where $K=\dim\mathfrak{g}$. Each column $x_{c}\in\mathbb{R}^{K}$ is the vector representation of a feature. We use the wedge ($\wedge$) and vee ($\vee$) operators to map between the vector form $x_{c}$ and the matrix form $X_{c}\in\mathfrak{g}$.
D.1 Equivariant Linear Layer
The ReLN-Linear layer applies a linear map to the channel dimension of the input tensor $x$:
$f_{\mathrm{ReLN-Lin}}(x;W)=xW,\quad\text{where }W\in\mathbb{R}^{C\times C^{\prime}}.$ (26)
We omit any bias term to preserve exact equivariance.
Proof of Equivariance.
The group action, $\mathrm{Ad}_{g}$ (defined in Equation 11), is a linear map that multiplies each feature channel from the left. The weight matrix $W$ multiplies the channel dimension from the right. These operations commute, ensuring strict $G$-equivariance for any $g\in G$:
$\begin{split}f_{\mathrm{ReLN-Lin}}(\mathrm{Ad}_{g}(x);W)&=(\mathrm{Ad}_{g}x)W\\ &=\mathrm{Ad}_{g}(xW)\\ &=\mathrm{Ad}_{g}(f_{\mathrm{ReLN-Lin}}(x;W)).\end{split}$ (27)
D.2 Equivariant Nonlinearities
Standard pointwise activations break equivariance under non-orthogonal transforms. We introduce two equivariant alternatives.
ReLN-ReLU.
This layer rectifies a feature based on its alignment with a learnable direction. Given the input tensor $x$, we first compute per-channel vector directions $d=xU$. The nonlinearity for the input $x$ is then defined as:
$f_{\mathrm{ReLN-ReLU}}(x)=\begin{cases}x,&\text{if }\widetilde{B}(x^{\wedge},d^{\wedge})\leq 0,\\ x+\widetilde{B}(x^{\wedge},d^{\wedge})d,&\text{otherwise}.\end{cases}$ (28)
Since all operations—the linear map to compute $d_{i}$, the bilinear form $\widetilde{B}$, and vector addition/scaling—are equivariant, the entire function is equivariant. The leaky variant $f_{\mathrm{ReLN-LeakyReLU}}(x)=\alpha\,x+(1-\alpha)\,f_{\mathrm{ReLN-ReLU}}(x)$ follows directly.
ReLN-Bracket (Lie-bracket nonlinearity).
This layer uses the matrix commutator, a natural $\mathrm{Ad}$-equivariant primitive, to create learnable interactions between channels. Let the input be a batch of features represented by their vector coordinates, $x\in\mathbb{R}^{B\times C_{\mathrm{in}}\times K}$. First, two independent linear maps (with learnable weights $W_{a},W_{b}\in\mathbb{R}^{C_{\mathrm{in}}\times C_{\mathrm{out}}}$) transform the input features into two intermediate tensors, $u,v\in\mathbb{R}^{B\times C_{\mathrm{out}}\times K}$. The Lie bracket is then computed channel-wise between the corresponding feature vectors of $u$ and $v$. This produces an update tensor, $\Delta x\in\mathbb{R}^{B\times C_{\mathrm{out}}\times K}$, where each vector is defined as:
$(\Delta x)_{b,c^{\prime},:}=\left[(u_{b,c^{\prime},:})^{\wedge},(v_{b,c^{\prime},:})^{\wedge}\right]^{\vee}.$ (29)
This update is added to the input for a residual connection (requiring $C_{\mathrm{in}}=C_{\mathrm{out}}$ for the shapes to match):
$f_{\mathrm{ReLN-Bracket}}(x)=x+\Delta x.$ (30)
Each step in this process (linear map, the Lie bracket, and vee/hat operations) is equivariant under the adjoint action, making the entire block equivariant $f_{\mathrm{ReLN-Bracket}}(\mathrm{Ad}_{g}x)=\mathrm{Ad}_{g}(f_{\mathrm{ReLN-Bracket}}(x))$ for all $g\in G$.
Appendix E Experimental Details
Training and evaluation for all presented experiments, Platonic solid classification, invariant function regression, top tagging, and drone state estimation, were conducted on a single NVIDIA GeForce RTX 4090 GPU.
E.1 Model Architectures and Implementation Details
Across all experiments, our proposed ReLN models are constructed by stacking ReLN-Linear, ReLN-ReLU, and ReLN-Bracket layers. The specific number of layers and channel widths are adapted for each task to ensure a fair comparison with baseline models in terms of parameter count.
Algebraic Benchmarks ($\mathfrak{sl}(3)$ and $\mathfrak{sp}(4)$).
For the Platonic solid classification and $\mathfrak{sp}(4)$ invariant regression tasks, our ReLN model directly adopts the architecture used by the Lie Neurons benchmark model from Lin et al. (2024a). The primary modification is the replacement of their Killing form-based nonlinearity and invariant layers with our proposed non-degenerate bilinear form $\widetilde{B}$ (Eq. 2). This setup allows for a direct comparison of the impact of the bilinear form, as all other architectural hyperparameters are kept identical to the baseline.
Top Tagging.
For the Top-Tagging task, our model is a modification of the LorentzNet architecture (Gong et al., 2022). We adapt its Lorentz Group Equivariant Blocks (LGEBs) by replacing the invariant feature computation with our proposed bilinear form. A detailed description of the architecture, our modifications, and training protocol is provided in Appendix F.
Drone State Estimation.
In this task, we compare our ReLN model against two baseline families: a non-equivariant 1D ResNet and an equivariant Vector Neurons (VN) model. The specific implementation details and architectural choices for each model are provided next in Appendix G.
E.2 Platonic Solid Classification on $\mathfrak{sl}(3)$
Overview.
All experiments evaluate classification of Platonic solids (tetrahedron, octahedron, icosahedron) from inter-face homographies computed in the image plane. For each model-family we train 5 independent runs with different random seeds and report mean ± standard deviation. Training uses fixed object and camera poses; at test time we report results on the in-distribution (ID) split and the rotated-camera (RC) split (RC applies ten random $\mathrm{SO}(3)$ rotations to the camera frame). The ‘MLP (wider)’ denotes a capacity-matched ($\approx$ 2× parameters) MLP used for a fairer comparison.
Table 6: Common training hyperparameters (used across model families unless noted).
Hyperparameter Value
Optimizer Adam
Batch size 100
Number of independent runs (seeds) 5
Max epochs / stopping criterion 5000 epochs
Data augmentation (train) Random camera rotations applied to training examples when enabled
RC evaluation 500 random $\mathrm{SO}(3)$ rotations applied to each test example
Metric reported Classification accuracy (mean ± std across runs)
Table 7: Model-specific hyperparameters and implementation notes.
Model family Key choices Notes
Latent Feature Size (MLP Baseline) 256 As in Lin et al. (2024a).
Latent Feature Size (MLP Wider) 386 Increased width total parameters $\approx 2\times$ baseline.
Learning rate (MLP models) $1\times 10^{-4}$
Learning rate (ReLNs/Lie Neurons models) $3\times 10^{-6}$ Lower LR chosen for stable training
Appendix F Top Tagging Experiment: Framework, Proof, and Implementation
This appendix provides the complete details for our jet tagging experiment. We first present the geometric framework and the mathematical proof of our Lorentz-equivariant embedding, and then describe the model architecture and training protocol.
F.1 Group Action Equivariance via Embedding Map
To process four-momenta within our Lie-algebraic framework, we require an embedding that translates the action of the Lorentz group into an adjoint action on a matrix space. This is achieved by lifting the four-vector into $\mathfrak{gl}(5)$.
Definition F.1 (Lorentz-Compatible Embedding).
Given a four-vector $p\in\mathbb{R}^{4}$ and the Minkowski metric $\eta=\operatorname{diag}(1,-1,-1,-1)$, we define its embedding $\varphi(p)$ into $\mathfrak{gl}(5)$ as:
$\varphi(p)=\begin{bmatrix}0_{4\times 4}&p\\ p^{\top}\eta&0\end{bmatrix}$. (31)
Theorem F.1 (Adjoint Equivariance).
The embedding $\varphi$ correctly models the Lorentz group action. For any $p\in\mathbb{R}^{4}$ and Lorentz transformation $\Lambda\in\operatorname{SO}(1,3)$, let $G=\operatorname{diag}(\Lambda,1)\in\operatorname{GL}(5)$. The map is equivariant in the sense that the standard action on $p$ corresponds to the adjoint action on $\varphi(p)$:
$\operatorname{Ad}_{G}(\varphi(p))=G\varphi(p)G^{-1}=\varphi(\Lambda p).$ (32)
Proof.
We compute the left-hand side (LHS) of Eq. 32, which is the adjoint action:
$\operatorname{Ad}_{G}(\varphi(p))$ $=\begin{bmatrix}\Lambda&0\\ 0&1\end{bmatrix}\begin{bmatrix}0&p\\ p^{\top}\eta&0\end{bmatrix}\begin{bmatrix}\Lambda^{-1}&0\\ 0&1\end{bmatrix}=\begin{bmatrix}0&\Lambda p\\ p^{\top}\eta\Lambda^{-1}&0\end{bmatrix}$. (33)
The right-hand side (RHS) is the lift of the transformed vector $\Lambda p$:
$\varphi(\Lambda p)=\begin{bmatrix}0&\Lambda p\\ (\Lambda p)^{\top}\eta&0\end{bmatrix}=\begin{bmatrix}0&\Lambda p\\ p^{\top}\Lambda^{\top}\eta&0\end{bmatrix}$. (34)
For the LHS and RHS to be equal, we must show that $\eta\Lambda^{-1}=\Lambda^{\top}\eta$. We start from the defining property of $\operatorname{SO}(1,3)$:
$\Lambda^{\top}\eta\Lambda=\eta.$ (35)
Right-multiplying Eq. 35 by $\Lambda^{-1}$ yields the desired identity:
$(\Lambda^{\top}\eta\Lambda)\Lambda^{-1}$ $=\eta\Lambda^{-1}\implies\Lambda^{\top}\eta(\Lambda\Lambda^{-1})=\eta\Lambda^{-1}\implies\Lambda^{\top}\eta=\eta\Lambda^{-1}.$ (36)
Since the condition holds, the proof is complete. ∎
Remark 3 (Generalization to Orthogonal Groups).
This embeddinging technique is not limited to the Lorentz group and can be readily generalized to any orthogonal group $\operatorname{O}(n)$ or special orthogonal group $\operatorname{SO}(n)$. For instance, in applications involving 3D point clouds where the symmetry is $\operatorname{SO}(3)$, a vector $p\in\mathbb{R}^{3}$ would be embedded into the Lie algebra $\mathfrak{gl}(4)$ as:
$\varphi(p)=\begin{bmatrix}0_{3\times 3}&p\\ p^{\top}&0\end{bmatrix}$ (37)
The proof of equivariance follows the same structure, using the property of orthogonal matrices, $R^{\top}R=I$ (which implies $R^{-1}=R^{\top}$), instead of the Minkowski metric identity. This highlights the broad applicability of our embedding strategy to any benchmark involving norm-preserving group transformations.
F.2 Experimental Implementation
Dataset
The experiment uses the Top-Tagging dataset (Kasieczka et al., 2019), which contains 2 million simulated proton-proton collision events. The dataset was generated with Pythia, Delphos, and FastJet to model the ATLAS detector response. We use the standard 60%/20%/20% splits for training, validation, and testing. Each jet is represented as a set of constituent particles, each with four-momentum $p=(E,p_{x},p_{y},p_{z})$.
Model
Our model leverages the established architecture of LorentzNet  (Gong et al., 2022), utilizing its stack of Lorentz Group Equivariant Blocks (LGEBs) for message passing on the jet’s particle cloud. While the original LorentzNet computes these features directly from the 4-momenta using the Minkowski inner product, our approach introduces a modified bilinear form based feature extraction. We first embed each pair of 4-momenta, $p_{i}$ and $p_{j}$, from the Minkowski space $\mathbb{R}^{1,3}$ into the Lie algebra $\mathfrak{gl}(5)$ via the map $p\mapsto X(p)$. The invariant features for the message passing are then derived from the bilinear form, $B(\cdot,\cdot)$, on this Lie algebraic space. The edge message $m_{ij}$ is thus constructed as:
$m_{ij}=\phi_{e}\Big(h_{i},h_{j},\psi\big(\widetilde{B}(X(p_{i}),X(p_{i}))\big),\psi\big(\widetilde{B}(X(p_{i}),X(p_{j}))\big)\Big)$ (38)
where $h_{i},h_{j}$ are scalar features, $\phi_{e}$ is an MLP, and $\psi$ is a stabilizing nonlinearity. As shown in the main results (Table 4), this approach leads to an advantage in background rejection when compared against a parameter-matched LorentzNet baseline. The architectural differences are summarized in Table 8.
Table 8: Architectural comparison for the Top-Tagging task.
Component LorentzNet (Original) Param-matched Baseline Ours (ReLN)
Number of LGEBs 6 5 5
Hidden feature dims 72 48 48
Edge feature computation Minkowski inner prod. Minkowski inner prod. Bilinear invariant form
Training Setup
For a fair comparison, our training procedure closely follows the protocol established in the LorentzNet (Gong et al., 2022). The model was trained for a total of 35 epochs on a single NVIDIA RTX 4090 GPU. We used the AdamW optimizer with a weight decay of 0.01 and a batch size of 128, matching the total effective batch size from the reference work. The learning rate was managed by the paper’s specific three-stage schedule: a 4-epoch linear warm-up to an initial rate of $1\times 10^{-3}$, followed by a 28-epoch CosineAnnealingWarmRestarts schedule, and a final 3-epoch exponential decay. After each epoch, the model with the highest validation accuracy was saved for final evaluation on the test set.
Appendix G Drone Experiment Details
This appendix provides the technical details for the drone state estimation experiment, including the theoretical framework, dataset generation, model implementations, and formal proofs.
G.1 Geometric Framework for Equivariant Covariance Processing
Our approach leverages the geometry of symmetric positive-definite matrices. A covariance matrix $C$ is symmetric positive-definite, residing on the manifold $\mathrm{SPD}(3)$. A non-degenerate covariance matrix $C\in\mathrm{SPD}(n)$ represents the anisotropic stretching of a general linear map, as seen via the polar decomposition $A=QP$ with $Q\in\mathrm{O}(n)$ and $P\in\mathrm{SPD}(n)$. Equivalently, there is a homogeneous-space isomorphism: $\mathrm{SPD}(n)\;\cong\;\mathrm{GL}(n)/\mathrm{O}(n),$ which motivates processing covariances in a $\mathrm{GL}(n)$-aware architecture.
While $\mathrm{SPD}(3)$ is not a Lie group, the matrix logarithm provides a canonical map to the vector space of symmetric matrices $\mathrm{Sym}(3)$, which is a linear subspace of $\mathfrak{gl}(3)$.
$\log:\mathrm{SPD}(3)\;\longrightarrow\;\mathrm{Sym}(3)\subset\mathfrak{gl}(3).$ (39)
This allows us to embed a geometric object from a curved manifold into a flat, Lie-algebra-compatible space. The following theorem proves that the congruence transformation on $C\in\mathrm{SPD}(n)$ becomes an adjoint action on its image $\log C\in\mathrm{Sym}(n)$, thus preserving the equivariant structure required by our model.
Theorem G.1 (Equivariance of the Logarithmic Map).
For any $C\in\mathrm{SPD}(n)$ and any rotation matrix $R\in\mathrm{SO}(n)$, the congruence transformation on $C$ corresponds to an adjoint action on its logarithm:
$\log(RCR^{\top})=R(\log C)R^{\top}.$ (40)
Proof.
The proof follows from the spectral theorem for real symmetric matrices.
  1. 1.
    Let the eigendecomposition of $C$ be $C=V\Lambda V^{\top}$, where $V$ is an orthogonal matrix ($V^{\top}V=I$) of eigenvectors and $\Lambda$ is the diagonal matrix of corresponding positive eigenvalues.
  2. 2.
    By definition, the matrix logarithm of $C$ is given by applying the logarithm to its eigenvalues:
    $\log C:=V(\log\Lambda)V^{\top}$ (41)
    where $\log\Lambda$ is the diagonal matrix of element-wise logarithms of the eigenvalues.
  3. 3.
    Consider the transformed matrix $C^{\prime}=RCR^{\top}$. Substituting the decomposition of $C$ yields:
    $C^{\prime}=R(V\Lambda V^{\top})R^{\top}=(RV)\Lambda(V^{\top}R^{\top})=(RV)\Lambda(RV)^{\top}$ (42)
    This is the eigendecomposition of $C^{\prime}$, where the new orthogonal matrix of eigenvectors is $V^{\prime}=RV$ and the eigenvalues $\Lambda$ are unchanged.
  4. 4.
    Applying the definition of the matrix logarithm to $C^{\prime}$ gives:
    $\log(C^{\prime})=V^{\prime}(\log\Lambda)(V^{\prime})^{\top}=(RV)(\log\Lambda)(RV)^{\top}$ (43)
  5. 5.
    Rearranging the terms, we arrive at the desired identity:
    $\log(C^{\prime})=R\left(V(\log\Lambda)V^{\top}\right)R^{\top}=R(\log C)R^{\top}$ (44)
This identity is critical, as it confirms that our adjoint-equivariant network can process either the raw covariance $C$ or its logarithm $\log C$ while perfectly preserving the $\mathrm{SO}(3)$ symmetry.
In the $\mathrm{SO}(3)$ regime used in our experiments, vectors (e.g., velocity $\mathbf{v}$) are represented in the Lie algebra $\mathfrak{so}(3)$ so that the adjoint action coincides with ordinary rotation, $\mathrm{Ad}_{R}(\mathbf{v})=R\mathbf{v}$. Conjugation then implements the covariance congruence $C\mapsto RCR^{\top}$. Consequently, ReLNs realize $\mathrm{SO}(3)$-equivariance by construction, avoiding the need for the model to learn these symmetries from data.
G.2 Dataset Generation.
We use the PyBullet engine to simulate 200 aggressive trajectories for a Crazyflie-like nano-quadrotor. To generate realistic measurements, the instantaneous velocity is corrupted by Gaussian noise, $\mathbf{v}_{\mathrm{noisy}}\sim\mathcal{N}(\mathbf{v}_{\mathrm{gt}},\,C_{v}),$ where the covariance $C_{v}$ varies with flight aggressiveness. The dataset provides time series of noisy velocities, ground-truth covariances, and ground-truth trajectories for evaluation.
Trajectory Generation.
The procedure begins with the procedural generation of a sequence of 20 to 40 random 3D waypoints within a flight volume of approximately $170\text{m}\times 170\text{m}\times 60\text{m}$. The waypoints are sampled from a uniform distribution to create diverse flight paths. To mimic the complex dynamics of aggressive flight, each trajectory is randomly generated using a path with random wiggles or a path featuring high-speed spiral maneuvers. These discrete waypoints are then interpolated using a Catmull-Rom spline to create a smooth, $C^{1}$ continuous target trajectory, which is densely sampled at an 80 Hz control frequency. Each of the 200 sequences results in a unique trajectory lasting approximately 2-4 minutes, totaling over 13 hours of simulated flight time. A sample generated trajectory is shown in Figure 5.
(a) A sample trajectory with spiral maneuvers.
(b) A sample trajectory with random wiggles.
Figure 5: Sample aggressive trajectories generated in the PyBullet simulator.
State-Dependent Noise Model.
To simulate realistic sensor characteristics, the ground-truth velocity is corrupted by zero-mean Gaussian noise, $\mathbf{v}_{\mathrm{noisy}}\sim\mathcal{N}(\mathbf{v}_{\mathrm{gt}},\,C_{v})$. The covariance matrix $C_{v}$ is state-dependent, designed to scale with the drone’s speed. The standard deviation $\sigma_{v}$ for each velocity axis is computed using a sigmoid function of the velocity magnitude $\|\mathbf{v}_{\mathrm{gt}}\|$:
$\sigma_{v}(\|\mathbf{v}_{\mathrm{gt}}\|)=\sigma_{\min}+(\sigma_{\max}-\sigma_{\min})\cdot\frac{1}{1+\exp(-\lambda(\|\mathbf{v}_{\mathrm{gt}}\|-v_{\text{mid}}))},$ (45)
where the variance on each axis is $\sigma_{v}^{2}$. We set the minimum and maximum standard deviations to $\sigma_{\min}=0.2$ m/s and $\sigma_{\max}=1.0$ m/s, respectively. The steepness $\lambda$ is set to 0.8, and the midpoint velocity $v_{\text{mid}}$ is dynamically adjusted based on the estimated average speed of each trajectory to ensure a realistic noise profile.
G.3 Baseline and Model Implementation Details
We compare ReLN against two baseline classes chosen to isolate the effect of geometric priors.
Non-equivariant baselines.
We use a standard 1D ResNet architecture with temporal convolutional blocks that processes flattened input sequences. The ResNet (velocity-only) model receives only the 3D velocity vector. The ResNet (velocity + covariance) model receives the flattened $3\times 3$ covariance matrix concatenated to the velocity vector.
Eigendecomposition-based $\mathrm{SO}(3)$-Equivariant Baseline.
This model adapts the 1D ResNet backbone for $\mathrm{SO}(3)$ equivariance using VN layers. Since VNs cannot directly ingest matrices, we decompose each covariance matrix $C=V\Lambda V^{\top}$ and use a dual-stream design:
  • an equivariant stream $\mathcal{F}_{\mathrm{eq}}=\{\mathbf{v},\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}\}$ comprising the measured velocity $\mathbf{v}$ and the three orthonormal eigenvectors $\mathbf{e}_{i}$, which together capture all directional information. This stream is handled by the VNs backbone.
  • an invariant stream $\mathcal{F}_{\mathrm{inv}}=\{\lambda_{1},\lambda_{2},\lambda_{3}\}$ processes the corresponding eigenvalues $\{\lambda_{1},\lambda_{2},\lambda_{3}\}$, which encode orientation-independent scale information, using a standard MLP.
The two latent features from both streams are fused at the final output layer. Eigenvector ambiguities (sign or multiplicities) are resolved via a deterministic, rotation-equivariant canonicalization.
Reductive Lie Neurons (ReLNs).
The ReLN model shares the same base architecture as the VN model but incorporates ReLN-Bracket layer as an additional source of nonlinearity after the initial feature extraction block. In contrast to the VN backbone, ReLNs provide a unified framework for velocity and covariance processing. Velocities $\mathbf{v}\in\mathbb{R}^{3}$ are lifted into the Lie algebra $\mathfrak{so}(3)$ via $K=\mathbf{v}^{\wedge}$, while covariance $C$ (or $\log C$) is treated as a structured geometric input. Both transform under the same adjoint action: $K^{\prime}=RKR^{\top}$ and $C^{\prime}=RCR^{\top}$, enabling joint equivariant processing. Although $\mathrm{SPD}(n)$ is not a Lie algebra or group, it is a subset of $\mathrm{GL}(n)$. By taking the matrix logarithm, $\log C\in\mathrm{Sym}(n)\subset\mathfrak{gl}(n)$, covariances are embedded into a linear subspace compatible with Lie-algebra processing.
The network $\Phi$ fuses these inputs into a single matrix $A\in\mathbb{R}^{3\times 3}$, from which we extract the velocity estimate:
$A_{\mathrm{skew}}\coloneqq\tfrac{1}{2}(A-A^{\top})\in\mathfrak{so}(3),\qquad\hat{\mathbf{v}}=\text{Vee}(A_{\mathrm{skew}}).$ (46)
The extracted velocity is provably $\mathrm{SO}(3)$-equivariant by construction. See Appendix H for the full statement and proof.
G.4 Training and Evaluation Protocol
Problem Formulation
The network is trained to predict the drone’s 3D position $\mathbf{p}_{t}\in\mathbb{R}^{3}$ at the end of a given time window, based on a sequence of noisy velocity measurements and their corresponding covariances within that window (e.g., a 1-second history). All models are trained by minimizing the Mean Squared Error (MSE) between the predicted position $\hat{\mathbf{p}}_{t}$ and the ground-truth position $\mathbf{p}_{t,\text{gt}}$. The loss function is defined as $\mathcal{L}=\|\hat{\mathbf{p}}_{t}-\mathbf{p}_{t,\text{gt}}\|_{2}^{2}$.
Dataset and Optimization.
We partition the dataset using a standard 80:10:10 train/validation/test split. All models are trained on identical splits to ensure fair comparison. Models are optimized using the AdamW optimizer with a ReduceLROnPlateau learning rate scheduler based on validation loss.
Evaluation Metrics.
We report the following pose-regression metrics over the test set:
  • Absolute Trajectory Error (ATE): The root-mean-square error between the ground-truth and predicted 3D positions over the entire trajectory, measured in meters.
  • ATE%: The ATE normalized by the total trajectory length and expressed as a percentage ($100\times\text{ATE}/\text{length}$). This metric provides a scale-invariant measure of error, which is crucial for fairly comparing performance across our aggressive flight trajectories of varying lengths.
  • Relative Pose Error (RPE): The error measured over fixed-length sub-trajectories, capturing local drift.
To explicitly validate equivariance, we also evaluate all models on the test set after applying a set of random $\mathrm{SO}(3)$ rotations to the entire input sequence.
G.5 Eigenvector Canonicalization for the VN Baseline
To resolve ambiguities in the eigendecomposition $C=V\Lambda V^{\top}$ for the VN baseline, we canonicalize the eigenvector matrix $V=[\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}]$ as follows:
  1. 1.
    Right-handed Frame: If $\det V<0$, we set $\mathbf{e}_{3}\leftarrow-\mathbf{e}_{3}$ to ensure $\det V=+1$.
  2. 2.
    Sign Disambiguation: For each eigenvector $\mathbf{e}_{i}$ with a distinct eigenvalue, we enforce a consistent sign by ensuring $\mathbf{v}^{\top}\mathbf{e}_{i}\geq 0$. If not, we set $\mathbf{e}_{i}\leftarrow-\mathbf{e}_{i}$.
  3. 3.
    Multiplicity Handling: In the rare case of repeated eigenvalues, we use the projection of the velocity vector $\mathbf{v}$ onto the corresponding eigenspace to deterministically define the first basis vector, then complete the basis via Gram-Schmidt.
All steps use only equivariant operations, preserving the overall symmetry of the baseline.
Appendix H Proof of $\mathrm{SO}(3)$-Equivariance for ReLN Velocity Extract with Covariance Inputs
This section provides a formal proof for the $\mathrm{SO}(3)$-equivariance of our Reductive Lie Neuron (ReLN) architecture when processing a velocity vector and a covariance matrix. We first establish the foundations for processing covariance matrices within a Lie-algebraic framework and then present the main proof.
H.1 $\mathrm{SO}(3)$-Equivariant Vector Extraction via Skew-Symmetric Projection
Our network, $\Phi$, is designed to be adjoint-equivariant. It maps geometric inputs—such as a embedded velocity $K\in\mathfrak{so}(3)$ and a covariance matrix $S\in\mathrm{SPD}(3)$—to a matrix feature $A\in\mathbb{R}^{3\times 3}$. The inputs transform under the adjoint action of any rotation $R\in\mathrm{SO}(3)$:
$K^{\prime}=\mathrm{Ad}_{R}(K)=RKR^{\top},\quad S^{\prime}=\mathrm{Ad}_{R}(S)=RSR^{\top}.$ (47)
By construction, the network’s output feature $A$ transforms according to the same law:
$\Phi(K^{\prime},S^{\prime})=\mathrm{Ad}_{R}\big(\Phi(K,S)\big)=R\,\Phi(K,S)\,R^{\top}.$ (48)
To obtain the final 3D velocity vector, we project the output matrix $A$ onto its skew-symmetric component and apply the vee operator. The following proposition formalizes the equivariance of this extraction mechanism.
Proposition H.1 (Equivariance of Skew-Symmetric Extraction).
Let a network $\Phi$ and its inputs transform according to Eqs. 47 and 48. If a vector $\hat{\mathbf{v}}\in\mathbb{R}^{3}$ is extracted from the output matrix $A=\Phi(K,S)$ via the projection
$A_{\mathrm{skew}}=\tfrac{1}{2}(A-A^{\top}),\qquad\hat{\mathbf{v}}=(A_{\mathrm{skew}})^{\vee},$ (49)
then the vector $\hat{\mathbf{v}}^{\prime}$ extracted from the transformed output $A^{\prime}=\Phi(K^{\prime},S^{\prime})$ transforms covariantly as $\hat{\mathbf{v}}^{\prime}=R\hat{\mathbf{v}}$.
Proof.
By the adjoint-equivariance property in Eq. 48, the network satisfies $\Phi(RKR^{\top},RCR^{\top})=R\,\Phi(K,C)\,R^{\top}=RAR^{\top}$. Let $A^{\prime}=RAR^{\top}$. The skew-symmetric component of the transformed output $A^{\prime}$ is:
$\begin{split}A^{\prime}_{\mathrm{skew}}&=\tfrac{1}{2}(A^{\prime}-A^{\prime\top})\\ &=\tfrac{1}{2}\big(RAR^{\top}-(RAR^{\top})^{\top}\big)\\ &=\tfrac{1}{2}\big(RAR^{\top}-RA^{\top}R^{\top}\big)\\ &=R\left(\tfrac{1}{2}(A-A^{\top})\right)R^{\top}\\ &=RA_{\mathrm{skew}}R^{\top}=\mathrm{Ad}_{R}(A_{\mathrm{skew}}).\end{split}$ (50)
The vee map, $(\cdot)^{\vee}:\mathfrak{so}(3)\to\mathbb{R}^{3}$, is itself an equivariant map satisfying $(\mathrm{Ad}_{R}(X))^{\vee}=R\,(X^{\vee})$ for any $X\in\mathfrak{so}(3)$. Applying this property yields the desired result:
$\hat{\mathbf{v}}^{\prime}=(A^{\prime}_{\mathrm{skew}})^{\vee}=(\mathrm{Ad}_{R}(A_{\mathrm{skew}}))^{\vee}=R\,(A_{\mathrm{skew}})^{\vee}=R\hat{\mathbf{v}}$. (51)
Remark 4.
The proof relies on three properties: (i) both inputs transform under the adjoint action $X\mapsto RXR^{\top}$; (ii) the network $\Phi$ is equivariant to this action; and (iii) the output is projected onto $\mathfrak{so}(3)$ before the vee operator is applied. As established previously, these conditions hold whether the network ingests the raw covariance $S$ or its logarithm $\log S$.
Large Language Model (LLM) Usage
We used a large language model (LLM) to aid in polishing the writing and improving grammatical clarity of the manuscript. The LLM did not contribute to the research ideation, experiments, or technical content; all scientific claims and results were generated solely by the authors.

### v2

@@
 Abstract
-Encoding symmetries is a powerful inductive bias for improving the generalization of deep neural networks.
-However, most existing equivariant models are limited to simple symmetries like rotations, failing to address the broader class of general linear transformations, $\mathrm{GL}(n)$, that appear in many scientific domains.
-We introduce Reductive Lie Neurons (ReLNs), a novel neural network architecture exactly equivariant to these general linear symmetries.
-ReLNs are designed to operate directly on a wide range of structured inputs, including general $n$-by-$n$ matrices.
-ReLNs introduce a novel adjoint-invariant bilinear layer to achieve stable equivariance for both Lie-algebraic features and matrix-valued inputs, without requiring redesign for each subgroup.
-This architecture overcomes the limitations of prior equivariant networks that only apply to compact groups or simple vector data.
-We validate ReLNs’ versatility across a spectrum of tasks: they outperform existing methods on algebraic benchmarks with $\mathfrak{sl}(3)$ and $\mathfrak{sp}(4)$ symmetries and achieve competitive results on a Lorentz-equivariant particle physics task.
-In 3D drone state estimation with geometric uncertainty, ReLNs jointly process velocities and covariances, yielding significant improvements in trajectory accuracy.
-ReLNs provide a practical and general framework for learning with broad linear group symmetries on Lie algebras and matrix-valued data.
+Many scientific and geometric problems exhibit general linear symmetries, yet most equivariant neural networks are built for compact groups or simple vector features, limiting their reuse on matrix-valued data such as covariances, inertias, or shape tensors.
+We introduce Reductive Lie Neurons (ReLNs), an exactly $\mathrm{GL}(n)$-equivariant architecture that natively supports matrix-valued and Lie-algebraic features.
+ReLNs resolve a central stability issue for reductive Lie algebras by introducing a non-degenerate adjoint (conjugation)-invariant bilinear form, enabling principled nonlinear interactions and invariant feature construction in a single architecture that transfers across subgroups without redesign.
+We demonstrate ReLNs on algebraic tasks with $\mathfrak{sl}(3)$ and $\mathfrak{sp}(4)$ symmetries, Lorentz-equivariant particle physics, uncertainty-aware drone state estimation via joint velocity–covariance processing, learning from 3D Gaussian-splat representations, and EMLP double-pendulum benchmark spanning multiple symmetry groups.
+ReLNs consistently match or outperform strong equivariant and self-supervised baselines while using substantially fewer parameters and compute, improving the accuracy–efficiency trade-off and providing a practical, reusable backbone for learning with broad linear symmetries.
 Project page: https://reductive-lie-neuron.github.io/
@@
 1 Introduction
-Leveraging the symmetries present in real-world data improves the generalization of neural networks.
-When a model is equivariant, i.e., its outputs transform predictably with inputs under certain group actions, it can learn more efficiently and generalize across scenarios with less data.
-Most advances in equivariant deep learning target compact symmetry groups, such as rotations, $\mathrm{SO}(3)$, or isometries, $\mathrm{E}(n)$, where strong geometric and algebraic tools support robust network design (Bronstein et al., 2021; Cohen & Welling, 2016; Thomas et al., 2018; Satorras et al., 2021; Finzi et al., 2020; Geiger & Smidt, 2022).
-These methods have led to practical gains.
-However, many real-world phenomena, such as affine distortions, sensor anisotropy, and general changes of basis, are governed by the broader general linear group $\mathrm{GL}(n)=\{A\in\mathbb{R}^{n\times n}:\det(A)\neq 0\}$, for which existing methods fail to provide exact equivariance.
-These more general symmetries are crucial in a wide array of scientific domains, including robotics, particle physics, and computer vision, a landscape we survey in Table 1.
-Naive solutions, such as flattening matrices into vectors, discard essential geometric structure, while methods based on spectral decompositions are brittle and numerically unstable  (Magnus, 1985).
-This leaves a gap for a numerically stable neural architecture capable of processing diverse geometric data types equivariantly under arbitrary invertible linear transformations (Bronstein et al., 2021).
-Figure 1: Examples of Lie groups and related manifolds in scientific applications.
-From left: the special linear group $\mathrm{SL}(3)$ (image homography), the Lorentz group $\mathrm{SO}(1,3)$ (spacetime symmetry), symplectic groups $\mathrm{Sp}(n)$ (Hamiltonian mechanics), the $\mathrm{SPD}(3)\oplus\mathbb{R}^{3}$ state space (probabilistic estimation), and the general linear group $\mathrm{GL}(3)$ (modeling stress-strain in continuum mechanics).
-See Table 1 for a more detailed survey.
-Table 1: A survey of common Lie groups and their applications in equivariant deep learning.
-Group / Data Structure Application Reference
-$\mathrm{SO}(3)$ 3D Point Clouds, State Estimation Lin et al. (2024a); Deng et al. (2021); Son et al. (2024)
-$\mathrm{SO}^{+}(1,3)$ Particle Physics, Jet Tagging Bogatskiy et al. (2020); Finzi et al. (2021); Batatia et al. (2023)
-$\mathrm{SU}(3)$ Quantum Chromodynamics (QCD) Favoni et al. (2022)
-$\mathrm{SL}(3)$ Homography Classification, 3D Vision Lin et al. (2024a); Finzi et al. (2021)
-$\mathrm{Sp}(4,\mathbb{R})$ Hamiltonian dynamics Lin et al. (2024a); Finzi et al. (2021)
-$\mathrm{SPD}(n)$ Geometric Uncertainty Processing -
-$\mathrm{GL}(n)$ General Linear Transformations Basu et al. (2025); Finzi et al. (2021)
-Not a group. $\mathrm{SPD}(n)$ is the manifold of symmetric positive-definite matrices, representable as the quotient space $\mathrm{GL}(n)/\mathrm{O}(n)$.
-This work introduces Reductive Lie Neurons (ReLNs), a unified neural architecture that provides exact equivariance to the adjoint action of the general linear group, $\mathrm{GL}(n)$, and its reductive subgroups, enabling robust learning on both Lie-algebraic features and structured geometric data like covariances.
-We resolve the central technical obstacle for non-semisimple algebras—the degeneracy of the canonical Killing form—by introducing a learnable, provably non-degenerate, and $\mathrm{Ad}$-invariant bilinear form.
-This provides a mathematically stable mechanism for defining symmetry-preserving nonlinearities, pooling, and invariant layers applicable to the reductive Lie algebra $\mathfrak{gl}(n)=\mathbb{R}^{n\times n}$.
+Exploiting symmetries in data is a fundamental principle of geometric deep learning.
+By enforcing equivariance, where model outputs transform predictably with inputs, neural networks can leverage strong inductive biases to improve data efficiency and generalization.
+Substantial progress has been made for compact symmetry groups such as rotations $\mathrm{SO}(3)$ and Euclidean isometries $\mathrm{E}(n)$ (Bronstein et al., 2021; Cohen and Welling, 2016; Thomas et al., 2018; Satorras et al., 2021; Geiger and Smidt, 2022).
+However, many geometric and physical systems exhibit broader linear symmetries that are naturally expressed by the non-compact general linear group $\mathrm{GL}(n)=\{A\in\mathbb{R}^{n\times n}:\det(A)\neq 0\}$.
+Despite its ubiquity, general-purpose, exact equivariant architectures for $\mathrm{GL}(n)$ remain comparatively underdeveloped.
+Appendix D summarizes these groups and applications.
+A key challenge for $\mathrm{GL}(n)$-equivariant learning is non-compactness.
+Unlike compact groups (e.g., $\mathrm{SO}(3)$), $\mathrm{GL}(n)$ does not come with a canonical $\mathrm{Ad}$-invariant positive-definite inner product: the Killing form is degenerate on $\mathfrak{gl}(n)$, and one loses the norm-preserving (unitary/orthogonal) structure that standard steerable irrep-based models rely on in practice (Thomas et al., 2018; Geiger and Smidt, 2022).
+At the Lie-algebra level, $\mathfrak{gl}(n)=\mathbb{R}^{n\times n}$ is reductive but not semisimple, so this degeneracy removes a canonical invariant metric and complicates stable invariant scalars for equivariant nonlinear.
+As a result, existing approaches either restrict attention to semisimple algebras or rely on ad hoc choices that do not provide a general, stable recipe for $\mathrm{GL}(n)$ adjoint equivariance.
+Beyond this algebraic obstacle, real-world geometric inputs are often heterogeneous, with different attributes obeying different transformation laws.
+Vectors transform by a left action (e.g., $v\mapsto Av$), whereas matrix-valued tensors such as uncertainty, inertia, or shape covariances transform by congruence (e.g., $\Sigma\mapsto A\Sigma A^{\top}$).
+Learning to jointly couple vector–matrix channels in a symmetry-consistent and numerically stable way remains challenging: naive flattening discards tensor structure, and eigen/spectral parameterizations can be ambiguous and fragile under differentiation (Magnus, 1985).
+These issues arise in robotics state estimation, where velocities and covariance uncertainty must be processed together, and in 3D Gaussian Splatting, which couples a mean $\mu\in\mathbb{R}^{3}$ with an anisotropic covariance $\Sigma\in\mathrm{SPD}(3)$.
+In this paper, we introduce Reductive Lie Neurons (ReLNs), a general-purpose architecture that is exactly equivariant to the adjoint action of $\mathrm{GL}(n)$ on $\mathfrak{gl}(n)$.
+ReLNs are built around a non-degenerate $\mathrm{Ad}$-invariant bilinear form that resolves Killing-form degeneracy on reductive algebras and enables stable invariant gating/normalization and expressive equivariant nonlinearities within a single backbone.
+We evaluate ReLNs on algebraic benchmarks ($\mathfrak{sl}(3)$, $\mathfrak{sp}(4)$), uncertainty-aware drone state estimation with joint velocity–covariance processing, learning from 3D Gaussian-splat representations with comparisons to Gaussian-MAE, and the EMLP double-pendulum benchmark; additional Lorentz-equivariant results are deferred to the appendix.
 Our main contributions are:
   1. 1.
-    We propose Reductive Lie Neurons (ReLNs), a novel, general-purpose, and numerically stable network architecture for exact $\mathrm{GL}(n)$ adjoint equivariance.
+    We propose Reductive Lie Neurons (ReLNs), a novel, general-purpose architecture for exact $\mathrm{GL}(n)$ adjoint equivariance, built upon a $\mathrm{Ad}$-invariant bilinear form that resolves the degeneracy issues of the standard Killing form on reductive algebras.
   2. 2.
-    We establish a connection between classical left-action equivariance and our adjoint-action framework for orthogonal groups.
-Through a provably-equivariant embedding map, we show that problems defined on standard vector actions—such as Lorentz transformations in particle physics or 3D point cloud processing—can be solved within our unified architecture, obviating the need for specialized model designs.
+    We introduce a unified lifting that embeds heterogeneous geometric inputs, left-acting vectors and congruence-transforming matrices (e.g., covariances, 3D Gaussian parameters), into a shared $\mathfrak{gl}(n)$ feature space, enabling a single adjoint-equivariant backbone with exact equivariance.
   3. 3.
-    We establish the framework for geometric uncertainty-aware equivariant learning, enabling models to treat matrix-valued data that transforms under congruence (e.g., covariance tensors) as geometric objects.
-  4. 4.
-    We demonstrate the effectiveness of ReLNs through extensive experiments, showing that they outperform prior methods on Lie-algebraic benchmarks and achieve significant improvements in accuracy and robustness on a challenging 3D drone state estimation task.
-2 Related Work
-Encoding symmetry into neural architectures is a powerful inductive bias that improves data efficiency and generalization (Bronstein et al., 2021).
-In geometric deep learning, the most mature area of research focuses on enforcing equivariance to Euclidean isometries, transformations such as rotations and rigid motions formalized by Lie groups like $\mathrm{SO}(n)$ and $\mathrm{SE}(n)$.
-For grid-like data, foundational works include Group-Equivariant CNNs (Cohen & Welling, 2016) and Steerable CNNs (Weiler et al., 2018; Weiler & Cesa, 2021).
-For unstructured data, a dominant paradigm uses features associated with irreducible group representations (tensorial methods, E(n)-GNNs, and transformers (Thomas et al., 2018; Fuchs et al., 2020; Satorras et al., 2021; Batatia et al., 2022; Battiloro et al., 2025; Liao & Smidt, 2023; Assaad et al., 2023; Hutchinson et al., 2021)), with lightweight vector-based alternatives also available (Deng et al., 2021; Son et al., 2024).
-Theoretical analyses have also explored the universal approximation capabilities of such invariant networks (Maron et al., 2019).
-Complementing these specialized layers are general, model-agnostic strategies like frame averaging and canonicalization (Puny et al., 2022; Lin et al., 2024b; Kaba et al., 2023; Panigrahi & Mondal, 2024).
-Despite their success, most practical implementations are centered on scalar and vector features.
-While frameworks such as Tensor Field Networks (Thomas et al., 2018) or EMLP (Finzi et al., 2021) can handle higher-order tensor representations, scalable architectural design for general matrix-valued quantities (e.g., covariances or inertia tensors transforming as $\Sigma\mapsto R\Sigma R^{\top}$) is still limited in practice.
-A major frontier is extending equivariance to non-compact groups like the general linear group $\mathrm{GL}(n)$ and the Lorentz group $\mathrm{SO}^{+}(1,3)$.
-Much of the work on the Lorentz group, which is critical in particle physics, has focused on designing specialized networks for the standard left action on vectors (Bogatskiy et al., 2020; Finzi et al., 2021; Batatia et al., 2023; Zhdanov et al., 2024).
-In contrast, we demonstrate that this left-action problem can be addressed within our universal, adjoint-equivariant framework via an embedding map.
-One line of work generalizes group convolution and kernel design, often leveraging Fourier analysis (Xu et al., 2022; Helwig et al., 2023).
-This includes general frameworks for constructing equivariant networks on arbitrary matrix groups (Basu et al., 2025), through matrix functions (Batatia et al., 2024), and on reductive Lie groups (Batatia et al., 2023).
-Other approaches define kernels in the Lie algebra (Finzi et al., 2020), use Lie group decompositions for integration (Mironenco & Forré, 2024), or adapt canonicalization using infinitesimal generators of the Lie algebra (Shumaylov et al., 2025).
-For these non-compact groups in general, the theoretical machinery underpinning many equivariant models does not readily apply.
-One line of work generalizes group convolution and kernel design, leveraging tools from Fourier analysis (Xu et al., 2022; Helwig et al., 2023), matrix functions (Batatia et al., 2024), or operating directly on reductive Lie groups (Batatia et al., 2023).
-Other methods focus on the Lie algebra, either by defining kernels within the algebra itself (Finzi et al., 2020) or using Lie group decompositions for integration (Mironenco & Forré, 2024; Shumaylov et al., 2025).
-Another approach leverages differential geometry, using tools such as partial differential operators (He et al., 2022; Shen et al., 2020; Jenner & Weiler, 2022) and the algebra of differential invariants (Sangalli et al., 2022; Li et al., 2024) to construct equivariant layers.
-A different generalist approach, taken by methods like the Equivariant MLP and G-RepsNet (Finzi et al., 2021; Basu et al., 2025), is to solve the equivariance constraint algebraically, though this often lacks specialized inductive biases like locality.
-In contrast to these manifold-focused approaches, a parallel line of work operates directly on the Lie algebra.
-Lin et al.
-(2024a) introduced Lie Neurons, establishing a framework for adjoint-equivariance, but their method is restricted to semisimple Lie algebras where the Killing form is non-degenerate.
-This limitation precludes direct application to the reductive but non-semisimple algebra $\mathfrak{gl}(n)$, whose degenerate Killing form poses a central challenge.
-The high computational cost of group convolution methods, combined with the semisimple constraints of existing Lie-algebraic techniques, limits their direct applicability to robotics tasks that involve real-time processing, noisy measurements, or uncertainty estimation (Eschmann et al., 2024; Yu & Sun, 2024).
-While traditional, model-based algorithms like Kalman filters are designed to respect this geometry (Barrau & Bonnabel, 2016; Hartley et al., 2020), they lack the flexibility.
-Our work confronts this challenge with ReLNs, a practical, numerically well-conditioned architecture for the full reductive group $\mathrm{GL}(n)$ and its subgroups.
-By introducing a learnable, non-degenerate, and $\mathrm{Ad}$-invariant bilinear form, our framework overcomes the key obstacle of the degenerate Killing form in $\mathfrak{gl}(n)$.
-This unified algebraic approach extends beyond strict Lie algebra elements to other geometric features, such as covariance matrices, that transform under congruence, providing a practical tool that sidesteps the complexities of prior methods requiring group integration or degenerate invariants.
-Figure 2 situates ReLNs as a general framework for $\mathrm{GL}(n)$ transformations, distinguishing our approach from architectures tailored to specialized subgroups.
-3 Preliminaries
-Our work builds equivariant networks on the Lie algebra $\mathfrak{g}=\mathrm{Lie}(G)$, the tangent space of a Lie group $G$ at the identity $e$.
-We focus on the general linear group $\mathrm{GL}(n)$ and its Lie algebra $\mathfrak{gl}(n)$, aiming for equivariance under the adjoint action $\mathrm{Ad}_{g}:X\mapsto gXg^{-1},\;g\in\mathrm{GL}(n),\;X\in\mathfrak{gl}(n)$.
-A central challenge arises from the structure of $\mathfrak{gl}(n)$: it is a reductive Lie algebra, and the canonical invariant inner product, the Killing form, is degenerate on $\mathfrak{gl}(n)$.
-This degeneracy poses a critical problem, as it prevents the construction of expressive nonlinear layers, causing them to collapse into linear maps and severely limiting the model’s expressive power.
-We address this by introducing a learnable, non-degenerate bilinear form, enabling fully nonlinear equivariant operations.
-For details of Lie theory and background, see Appendix A.
+    We demonstrate practical utility and efficiency across diverse regimes, matching Lie-algebra baselines on $\mathfrak{sl}(3)$ and $\mathfrak{sp}(4)$, achieving strong robustness gains on uncertainty-aware drone state estimation and 3D Gaussian-splat representation learning, substantially reducing per-step compute compared to EMLP on the double-pendulum benchmark, matching baseline performance on jet top-tagging with Lorentz-equivariant embeddings.
+Figure 1: Examples of Lie groups and related manifolds in scientific applications.
+From left: the special linear group $\mathrm{SL}(3)$ (image homography), the Lorentz group $\mathrm{SO}^{+}(1,3)$ (spacetime symmetry), symplectic groups $\mathrm{Sp}(n)$ (Hamiltonian mechanics), the $\mathrm{SPD}(3)\times\mathbb{R}^{3}$ state space (probabilistic estimation), and the general linear group $\mathrm{GL}(3)$ (modeling stress-strain in continuum mechanics).
 Figure 2: A taxonomy of selected representative equivariant neural architectures, categorized by the symmetries to which they are equivariant.
@@
 An asterisk () denotes methods equivariant to the group’s adjoint action.
+2 Related Work
+Encoding symmetry into neural architectures is a well-established inductive bias for improving data efficiency and generalization (Bronstein et al., 2021).
+Most geometric deep learning work targets Euclidean isometries such as rotations and rigid motions (e.g., $\mathrm{SO}(n)$ and $\mathrm{SE}(n)$).
+For grid data, representative approaches include group-equivariant and steerable CNNs (Cohen and Welling, 2016; Weiler et al., 2018; Weiler and Cesa, 2021).
+For sets and graphs, many methods build features from irreducible representations and tensorial message passing, including TFNs, $\mathrm{E}(n)$-GNNs, and equivariant transformers (Thomas et al., 2018; Fuchs et al., 2020; Satorras et al., 2021; Batatia et al., 2022; Liao and Smidt, 2023; Assaad et al., 2023; Hutchinson et al., 2021), alongside lighter vector-based variants (Deng et al., 2021; Son et al., 2024).
+Related theoretical work studies universality of invariant/equivariant architectures (Maron et al., 2019), and model-agnostic strategies such as canonicalization and frame averaging provide complementary routes to equivariance (Puny et al., 2022; Lin et al., 2024b; Kaba et al., 2023; Panigrahi and Mondal, 2024).
+While several frameworks can represent higher-order tensors (e.g., TFN (Thomas et al., 2018) and EMLP (Finzi et al., 2021)), scalable architectural design for general matrix-valued quantities (e.g., covariances transforming as $\Sigma\mapsto R\Sigma R^{\top}$) remains less standardized in practice.
+Non-compact groups and $\mathrm{GL}(n)$.
+Extending equivariance to non-compact symmetries remains an active area, notably exemplified by Lorentz-equivariant architectures that operate on Minkowski geometry in particle physics (Bogatskiy et al., 2020; Gong et al., 2022; Batatia et al., 2023; Zhdanov et al., 2024).
+Complementary research utilizes geometric (Clifford) algebras to represent features as multivectors, leading to frameworks such as Geometric Algebra Transformers (GATr) Brehmer et al.
+(2023), Clifford-equivariant simplicial message passing Liu et al.
+(2024), and lightweight designs like GLGENN Filimoshina and Shirokov (2025).
+Approaches to non-compact groups include generalizations of group convolutions and kernels (Xu et al., 2022; Helwig et al., 2023), constructions based on matrix functions or reductive groups (Batatia et al., 2024; 2023), Lie-algebra-based kernels and decompositions (Finzi et al., 2020; Mironenco and Forré, 2024; Shumaylov et al., 2025), differential-operator and differential-invariant methods (He et al., 2022; Shen et al., 2020; Jenner and Weiler, 2022; Sangalli et al., 2022; Li et al., 2024), and algebraic constraint-solving frameworks such as EMLP and G-RepsNet (Finzi et al., 2021; Basu et al., 2025).
+These lines provide general tools but can incur substantial overhead (e.g., integration/basis computation) or lack inductive biases such as locality.
+Within Lie-algebraic adjoint-equivariant learning, Lie Neurons (Lin et al., 2024a) establish $\mathrm{Ad}$-equivariant neural network but rely on semisimple algebras where the Killing form is non-degenerate, and thus do not directly cover reductive algebras.
+This gap is relevant when inputs include general matrix-valued quantities that transform by adjoint action.
+Classical invariant filtering explicitly respects these transformation rules (Barrau and Bonnabel, 2016; Hartley et al., 2020), but is model-based and does not provide a learned backbone.
+Learning on 3D Gaussian splats.
+3D Gaussian Splatting represents geometry using anisotropic primitives parameterized by a mean $\mu\in\mathbb{R}^{3}$ and covariance $\Sigma\in\mathrm{SPD}(3)$.
+Recent learning frameworks over such primitives, including self-supervised learning and dynamics/segmentation models (Ma et al., 2025; Lu et al., 2024; Zhang et al., 2024; Ye et al., 2024), often treat heterogeneous attributes as loosely coupled channels (e.g., concatenating $\mu$ with rotation parameters and applying generic architectures), which does not explicitly enforce the coupled transformation of $(\mu,\Sigma)$.
+This decoupled treatment disregards the underlying geometry, representing vectors and tensors as mixed latent feature that fail to recognize their transformation laws.
+Consequently, such models struggle to capture the symmetry of the scene without massive data augmentation.
+ReLN formulation targets this gap by embedding these heterogeneous attributes into a unified Lie-algebraic space, enforcing the coupled equivariant structure of $(\mu,\Sigma)$, exploiting the isomorphism between congruence and adjoint actions under $\mathrm{SO(3)}$.
+ReLNs overcome these limitations by establishing exact adjoint equivariance on $\mathfrak{gl}(n)$ via a non-degenerate $\mathrm{Ad}$-invariant bilinear form.
+This approach bypasses the reliance on degenerate invariants and group-specific integration required by previous methods.
+Figure 2 situates ReLNs relative to prior equivariant frameworks.
+3 Preliminaries
+We build equivariant networks on a reductive Lie algebra $\mathfrak{gl}(n)$, requiring equivariance under the adjoint action $\mathrm{Ad}_{g}(X)=gXg^{-1}$ for $g\in\mathrm{GL}(n)$.
+A key obstruction is that $\mathfrak{gl}(n)$ is reductive but not semisimple: its canonical $\mathrm{Ad}$-invariant symmetric bilinear form (the Killing form) is degenerate.
+Many equivariant architectures rely on a non-degenerate invariant form to produce invariant scalars (e.g., gating/normalization) and to parameterize genuinely nonlinear equivariant operations (e.g., Vector Neurons (Deng et al., 2021), Lie Neurons (Lin et al., 2024a)); degeneracy makes these constructions ill-conditioned or effectively linear.
+We resolve this by introducing a non-degenerate $\mathrm{Ad}$-invariant bilinear form on $\mathfrak{gl}(n)$, enabling fully nonlinear adjoint-equivariant layers.
+Our scope is reductive Lie algebras; extending to non-reductive cases (e.g., $\mathfrak{aff}(n)$) generally requires additional non-canonical completions.
+For details of Lie theory and background, see Appendix A.
 4 Reductive Lie Neurons: Architecture
-We present ReLNs, a framework for building deep networks equivariant to the adjoint action of the general linear group $\mathrm{GL}(n)$.
-The design centers on a learnable, non-degenerate, $\mathrm{Ad}$-invariant bilinear form on the reductive Lie algebra $\mathfrak{gl}(n)$ and a complete toolbox of equivariant linear maps, nonlinearities, pooling, and invariant readouts.
-4.1 A General $\mathrm{Ad}$-invariant Bilinear Form for Reductive Lie Algebras
-The primary obstacle to applying Lie-algebraic methods such as Lie Neurons (Lin et al., 2024a) to $\mathfrak{gl}(n)$ is the degeneracy of its Killing form.
-We resolve this by constructing a modified bilinear form $\widetilde{B}$ that restores non-degeneracy while preserving the crucial $\mathrm{Ad}$-invariance property.
+We present ReLNs, a framework for building deep networks equivariant to the adjoint action of $\mathrm{GL}(n)$ on its Lie algebra $\mathfrak{gl}(n)$.
+The design centers on a non-degenerate, $\mathrm{Ad}$-invariant bilinear form on the reductive Lie algebra $\mathfrak{gl}(n)$ and a complete toolbox of equivariant linear maps, nonlinearities, pooling, and invariant readouts.
+4.1 An $\mathrm{Ad}$-invariant bilinear form for reductive Lie algebras
+A key obstruction to extending Lie-algebraic designs (e.g., Lie Neurons (Lin et al., 2024a)) from semisimple Lie algebras to $\mathfrak{gl}(n)$ is that the Killing form may be degenerate on non-semisimple algebras.
+A finite-dimensional Lie algebra is reductive if it admits an ideal decomposition $\mathfrak{g}=\mathfrak{z}(\mathfrak{g})\oplus[\mathfrak{g},\mathfrak{g}]$ with $\mathfrak{g},\mathfrak{g}]$ semisimple, and the Killing form is $B_{\mathfrak{g}}(X,Y)=\operatorname{Tr}(\mathrm{ad}_{X}\circ\mathrm{ad}_{Y})$; see Appendix A.4 for detailed definitions and the Killing form on $\mathfrak{g}$.
+In particular, $\mathfrak{gl}(n)$ is reductive but not semisimple: it contains a nontrivial center on which $\mathrm{ad}$ vanishes, forcing degeneracy of the Killing form.
+To obtain well-posed invariant contractions and nonlinearities, we construct an $\mathrm{Ad}$-invariant non-degenerate completion of the Killing form on any reductive Lie algebra.
 Definition 4.1 (Modified Bilinear Form on a Reductive Lie Algebra).
@@
 $\widetilde{B}(Z_{1}{+}X_{1},\,Z_{2}{+}X_{2}):=\langle Z_{1},Z_{2}\rangle_{\mathfrak{z}}+B(X_{1},X_{2}),\qquad$ (1)
-where $B$ denotes the Killing form on $\mathfrak{g},\mathfrak{g}]$.
-Proposition 4.1.
-The bilinear form $\widetilde{B}$ is symmetric, $\mathrm{Ad}$-invariant, and nondegenerate.
+where $B$ denotes the Killing form on the semisimple ideal $\mathfrak{g},\mathfrak{g}]$.
+Proposition 4.2 (Non-degeneracy and $\mathrm{Ad}$-invariance).
+The form $\widetilde{B}$ is symmetric, $\mathrm{Ad}$-invariant, and non-degenerate.
 Moreover, $\mathfrak{z}(\mathfrak{g})$ and $\mathfrak{g},\mathfrak{g}]$ are $\widetilde{B}-$orthogonal, with $\widetilde{B}|_{[\mathfrak{g},\mathfrak{g}]}=B$ and $\widetilde{B}|_{\mathfrak{z}(\mathfrak{g})}=\langle\cdot,\cdot\rangle_{\mathfrak{z}}$.
@@
 Nondegeneracy follows since $B$ is nondegenerate on the semisimple ideal and $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ is nondegenerate on the center; orthogonality holds because $B(\mathfrak{z}(\mathfrak{g}),[\mathfrak{g},\mathfrak{g}])=0$.
+Detailed proofs are provided in Appendix C.
 ∎
-For our primary case $\mathfrak{g}=\mathfrak{gl}(n,\mathbb{R})=\mathbb{R}I\oplus\mathfrak{sl}(n,\mathbb{R})$ we choose the canonical trace-based form
-$\widetilde{B}(X,Y)\;=\;2n\cdot\mathrm{tr}(XY)\;-\;\mathrm{tr}(X)\,\mathrm{tr}(Y),$ (2)
-This form $\widetilde{B}$ is the fundamental tool used throughout the ReLN architecture.
+For our primary case $\mathfrak{g}=\mathfrak{gl}(n)=\mathbb{R}I\oplus\mathfrak{sl}(n)$, we use the closed form
+$\widetilde{B}(X,Y)\;=\;2n\,\operatorname{Tr}(XY)\;-\;\operatorname{Tr}(X)\operatorname{Tr}(Y).$ (2)
+This $\widetilde{B}$ is our basic $\mathrm{Ad}$-invariant contraction throughout ReLN architecture.
 Verification and Relation to Prior Bilinear Forms.
 Our concrete form in Eq.
-2 satisfies the conditions of Proposition 4.1.
+2 satisfies the conditions of Definition 4.1.
 Decomposing a matrix $X=X_{0}+\frac{1}{n}\mathrm{tr}(X)I$ (where $X_{0}\in\mathfrak{sl}(n)$) reveals that our form separates orthogonally:
@@
 The second term is a standard inner product on the center, which, under the isomorphism $\mathfrak{so}(3)\simeq\mathbb{R}^{3}$, recovers the dot product used in Vector Neurons (Deng et al., 2021).
-Our single form thus unifies these approaches, extending to the full reductive algebra $\mathfrak{gl}(n)$ (details can be found in Appendix C).
+For $\mathfrak{so}(3)$, the hat map intertwines the vector action and the adjoint action, so $\mathrm{Ad}$-equivariance on $\mathfrak{so}(3)$ corresponds to standard $\mathrm{SO}(3)$-equivariance on $\mathbb{R}^{3}$; see Lemma B.7.
+Our single form thus unifies these approaches, extending to the full reductive algebra $\mathfrak{gl}(n)$ (details can be found in Appendix B).
 Figure 3: Adjoint equivariance using a unified representation for diverse geometric inputs.
@@
 ReLN-Linear.
-A linear map applied to the channel dimension $f(x;W)=xW$ with $W\in\mathbb{R}^{C\times C^{\prime}}$ is strictly equivariant: the group acts on the left (geometric dimension) while $W$ acts on the right (channel dimension), and thus these operations commute (formal proof in Appendix D).
+A linear map applied to the channel dimension $f(x;W)=xW$ with $W\in\mathbb{R}^{C\times C^{\prime}}$ is equivariant: the group acts on the left (geometric dimension) while $W$ acts on the right (channel dimension), and thus these operations commute (formal proof in Appendix E).
 Equivariant Nonlinearities.
-We introduce two complementary nonlinear primitives; full definitions, parameterizations, and stability prescriptions are deferred to Appendix D.
-  • ReLN-ReLU: To overcome the non-equivariance of elementwise activations, we use our form $\widetilde{B}$ to define a directional nonlinearity.
-Each vector feature $x_{c}$ is rectified along a learnable direction $d_{c}$ via update rule:
-  • ReLN-ReLU Using $\widetilde{B}$ to build an invariant gate, each channel feature is updated by
-    $x_{c}^{\prime}\;=\;x_{c}\;+\;\max\big(0,\;\widetilde{B}(x_{c}^{\wedge},d_{c}^{\wedge})\big)\,d_{c},$ (4)
+We introduce two complementary nonlinear primitives; full definitions, parameterizations, and stability prescriptions are deferred to Appendix E.
+  • ReLN-ReLU: To avoid basis-dependent elementwise activations, we use $\widetilde{B}$ to build an invariant scalar gate and apply it along an equivariant direction.
+For input features $x\in\mathbb{R}^{K\times C}$, we explicitly define the equivariant reference direction $d=xW$, where $W\in\mathbb{R}^{C\times C}$ is a learnable weight matrix.
+The feature is then updated via the following gated mechanism:
+    $x_{c}^{\prime}=\begin{cases}x_{c},&\text{if }\widetilde{B}(x_{c}^{\wedge},d_{c}^{\wedge})\leq 0\\ x_{c}+\widetilde{B}(x_{c}^{\wedge},d_{c}^{\wedge})d_{c},&\text{otherwise}\end{cases}$ (4)
 […]
@@
 The final components of the ReLN toolbox enable feature aggregation and the production of invariant outputs.
-  • Max-Killing Pooling: To aggregate a set of features $\{X_{n}\}_{n=1}^{N}$, where each $X_{n}$ is a multi-channel feature tensor, this layer selects the representative feature with the maximal projection onto a learnable direction.
-For each channel $c$, the index is found via $n^{*}(c)=\arg\max_{n}\widetilde{B}(X_{n,c},D_{n,c})$, and the pooled feature is $X^{\max}_{c}=X_{n^{*}(c),c}$.
+[…]
   • Invariant Layer: To produce a group-invariant output, this layer contracts feature $X_{c}$ using the form $\widetilde{B}$.
@@
 5 Experiments
-We evaluate ReLNs on a suite of tasks designed to highlight two complementary strengths: algebraic generality on benchmarks and practical efficacy on a challenging, uncertainty-aware robotics task.
-We compare against standard non-equivariant baselines (MLP, ResNet), the original Lie Neurons, and a Vector Neurons–style baseline adapted for covariance inputs by eigendecomposition.
-For the Top-Tagging benchmark we also report results versus established physics models used in prior work.
+We evaluate ReLNs along three axes: (i) exact adjoint equivariance under broad linear symmetries (including reductive and non-compact groups), (ii) learning from coupled heterogeneous geometry (e.g., velocities with covariances), and (iii) accuracy–efficiency trade-offs against widely used equivariant and self-supervised baselines.
+We consider (a) algebraic benchmarks for adjoint-equivariant learning on $\mathfrak{sl}(3)$ and $\mathfrak{sp}(4)$, (b) uncertainty-aware drone state estimation with joint velocity–covariance processing under $\mathrm{SO}(3)$ frame changes, (c) the EMLP double-pendulum benchmark assessing accuracy and per-step efficiency, and (d) learning from 3D Gaussian-splat primitives in ShapeSplat Ma et al.
+(2025) with comparisons to Gaussian-MAE.
+We compare against non-equivariant baselines (MLP, ResNet), Lie-algebraic equivariant models (Lie Neurons), and strong equivariant architectures including Vector Neurons (VN), Tensor Field Networks (TFN), and an $\mathrm{SE}(3)$-Transformer.
+Additional Lorentz-equivariant top-tagging results are deferred to Appendix H.
 5.1 Algebraic Benchmarks on Semisimple Lie Algebras
@@
 5.1.1 Platonic Solid Classification
-We first validate our model on the Platonic solid classification benchmark from (Lin et al., 2024a), testing the adjoint-equivariance where camera rotations induce a conjugation action on inter-face homographies in $\mathrm{SL}(3)$.
-Full experimental details are provided in Appendix E.2.
+We first validate our model on the Platonic solid classification benchmark from (Lin et al., 2024a), testing the adjoint-equivariance where camera rotations induce a conjugation action on the homographies between the projected faces of the solid, represented as $\mathrm{SL}(3)$.
+Full experimental details are provided in Appendix G.2.
 The results, summarized in Table 1, confirm that non-equivariant baselines fail to generalize to rotated camera views.
-This fundamental weakness is not resolved by data augmentation or increased model capacity, as our wider MLP variant with approximately double the parameters shows negligible improvement on the out-of-distribution test set.
+This generalization gap persists even with augmentation or increased capacity, as our wider MLP variant with approximately double the parameters shows negligible improvement on the out-of-distribution test set.
 In contrast, the ReLN model achieves near-perfect accuracy with robustness, matching the performance of the Lie Neurons while demonstrating improved results.
-Importantly, this result validates that our general $\mathfrak{gl}(n)$ framework operates effectively on common semisimple subalgebras, as its built-in adjoint-equivariance on the parent group yields robust behavior when restricted to subgroups like $\mathrm{SO}(3)$ and $\mathrm{SL}(3)$.
+This result validates that our general $\mathfrak{gl}(n)$ framework operates effectively on common semisimple subalgebras, as its built-in adjoint-equivariance on the parent group yields robust behavior when restricted to subgroups like $\mathrm{SO}(3)$ and $\mathrm{SL}(3)$.
 5.1.2 Invariant Function Regression on $\mathfrak{sp}(4)$.
-To further probe our framework’s algebraic generality, our second benchmark involves regressing a highly nonlinear invariant function on the real symplectic Lie algebra $\mathfrak{sp}(4,\mathbb{R})$.
-The symplectic algebra $\mathfrak{sp}(2n)$ is the mathematical foundation of Hamiltonian mechanics, which describes any physical system where energy is conserved.
-Our target is a scalar invariant defined for pairs $X,Y\in\mathfrak{sp}(4,\mathbb{R})$:
-$g(X,Y)=\sin\!\big(\operatorname{Tr}(XY)\big)+\cos\!\big(\operatorname{Tr}(YY)\big)-\tfrac{1}{2}\operatorname{Tr}(YY)^{3}+\det(XY)+\exp\!\big(\operatorname{Tr}(XX)\big).$ (6)
-We generate a dataset of 10k training and 10k test pairs by sampling from $\mathfrak{sp}(4,\mathbb{R})$.
-We compare ReLN against MLP baselines (trained with and without $\mathrm{Sp}(4)$ data augmentation) and the original Lie Neurons.
-At test time, we report the standard MSE, MSE averaged over 500 random adjoint actions, and the invariance error.
-As shown in Table 3, non-equivariant MLPs are orders of magnitude less accurate and exhibit high invariance error, failing to learn the group structure even with data augmentation.
-Our ReLN model not only achieves the lowest MSE and near-zero invariance error, but also shows a modest but consistent improvement over Lie Neurons.
-This suggests that our non-degenerate bilinear form provides not only theoretical generality but also superior numerical conditioning in practice.
-Table 3: Regression performance and invariance error on $\mathfrak{sp}(4)$. “Training Aug.” indicates whether $\mathrm{Sp}(4)$ conjugation was applied during training.
-Model Training Aug. # Params Test Aug. Invariance Error
+To probe algebraic generality beyond $\mathfrak{sl}(n)$, we regress a highly nonlinear $\mathrm{Sp}(4)$-invariant scalar on the real symplectic Lie algebra $\mathfrak{sp}(4,\mathbb{R})$ (the Lie algebra underlying Hamiltonian/symplectic symmetry).
+For $X,Y\in\mathfrak{sp}(4,\mathbb{R})$, the target is
+$g(X,Y)=$ $\sin\!\big(\operatorname{Tr}(XY)\big)+\cos\!\big(\operatorname{Tr}(YY)\big)$
+$-\tfrac{1}{2}\operatorname{Tr}(YY)^{3}+\det(XY)+\exp\!\big(\operatorname{Tr}(XX)\big).$ (6)
+We sample 10k training and 10k test pairs and compare ReLN to MLP baselines (with/without $\mathrm{Sp}(4)$ conjugation augmentation) and Lie Neurons.
+We report test MSE, MSE averaged over 500 random adjoint actions, and the invariance error.
+As shown in Table 2, non-equivariant MLPs are orders of magnitude less accurate and exhibit substantially larger invariance error, indicating that they fail to reliably capture the underlying group structure even with $\mathrm{Sp}(4)$ conjugation augmentation.
+In contrast, Lie Neurons and ReLN achieve low test MSE and near-zero invariance error, demonstrating stable invariant regression on $\mathfrak{sp}(4)$.
+At this parameter scale, the performance of the two equivariant models is comparable.
+These results support that our model matches the accuracy of specialized Lie-algebraic baselines while retaining a unified construction across diverse Lie algebras.
+Table 2: Regression performance and invariance error on $\mathfrak{sp}(4)$. “Tr. Aug.” indicates training augmentation.
+Model Tr. Aug. # Params Test Aug. Inv. Error
 ID SP(4)
@@
 ReLN (ours) Id 263,170 $5.14\times 10^{-4}$} $5.14\times 10^{-4}$} $4.73\times 10^{-7}$
-5.2 Particle Physics with Lorentz Group $\mathrm{SO}(1,3)$ Equivariance
-We test our framework on the Top-Tagging benchmark (Kasieczka et al., 2019), a task to distinguish particle jets originating from top quarks against a large background from Quantum Chromodynamics.
-Because these relativistic collisions are subject to the symmetries of spacetime, the task requires equivariance under the Lorentz group $\mathrm{SO}(1,3)$.
-We solve this left-action equivariant learning problem by introducing a map that embeds the four-momentum $p\in\mathbb{R}^{1,3}$ of each constituent particle, which combines its energy and 3D momentum, into the Lie algebra $\mathfrak{gl}(5)$ as $\varphi(p)=\begin{pmatrix}0_{4\times 4}&p\\ p^{\top}\eta&0\end{pmatrix}$, where $\eta=\mathrm{diag}(-1,1,1,1)$ is the Minkowski metric.
-As proven in Appendix F.1, this embedding unifies left- and adjoint-action equivariance within a single Lie-algebraic framework.
-We adapt the LorentzNet architecture by replacing its invariant feature computation with our proposed bilinear form.
-To ensure a fair comparison, we created a parameter-matched version of LorentzNet.
-As shown in Table 4, our model achieves competitive overall performance while demonstrating an advantage on the background rejection metric.
-This result shows that our general Lie-algebraic approach can effectively unify adjoint- and left-action equivariance in a parameter-efficient manner.
-Table 4: Comparison of performance on the Top-Tagging dataset.
-Rej@30% denotes the background rejection at 30% signal efficiency (higher is better).
-Benchmark scores are as reported in the original publications.
-Architecture #Params Accuracy AUC Rej@30% Reference
-PELICAN 45k 0.943 0.987 $2289\pm 204$ Bogatskiy et al. (2022)
-LorentzNet (original) 224k 0.942 0.987 $2195\pm 173$ Gong et al. (2022)
-LorentzNet (param-matched) 84k 0.942 0.987 $1821\pm 94$ Our reproduction
-LGN 4.5k 0.929 0.964 $435\pm 95$ Bogatskiy et al. (2020)
-BIP 4k 0.931 0.981 $853\pm 68$ Munoz et al. (2022)
-partT 2.14M 0.940 0.986 $1602\pm 81$ Qu et al. (2022)
-ParticleNet 498k 0.938 0.985 $1298\pm 46$ Qu & Gouskos (2020)
-EFN 82k 0.927 0.979 $633\pm 31$ Komiske et al. (2019)
-TopoDNN 59k 0.916 0.972 $295\pm 5$ Pearkes et al. (2017)
-LorentzMACE 228k 0.942 0.987 $1935\pm 85$ Batatia et al. (2023)
-ReLN (Ours) 84k 0.942 0.987 $1979\pm 87$
 5.2 Drone State Estimation with Geometric Uncertainty
-We test our framework on a challenging drone state estimation task using a large-scale dataset of aggressive, highly dynamic flights.
-The objective is to recover a 3D trajectory from a sequence of noisy velocity measurements and their corresponding time-varying covariances, where each covariance matrix quantifies the uncertainty of its associated velocity measurement.
-This setup tests a model’s ability to jointly process vector ($\mathbf{v}$) and matrix ($C$) data in a geometrically consistent and uncertainty-aware manner.
+We evaluate ReLN on a drone state-estimation task that reconstructs 3D trajectories from noisy velocity measurements $\mathbf{v}\in\mathbb{R}^{3}$ and time-varying uncertainty covariances $C\in\mathrm{SPD}(3)$ during highly dynamic flights.
+This setup requires a model to jointly process vector and matrix data in a geometrically consistent and uncertainty-aware manner.
 Experimental Setup.
-We created a large-scale synthetic dataset of 200 aggressive drone trajectories, over 13 hours of challenging, high-speed flight (details in Appendix G).
-The network is trained to regress the 3D position from a sequence of velocity and covariance measurements within a time window.
-Our model is compared against non-equivariant ResNets and an $\mathrm{SO}(3)$-equivariant baseline using Vector Neurons (VN).
-Since VNs cannot directly process matrix inputs, we implement a method that handles covariance matrices via an eigendecomposition-based strategy (details in Appendix G.3).
-In contrast, our ReLN architecture treats both velocity and covariance as unified geometric objects within a single algebraic space, $\mathfrak{gl}(3)$.
-Specifically, the velocity vector $\mathbf{v}\in\mathbb{R}^{3}$ is lifted to its matrix representation in $\mathfrak{so}(3)\subset\mathfrak{gl}(3)$.
-For the time-varying covariance $C$, we explore two representations that also reside in $\mathfrak{gl}(3)$: (1) the matrix $C$ directly, and (2) its matrix logarithm $\log C$, which respects the geometry of the $\mathrm{SPD}(3)$ manifold.
-This unified representation ensures a measurement and its uncertainty transform consistently under the adjoint action of $\mathrm{SO}(3)$.
-The final velocity estimate is then equivariantly extracted by projecting the network’s matrix output onto its skew-symmetric component.
-We test three ReLN variants: (1) velocity only, (2) velocity + covariance, and (3) velocity + log-covariance.
-Full implementation details are provided in Appendix G.
+We use a large-scale synthetic dataset of 200 aggressive trajectories (over 13 hours of flight); details are in Appendix I.
+We evaluate ReLN against: non-equivariant 1D ResNets, VN (Deng et al., 2021), and two spherical-harmonics-based equivariant architectures that use steerable bases and tensor-product coupling to construct equivariant messages: TFN (Thomas et al., 2018) and SE(3)-Transformers (Fuchs et al., 2020).
+To isolate the effect of uncertainty fusion, we test three variants for each equivariant model: velocity-only, velocity + covariance, and velocity + log-covariance.
+For baselines such as VN that cannot natively process matrix inputs, we implement an eigendecomposition-based strategy to ingest covariance data (details in Appendix I.3).
+To ingest covariance data in TFN, SE(3)-Transformer, we apply irreducible representation (irrep) decomposition.
+Irrep decomposition vs. unified adjoint processing.
+Steerable baselines (TFN, SE(3)-Transformer) represent $C$ through irreducible components (trace and traceless-symmetric parts, corresponding to $\ell=0$ and $\ell=2$), and learn interactions with $\mathbf{v}$ via tensor-product coupling between fibers.
+This decomposition ignores the holistic geometry of the measurement-uncertainty pair, relying on tensor-product coupling to reconnect these components.
+See Appendix I.3 for the detailed architecture design and decomposition.
+In contrast, ReLN embeds both $\mathbf{v}$ and $C$ into a common matrix-valued representation with a single conjugation rule: $\mathbf{v}$ is lifted to $\mathfrak{so}(3)\subset\mathfrak{gl}(3)$ and $C$ (or $\log C$) is provided as a matrix feature.
+This yields a uniform $\mathrm{Ad}$-action for the pair and allows gating/normalization using the same $\mathrm{Ad}$-invariant non-degenerate form $\widetilde{B}$.
+The final velocity estimate is obtained by projecting the network output to its skew-symmetric component; implementation details are deferred to Appendix I.
 5.2.1 Results and Analysis
-Table 5: Performance on the drone trajectory dataset. Best result in each column is shown in bold.
-Model ID $\mathrm{SO}(3)$
-ATE ATE% RPE ATE ATE% RPE
-Non-Equivariant Baselines
-   ResNet (Velocity only) $208.07$ $95.06$ $107.60$ $217.02$ $100.39$ $111.29$
-   ResNet (Velocity + Covariance) $205.11$ $94.94$ $106.07$ $213.26$ $98.90$ $109.37$
-Equivariant Baselines
-   VN (Velocity only) $17.36$ $7.52$ $13.51$ $17.36$ $7.52$ $13.51$
-   VN (Velocity + Covariance) $191.78$ $88.66$ $98.39$ $190.22$ $88.47$ $98.26$
+Table 3: Drone state estimation performance (ATE and RTE in meters).
+$v,C,\log C$ denote velocity, covariance, and log-covariance inputs.
+Refer to Appendix I.6 for the exhaustive performance report including all variants and metrics.
+Model Input ID SO(3)
+ATE $\downarrow$ RTE $\downarrow$ ATE $\downarrow$ RTE $\downarrow$
+Non-Equiv.
+ResNet $(v,C)$ $205.11$ $106.07$ $213.26$ $109.37$
+Equiv. Baselines
+VN (Deng et al., 2021) $v$ $17.36$ $13.51$ $17.36$ $13.51$
+VN (Deng et al., 2021) $(v,C)$ $191.78$ $98.39$ $190.22$ $98.26$
+TFN (Thomas et al., 2018) $(v,C)$ $17.56$ $14.40$ $17.56$ $14.40$
+SE(3)-Tr. (Fuchs et al., 2020) $(v,C)$ $21.67$ $16.77$ $21.67$ $16.77$
+Lie Neurons (Lin et al., 2024a) $(v,C)$ $16.86$ $13.65$ $16.86$ $13.65$
 Our Equivariant Models
-   ReLN (Velocity only) $16.85$ $7.31$ $12.7$ $16.85$ $7.31$ $12.7$
-   ReLN (Velocity + Covariance) $16.49$ $7.21$ $13.02$ $16.49$ $7.21$ $13.02$
-   ReLN (Velocity + $\log$-Covariance) 13.92 5.99 11.04 13.92 5.99 11.04
-Abbreviations and units: ID = in-distribution test (no rotation), $\mathrm{SO}(3)$ = test-time random $\mathrm{SO}(3)$ rotations.
-ATE = absolute trajectory error (meters), ATE% = ATE relative to trajectory length (%), RPE = relative pose error (meters).
-Figure 4: Qualitative comparison of the best-case (middle), average (left), and challenging (right) test sequences.
+ReLN (Ours) $v$ $16.85$ $12.70$ $16.85$ $12.70$
+ReLN (Ours) $(v,C)$ $16.49$ $13.02$ $16.49$ $13.02$
+ReLN (Ours) $(v,\log C)$ 13.92 11.04 13.92 11.04
+ReLN (Ours) (no semisimple) $(v,\log C)$ $16.27$ $12.65$ $16.27$ $12.65$
+Figure 4: Trajectory reconstruction across flight difficulties: high (left), best-case (middle), and medium (right).
 ReLN models consistently track the ground truth (black) with high fidelity, especially when leveraging covariance.
-Insets provide a magnified view of the two best-performing variants (ours) to highlight their accuracy.
-Our experiments, summarized in Table 5 and Figure 4, reveal a clear performance hierarchy where the geometric representation of features is the critical factor.
-Non-equivariant models fail to generalize. Non-equivariant models like ResNets fail to generalize to rotated trajectories, and their performance does not improve when provided with covariance data.
-Equivariance is necessary, but implementation design critically affects performance.
-Both VNs and ReLNs establish strong velocity-only equivariant baselines.
-ReLNs show improved performance due to its richer feature learning from both the Lie bracket and our bilinear nonlinearity form.
-However, a method for incorporating covariance is crucial.
-The VN baseline, which decomposes covariance, degrades the performance.
-This result confirms that separating the principle axes and its corresponding scales (eigenvalues) prevents the network from learning the structure between a measurement and its uncertainty.
-Joint processing of velocity and covariance further improves performance.
-Integrating covariance as a geometric object in our framework further reduces the ATE, showcasing the performance gains from unified geometric representation on symmetry-preserving uncertainty handling.
-Log-covariance ablation highlights mathematical and practical contributions.
-Finally, the ReLN (Velocity + $\log$-Covariance) model achieves the best performance.
-By processing covariance matrices via the matrix logarithm, we demonstrate the power of extending Lie-algebraic architectures to operate on manifold-valued data, such as the $\mathrm{SPD}(3)$ matrices representing geometric uncertainty.
-This success establishes ReLNs as a practical and high-performance framework for uncertainty-aware learning in dynamical systems.
-Discussion Our results indicate that ReLNs operate effectively as a geometry- and uncertainty-aware estimator that generalizes across random measurement-frame changes (i.e., arbitrary 3D rotations).
-While the network is not a classical recursive Markovian filter, it learns to integrate velocity measurements with their associated covariances using uncertainty-dependent weighting, producing robust and accurate trajectory estimates.
-This behavior suggests ReLNs are suitable as modular components in downstream systems that require handling of matrix-valued uncertainty.
-Future work will investigate theoretical guarantees for the learned weighting, extensions to other geometric matrix representations, and scalability to higher-dimensional structured geometric inputs.
+Insets provide a magnified view of the two best-performing variants (ReLN) to highlight their accuracy.
+Table 3 summarizes the main findings (full results in Appendix I.6).
+Equivariance is necessary for robustness and accuracy.
+Non-equivariant ResNets exhibit large trajectory errors and do not reliably benefit from adding covariance, indicating that simply concatenating matrix-valued uncertainty does not reliably yield geometry-consistent fusion.
+Moreover, their errors remain high under test-time $\mathrm{SO}(3)$ rotations, confirming poor robustness to measurement-frame changes.
+The Necessity of Structural Compatibility in Covariance Fusion.
+Among equivariant baselines, velocity-only models already attain low error (e.g., VN), but performance becomes highly sensitive once covariance is introduced.
+In particular, the VN $(v,C)$ variant collapses toward non-equivariant performance (Table 3; Table 13), consistent with the fact that an eigendecomposition-based interface can separate coupled degrees of freedom (eigenvectors vs.
+eigenvalues) and yield an unstable or non-geometric uncertainty signal for learning.
+By contrast, TFN and SE(3)-Transformer can incorporate $(v,C)$ through steerable tensor representations, leading to moderate improvements.
+Local equivariant operators offer superior robustness for uncertainty fusion.
+While SE(3)-Transformer shows a slight performance advantage in the velocity-only regime ($v$), the introduction of covariance uncertainty shifts the advantage to TFN.
+As shown in Table 13, TFN consistently improves over SE(3)-Transformer across all covariance-integrated variants (e.g., $(v,\log C)$).
+This pattern suggests that while attention mechanisms can effectively model global dependencies in simple velocity sequences, local convolutional aggregation provides a stronger and more stable inductive bias for processing high-frequency dynamics and uncertainty.
+ReLN yields the most consistent uncertainty-aware gains.
+ReLN achieves the best overall performance, and its strongest configuration uses $(v,\log C)$ (ATE $13.92$, RTE $11.04$).
+Unlike tensorial pipelines, ReLN represents velocity and covariance in a shared matrix-algebraic space with a single conjugation rule, enabling uncertainty-conditioned contractions and gating without violating equivariance.
+The improvement from $C$ to $\log C$ further supports that respecting the intrinsic geometry of $\mathrm{SPD}(3)$ via the matrix logarithm improves conditioning for learning.
+Ablative Validation of Reductive Decomposition.
+ReLN improves over its semisimple-only variant (the no-center model, equivalent to Lie Neurons), indicating that incorporating the central component $\mathfrak{z}(\mathfrak{g})$ is beneficial for uncertainty-aware estimation.
+Conversely, removing the semisimple ideal (retaining only the center) degrades performance, increasing ATE from $13.92$ to $16.27$ (Table 3).
+Together, these ablations support that the best performance arises from modeling the full reductive structure.
+Discussion.
+Overall, the results indicate that ReLNs act as a geometry- and uncertainty-aware estimator, remaining stable under random measurement-frame changes (test-time $\mathrm{SO}(3)$ rotations).
+While ReLN is not a classical recursive (Markovian) filter, it learns to fuse velocity sequences with their time-varying covariances through uncertainty-dependent modulation, improving trajectory reconstruction.
+This suggests ReLNs as a modular backbone for estimation pipelines that must handle matrix-valued uncertainty.
+A natural direction is to characterize when the learned fusion approximates uncertainty-adaptive integration, and to integrate ReLNs with classical state-estimation frameworks.
+5.3 Equivariance for 3D Gaussian Splatting (3DGS)
+Figure 5: The ReLN-integrated Gaussian-MAE Framework.
+Our architecture pipelines raw 3D Gaussian splats through a Geometric Lifting stage to map them into the $\mathfrak{gl}(3)$ Lie algebra.
+The model explicitly bifurcates the data flow: (1) Active Geometry (position $\mu$, covariance $\Sigma$) is processed via the ReLN Encoder and Decoder to maintain $\mathrm{GL}(3)$-equivariance, while (2) Invariant Attributes (color $c$, opacity $\alpha$) are integrated as Type-0 features.
+A dedicated Central Positional Embedding ensures stable spatial context.
+Finally, the Geometric Projector utilizes the Vee map and Killing form to reconstruct the physical Gaussian parameters while preserving holistic symmetry.
+Standard point cloud networks typically process only spatial coordinates.
+However, 3DGS represents scenes using anisotropic 3D Gaussians parameterized by a mean position $\mu\in\mathbb{R}^{3}$ and a covariance matrix $\Sigma\in\mathrm{SPD}(3)$.
+This introduces a geometric challenge: a global rotation $R\in\mathrm{SO}(3)$ acts differently on these components, where $\mu$ transforms as a vector ($R\mu$) and $\Sigma$ transforms via congruence ($R\Sigma R^{\top}$).
+To evaluate ReLNs in this multi-modal geometric setting, we adapt the masked autoencoder framework proposed by Ma et al.
+(2025), which establishes a self-supervised pretraining baseline for 3D Gaussian Splats.
+While their original architecture processes Gaussian parameters as loosely coupled features using standard Transformers, we redesign the encoder and decoder using ReLN blocks to enforce $\mathrm{GL}(3)$-equivariance.
+The overall architecture of our ReLN-integrated Gaussian-MAE is illustrated in Figure 5.
+Specifically, we unify these geometric types by embedding them into the Lie algebra $\mathfrak{gl}(3)$: the mean position $\mu$ is treated as a translation, while the covariance $\Sigma$ is mapped to the linear subspace $\mathrm{Sym}(3)\subset\mathfrak{gl}(3)$ via the matrix logarithm.
+This ensures that the network learns consistent geometric representations of 3D shapes under arbitrary orientations.
+Experimental Setup.
+We utilize the ShapeNet dataset processed into 3D Gaussian primitives for self-supervised pre-training, following the protocol of Ma et al.
+(2025).
+The model is then fine-tuned on ModelNet10 for classification.
+We compare our ReLN-based architecture against the baseline model from Ma et al.
+(2025).
+Refer to Appendix K for full experimental details.
+Experimental Results.
+Table 4 evaluates classification accuracy on aligned and randomly rotated ($0^{\circ}$$180^{\circ}$) test sets.
+The baseline drops from $93.39\%$ to $18.28\%$.
+In contrast, ReLN maintains a stable accuracy of $95.15\%$, representing a marginal $0.33$ point variance compared to its aligned performance.
+Furthermore, ReLN converges in fewer epochs during pretraining and achieves lower reconstruction error across attributes (including rotation and scale); see Appendix Figure 8.
+Together with the rotation-stable accuracy in Table 4, these results support that enforcing the coupled transformation rules of $(\mu,\Sigma)$ via a reductive Lie-algebraic backbone improves robustness to orientation changes.
+Table 4: Rotation robustness on 3D Gaussian splats.
+Classification accuracy (%) on ModelNet10 for the baseline (Ma et al., 2025) and ReLN.
+Standard uses the aligned test set; Rotated applies random rotations ($0^{\circ}$$180^{\circ}$).
+Method Standard Acc. (%) Rotated Acc. (%)
+Baseline (Ma et al., 2025) 93.39 18.28
+ReLN (Ours) 94.82 95.15
+5.4 EMLP Double-Pendulum Benchmark and Efficiency
+We evaluate ReLN on the EMLP double-pendulum benchmark Finzi et al.
+(2021), which learns Hamiltonian dynamics under multiple symmetry groups ($O(2)$, $SO(2)$, and $D_{6}$).
+Following the standard protocol, we report test rollout errors in Table 5; additional details are provided in Appendix L.
+Accuracy. ReLN achieves comparable or lower rollout error than EMLP across all symmetry settings (e.g., $SO(2)$: 0.010 vs. $0.015$), without introducing group-specific architectural modifications.
+Compute and latency.
+With matched hidden size and representation dimension, ReLN-HNN is substantially cheaper per step than EMLP-HNN (Table 6): 142,190142{,}190 vs.
+1,589,9091{,}589{,}909 FLOPs ($11.18\times$) and $2.216$ vs.
+$61.635$ ms ($27.81\times$).
+This gap is consistent with ReLN using closed-form matrix operations from exact $\mathrm{Ad}$-equivariant primitives on $\mathfrak{gl}(n)$, whereas EMLP relies on symmetry-dependent basis construction and projection.
+Overall, these results support ReLN as a reusable equivariant backbone that attains EMLP-level accuracy while substantially reducing per-step computation across different symmetry choices.
+Table 5: Test rollout error on the EMLP Double-Pendulum Benchmark.
+ReLN matches or improves upon EMLP across different symmetry groups ($O(2)$, $SO(2)$, $D_{6}$).
+Values represent the geometric mean of relative errors, with the standard deviation over 3 trials in parentheses; e.g., the notation $0.012(2)$ denotes $0.012\pm 0.002$.
+Metric $O(2)$ $SO(2)$ $D_{6}$ MLP-HNNs
+EMLP ReLN EMLP ReLN EMLP ReLN
+Rollout Error 0.012 (2) 0.011 (2) 0.015 (3) 0.010 (4) 0.013 (2) 0.011 (2) 0.028
+Table 6: Computational cost on the HNN task. ReLN-HNN reduces FLOPs/step and inference latency relative to EMLP-HNN.
+Model # Params FLOPs / step Inference (ms)
+MLP-HNN 34,817 70,400 0.159
+EMLP-HNN 55,569 1,589,909 61.635
+ReLN-HNN (Ours) 69,889 142,190 2.216
+6 Discussion and Limitations.
+ReLNs are exactly $\mathrm{GL}(n)$-adjoint equivariant by construction.
+For covariance-like quantities, however, the physically standard transformation is congruence, $\Sigma\mapsto A\,\Sigma\,A^{\top}$, which agrees with adjoint conjugation $X\mapsto AXA^{-1}$ only for orthogonal changes of frame ($A^{-1}=A^{\top}$).
+Thus, while our unified Lie-algebraic interface is most directly interpretable under rigid rotations, extending the same measurement–uncertainty coupling to general $\mathrm{GL}(n)$ coordinate changes requires additional geometric justification and may call for alternative liftings or invariants.
+A further direction is to extend the framework to groups with translations, such as affine and $\mathrm{SE}(3)$ symmetries, by leveraging their semidirect-product structure and the induced actions on the corresponding Lie algebras.
 7 Conclusion
-This work introduces ReLNs, a unified neural architecture that provides exact equivariance to the adjoint action of the general $n\times n$ matrix algebra $\mathfrak{gl}(n)$ and its subgroups.
-ReLNs enable efficient learning on Lie-algebraic features and structured geometric data, such as covariance matrices.
-Furthermore, our work establishes a unified Lie-algebraic framework that handles both classical left-action symmetries on vectors and native adjoint-actions on matrices within a single architecture.
-ReLNs achieve state-of-the-art results on benchmarks and deliver large gains in a challenging drone state estimation task by integrating uncertainty.
-We’ll apply our equivariant matrix processing capability to a wider array of physical systems, including the dynamics of articulated robots and large-scale sensor fusion, to further expand the boundaries of geometric deep learning.
+We introduced Reductive Lie Neurons (ReLNs), a unified architecture that is exactly equivariant to the adjoint action on $\mathfrak{gl}(n)$ and its reductive matrix subalgebras.
+Using a non-degenerate $\mathrm{Ad}$-invariant contraction, ReLNs provide an efficient Lie-equivariant framework for learning with structured geometric quantities, including matrix-valued uncertainty.
+The same viewpoint links classical left actions on vectors to adjoint actions on matrices via explicit Lie-algebraic liftings, enabling a single backbone to process heterogeneous geometric inputs in a consistent manner.
+Empirically, ReLNs achieve strong performance on algebraic benchmarks and 3D Gaussian-splat representation learning, and yield substantial gains on drone state estimation by incorporating uncertainty equivariantly.
+We expect these capabilities to benefit data-efficient representation learning in vision and robotics, particularly for structured 3D scene models—such as 3D Gaussian splats and world models—where coupled transformation rules are central.
+Future work will scale ReLNs to larger physical systems (e.g., multimodal fusion and articulated dynamics) and further explore efficient equivariant architectures for matrix-parameterized representations.
 Acknowledgments
@@
 Appendix A Lie-theoretic preliminaries
-This appendix provides an overview of key concepts and derivations from Lie group theory relevant to our construction of $\mathrm{GL}(n)$-equivariant neural networks referenced in the main text.
+This appendix provides an overview of key concepts and derivations from Lie group theory relevant to our construction of $\mathrm{GL}(n)$ adjoint-equivariant neural networks.
 A.1 Lie groups, Lie algebras, hat/vee
@@
 The bracket quantifies non-commutativity of generators: $[A,B]=0$ implies commutativity, whereas $A,B]\neq 0$ indicates a non-trivial interaction.
-A.2 Representations and the adjoint
+A.2 Representations and the adjoint action
 A representation $\Phi:G\to\mathrm{GL}(V)$ differentiates to $\phi:\mathfrak{g}\to\mathfrak{gl}(V)$ by
@@
 A.3 Vectorized adjoint
-Using the hat/vee maps, the adjoint action on the Lie algebra induces a corresponding action on the vector coordinates. This vectorized action is a linear map represented by a matrix:
+Using the hat ($\wedge$) / vee ($\vee$) maps, the adjoint action on the Lie algebra induces a corresponding action on the vector coordinates.
+This vectorized action is a linear map represented by a matrix:
 $\mathrm{Ad}^{m}_{g}:\mathbb{R}^{m}\to\mathbb{R}^{m},\qquad\mathrm{Ad}^{m}_{g}(x)=(\mathrm{Ad}_{g}(x^{\wedge}))^{\vee}.$ (13)
+Equivalently, $\mathrm{Ad}^{m}_{g}$ is the $m\times m$ matrix representation of $\mathrm{Ad}_{g}$ in the chosen basis $\{E_{i}\}$.
 In practice we precompute or assemble the $m\times m$ matrix representing $\mathrm{Ad}^{m}_{g}$ (or apply it implicitly) to implement left-multiplicative equivariant layers that act on vector features.
@@
 Definition A.1 (Semisimple and Reductive Lie Algebras).
-A Lie algebra $\mathfrak{g}$ is:
-  • Semisimple if it is a direct sum of simple Lie algebras (i.e., non-abelian and having no nontrivial ideals).
-  • Reductive if it decomposes as $\mathfrak{g}=\mathfrak{s}\oplus\mathfrak{z}$, where $\mathfrak{s}$ is semisimple and $\mathfrak{z}$ is the center (an abelian Lie subalgebra).
+A finite-dimensional Lie algebra $\mathfrak{g}$ is
+  • semisimple if it has no nonzero solvable ideals (equivalently, its radical is zero);
+  • reductive if it decomposes as a direct sum of ideals
+    $\mathfrak{g}\;=\;\mathfrak{z}(\mathfrak{g})\;\oplus\;[\mathfrak{g},\mathfrak{g}],$ (14)
+    where $\mathfrak{z}(\mathfrak{g})$ is the center and $\mathfrak{g},\mathfrak{g}]$ is semisimple.
 Example A.1.
 The Lie algebra $\mathfrak{gl}(n)$ decomposes as:
-$\mathfrak{gl}(n)=\mathfrak{sl}(n)\oplus\mathbb{R}I,$ (14)
-where $\mathfrak{sl}(n)$ (traceless matrices) is semisimple, and $\mathbb{R}I$ (scalar matrices) forms the center.
+$\mathfrak{gl}(n)=\mathbb{R}I\oplus\mathfrak{sl}(n),\qquad\mathbb{R}I=\mathfrak{z}(\mathfrak{gl}(n)),\quad\mathfrak{sl}(n)=[\mathfrak{gl}(n),\mathfrak{gl}(n)].$ (15)
+where $\mathfrak{sl}(n)$ (traceless matrices) is semisimple, and $\mathbb{R}I$ (scalar matrices) forms the center. Thus $\mathfrak{gl}(n)$ is reductive but not semisimple (it has a nontrivial center).
 This decomposition highlights the non-semisimple nature of $\mathfrak{gl}(n)$, which plays a critical role in understanding the degeneracy of certain invariant forms such as the Killing form.
@@
 Our work addresses this issue in the context of $\mathrm{GL}(n)$-equivariant architectures in Lie algebra $\mathfrak{gl}(n)$.
-Theorem A.1 (Cartan criterion; standard).
-The Killing form $B(X,Y)=\operatorname{tr}(\mathrm{ad}_{X}\circ\mathrm{ad}_{Y})$ is non-degenerate iff $\mathfrak{g}$ is semisimple.
-In particular, because $\mathfrak{gl}(n)$ contains the central scalar direction $\mathbb{R}I$ with $\mathrm{ad}_{I}=0$, the Killing form is degenerate on $\mathfrak{gl}(n)$.
-A.5 Invariant bilinear forms; trace form
-Definition A.2.
+The Killing form on $\mathfrak{g}$ is
+$B_{\mathfrak{g}}(X,Y):=\operatorname{tr}(\mathrm{ad}_{X}\circ\mathrm{ad}_{Y}).$ (16)
+Theorem A.2 (Cartan criterion).
+For a finite-dimensional Lie algebra $\mathfrak{g}$, the Killing form $B_{\mathfrak{g}}$ is non-degenerate if and only if $\mathfrak{g}$ is semisimple.
+In particular, if $\mathfrak{g}$ has nontrivial center then $B_{\mathfrak{g}}$ is degenerate, since $\mathrm{ad}_{Z}=0$ for all $Z\in\mathfrak{z}(\mathfrak{g})$.
+A.5 $\mathrm{Ad}$-invariant bilinear forms and the trace form
+Definition A.3 ($\mathrm{Ad}$-invariance).
 A bilinear form $B:\mathfrak{g}\times\mathfrak{g}\to\mathbb{R}$ is $\mathrm{Ad}$-invariant if
-$B(\mathrm{Ad}_{g}X,\mathrm{Ad}_{g}Y)=B(X,Y)\quad\forall g\in G.$ (15)
-On semisimple algebras the Killing form provides such an invariant, non-degenerate form. On $\mathfrak{gl}(n)$ we instead use the trace form:
-$\langle X,Y\rangle_{\mathrm{tr}}=\operatorname{tr}(XY).$ (16)
-Proposition A.1.
-The trace form is $\mathrm{Ad}$-invariant on $\mathfrak{gl}(n)$:
-$\operatorname{tr}((gXg^{-1})(gYg^{-1}))=\operatorname{tr}(gXYg^{-1})=\operatorname{tr}(XY).$ (17)
+$B(\mathrm{Ad}_{g}X,\mathrm{Ad}_{g}Y)=B(X,Y)\qquad\forall g\in G,\;\forall X,Y\in\mathfrak{g}.$ (17)
+On semisimple $\mathfrak{g}$, the Killing form is $\mathrm{Ad}$-invariant and non-degenerate.
+On reductive but non-semisimple algebras (notably $\mathfrak{gl}(n)$), one often uses alternative $\mathrm{Ad}$-invariant forms.
+A basic example on $\mathfrak{gl}(n)$ is the trace pairing
+$\langle X,Y\rangle_{\mathrm{tr}}:=\operatorname{tr}(XY).$ (18)
+Proposition A.4 ($\mathrm{Ad}$-invariance of the trace pairing on $\mathfrak{gl}(n)$).
+For $g\in\mathrm{GL}(n)$ and $X,Y\in\mathfrak{gl}(n)$,
+$\operatorname{tr}\bigl((gXg^{-1})(gYg^{-1})\bigr)=\operatorname{tr}(XY).$ (19)
 Proof.
-$\operatorname{tr}((gXg^{-1})(gYg^{-1}))=\operatorname{tr}(gXYg^{-1})=\operatorname{tr}(XY)$ by cyclicity of trace. ∎
+By cyclicity of trace, $\operatorname{tr}((gXg^{-1})(gYg^{-1}))=\operatorname{tr}(gXYg^{-1})=\operatorname{tr}(XY)$. ∎
 The trace form is non-degenerate as a bilinear form on the vector space $\mathfrak{gl}(n)$ and therefore provides a practical substitute for the Killing form when designing $\mathrm{Ad}$-invariant bilinear layers on $\mathfrak{gl}(n)$.
-A.6 Killing form on $\mathfrak{sl}(n)$
-Restricted to $\mathfrak{sl}(n)$ the Killing form simplifies and is non-degenerate; one frequently uses the proportionality $B(X,Y)\propto\operatorname{tr}(XY)$ on $\mathfrak{sl}(n)$.
-Appendix B Proofs of Key Theorems
-B.1 Proof of Non-degeneracy and Ad-invariance of Modified Killing Form $B_{e}$
+We will use explicit $\mathfrak{gl}(n)$ identities (including the closed form of $B_{\mathfrak{gl}(n)}$, the resulting degeneracy, and our non-degenerate $\mathrm{Ad}$-invariant modification $\widetilde{B}$) in Appendix B.
+Appendix B Connections for $\mathfrak{gl}(n)$ and $\mathfrak{so}(3)\simeq\mathbb{R}^{3}$
+This appendix collects two specialization facts used in the main text: (i) an explicit expression for the (degenerate) Killing form on $\mathfrak{gl}(n)$ and its relation to our modified form $\widetilde{B}$; (ii) the standard intertwining between the $\mathrm{SO}(3)$ vector action and the adjoint action on $\mathfrak{so}(3)$.
+B.1 The classical Killing form on $\mathfrak{gl}(n)$
+Proposition B.1 (Closed form and degeneracy).
+For $X,Y\in\mathfrak{gl}(n)$,
+$B_{\mathfrak{gl}(n)}(X,Y)=2n\,\operatorname{tr}(XY)-2\,\operatorname{tr}(X)\operatorname{tr}(Y).$ (20)
+In particular, $B_{\mathfrak{gl}(n)}$ is degenerate, and $\mathbb{R}I$ lies in its radical.
+Proof sketch.
+The identity equation 20 is standard for $\mathfrak{gl}(n)$ (up to an overall normalization convention).
+Degeneracy follows since $\mathrm{ad}_{\lambda I}=0$ for all $\lambda I\in\mathbb{R}I$, hence $B_{\mathfrak{gl}(n)}(\lambda I,\cdot)=\operatorname{tr}(\mathrm{ad}_{\lambda I}\circ\mathrm{ad}_{\cdot})=0$.
+∎
+Proposition B.2 (Restriction to $\mathfrak{sl}(n)$).
+If $X,Y\in\mathfrak{sl}(n)$, then
+$B_{\mathfrak{gl}(n)}(X,Y)=2n\,\operatorname{tr}(XY),$ (21)
+which coincides with the Killing form on the semisimple ideal $\mathfrak{sl}(n)$ under the same normalization.
+Proof.
+For $X,Y\in\mathfrak{sl}(n)$, $\operatorname{tr}(X)=\operatorname{tr}(Y)=0$ and equation 20 reduces to $2n\,\operatorname{tr}(XY)$. ∎
+B.2 Our modified $\mathrm{Ad}$-invariant form on $\mathfrak{gl}(n)$
+In the main text we use
+$\widetilde{B}(X,Y)=2n\,\operatorname{tr}(XY)-\operatorname{tr}(X)\operatorname{tr}(Y).$ (22)
+Proposition B.3 (Relation to $B_{\mathfrak{gl}(n)}$).
+With $B_{\mathfrak{gl}(n)}$ as in equation 20,
+$\widetilde{B}(X,Y)=B_{\mathfrak{gl}(n)}(X,Y)+\operatorname{tr}(X)\operatorname{tr}(Y).$ (23)
+Proof.
+Subtract equation 20 from equation 22. ∎
+Proposition B.4 (Orthogonal splitting).
+Write $X=X_{0}+\frac{1}{n}\operatorname{tr}(X)I$ and $Y=Y_{0}+\frac{1}{n}\operatorname{tr}(Y)I$ with $X_{0},Y_{0}\in\mathfrak{sl}(n)$. Then
+$\widetilde{B}(X,Y)=2n\,\operatorname{tr}(X_{0}Y_{0})+\operatorname{tr}(X)\operatorname{tr}(Y),$ (24)
+and $\mathbb{R}I$ and $\mathfrak{sl}(n)$ are $\widetilde{B}-$orthogonal.
+Proof.
+Using $\operatorname{tr}(I)=n$ and $\operatorname{tr}(X_{0})=\operatorname{tr}(Y_{0})=0$,
+$\operatorname{tr}(XY)=\operatorname{tr}(X_{0}Y_{0})+\frac{1}{n}\operatorname{tr}(X)\operatorname{tr}(Y).$ (25)
+Substitute into equation 22. Orthogonality follows by setting $X=\lambda I$ and $Y\in\mathfrak{sl}(n)$. ∎
+Proposition B.5 ($\mathrm{Ad}$-invariance).
+For all $g\in\mathrm{GL}(n)$ and $X,Y\in\mathfrak{gl}(n)$,
+$\widetilde{B}(\mathrm{Ad}_{g}X,\mathrm{Ad}_{g}Y)=\widetilde{B}(X,Y).$ (26)
+Proof.
+Use $\mathrm{Ad}_{g}X=gXg^{-1}$, cyclicity $\operatorname{tr}(gXYg^{-1})=\operatorname{tr}(XY)$, and $\operatorname{tr}(gXg^{-1})=\operatorname{tr}(X)$ in equation 22. ∎
+Proposition B.6 ($\widetilde{B}$ as an $\mathrm{Ad}$-invariant non-degenerate completion on the center).
+The classical Killing form $B_{\mathfrak{gl}(n)}$ is degenerate with radical containing the center $\mathfrak{z}(\mathfrak{gl}(n))=\mathbb{R}I$.
+The modified form $\widetilde{B}$ in equation 22 is $\mathrm{Ad}$-invariant and non-degenerate, and it agrees with $B_{\mathfrak{gl}(n)}$ on the semisimple ideal $\mathfrak{sl}(n)$ while supplying a non-degenerate inner product on the center:
+$\widetilde{B}|_{\mathfrak{sl}(n)\times\mathfrak{sl}(n)}=B_{\mathfrak{sl}(n)},\qquad\widetilde{B}(\lambda I,\mu I)=n^{2}\,\lambda\mu\quad(\lambda,\mu\in\mathbb{R}).$ (27)
+In particular, $\widetilde{B}$ resolves the degeneracy of $B_{\mathfrak{gl}(n)}$ precisely along $\mathbb{R}I$ without altering the semisimple part.
+Proof.
+$\mathrm{Ad}$-invariance is Proposition B.5.
+For $\lambda I,\mu I\in\mathbb{R}I$, equation 22 gives $\widetilde{B}(\lambda I,\mu I)=2n\,\operatorname{Tr}(\lambda\mu I)-\operatorname{Tr}(\lambda I)\operatorname{Tr}(\mu I)=2n(\lambda\mu n)-(\lambda n)(\mu n)=n^{2}\lambda\mu$, which is non-degenerate.
+The restriction to $\mathfrak{sl}(n)$ follows since $\operatorname{Tr}(X)=0$ on $\mathfrak{sl}(n)$, so equation 22 reduces to $2n\,\operatorname{Tr}(XY)$, which matches Proposition B.2 under the same normalization.
+∎
+B.3 $\mathfrak{so}(3)\simeq\mathbb{R}^{3}$: vector action vs. adjoint action
+Let $\hat{\cdot}:\mathbb{R}^{3}\to\mathfrak{so}(3)$ be the standard hat map defined by $\hat{v}\,w=v\times w$, with inverse $(\cdot)^{\vee}:\mathfrak{so}(3)\to\mathbb{R}^{3}$.
+Lemma B.7 (Intertwining identity).
+For all $R\in\mathrm{SO}(3)$ and $v\in\mathbb{R}^{3}$,
+$\widehat{Rv}=R\hat{v}R^{\top}=\mathrm{Ad}_{R}(\hat{v}),\qquad\mathrm{Ad}_{R}(X)=RXR^{-1}=RXR^{\top}.$ (28)
+[…]
+Proof.
+For any $w\in\mathbb{R}^{3}$, using $(Ra)\times(Rb)=R(a\times b)$,
+$(R\hat{v}R^{\top})w=R\hat{v}(R^{\top}w)=R\bigl(v\times(R^{\top}w)\bigr)=(Rv)\times w=\widehat{Rv}\,w.$ (29)
+Thus $\widehat{Rv}=R\hat{v}R^{\top}$. The equivariance statement follows by applying $(\cdot)^{\vee}$ and using $\widehat{Ra}=R\hat{a}R^{\top}$. ∎
+Remark B.8 (Trace contraction recovers the Vector Neuron inner product).
+Under the hat identification $\hat{\cdot}:\mathbb{R}^{3}\to\mathfrak{so}(3)$,
+$\hat{v}=\begin{bmatrix}0&-v_{3}&v_{2}\\ v_{3}&0&-v_{1}\\ -v_{2}&v_{1}&0\end{bmatrix},\qquad\operatorname{Tr}(\hat{v}\,\hat{w})=-2\,v^{\top}w\quad(v,w\in\mathbb{R}^{3}).$ (30)
+Hence the $\mathrm{Ad}$-invariant trace pairing on $\mathfrak{so}(3)$, $\langle X,Y\rangle_{\mathrm{tr}}:=-\tfrac{1}{2}\operatorname{Tr}(XY)$, corresponds exactly to the Euclidean dot product on $\mathbb{R}^{3}$: $\langle\hat{v},\hat{w}\rangle_{\mathrm{tr}}=v^{\top}w$.
+In particular, $\mathrm{Ad}$-invariant scalar contractions on $\mathfrak{so}(3)$ recover the invariants used by Vector Neurons  Deng et al.
+(2021)(up to a fixed global scaling, which can be absorbed by adjacent learnable linear maps or normalization).
+Appendix C Proofs of Key Theorems
+In this section, we provide a generalized treatment and proofs for the modified bilinear form $\widetilde{B}$ introduced in Section 4 and Appendix B.
+Our discussion here establishes the formal properties of $\widetilde{B}$ for any real reductive Lie algebra.
+C.1 Proof of Non-degeneracy and Ad-invariance of Modified Killing Form $B_{e}$
 Let $\mathfrak{g}$ be a real reductive Lie algebra.
-Definition B.1 (Reductive decomposition).
+Definition C.1 (Reductive decomposition).
 A Lie algebra $\mathfrak{g}$ is reductive if $\mathfrak{g}=\mathfrak{z}(\mathfrak{g})\oplus[\mathfrak{g},\mathfrak{g}]$, where $\mathfrak{z}(\mathfrak{g})$ is the center and $\mathfrak{g},\mathfrak{g}]$ is semisimple.
 This decomposition is canonical (both summands are ideals).
-Definition B.2 (Modified Killing form on a reductive Lie algebra).
+Definition C.2 (Modified Killing form on a reductive Lie algebra).
 Fix any symmetric, positive–definite inner product $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ on $\mathfrak{z}(\mathfrak{g})$, and let $B$ denote the Killing form on the semisimple ideal $\mathfrak{g},\mathfrak{g}]$.
@@
 $\widetilde{B}(Z_{1}{+}X_{1},\,Z_{2}{+}X_{2})\;:=\;\langle Z_{1},Z_{2}\rangle_{\mathfrak{z}}\;+\;B(X_{1},X_{2}).$ (31)
-Remark 1 (Canonicity).
+Remark C.3 (Canonicity).
 On $\mathfrak{g},\mathfrak{g}]$ the restriction (Killing form) is canonical.
@@
 The choice we make in the case of $\mathfrak{gl}(n)$ ensures that it agrees with the Killing form on the semisimple part $\mathfrak{sl}(n)$, and the center $\mathbb{R}I$ is normalized by a natural trace scale.
-Proposition B.1 (Block–orthogonality and restrictions).
+Proposition C.4 (Block–orthogonality and restrictions).
 With notation as above,
 $\widetilde{B}\big(\mathfrak{z}(\mathfrak{g}),[\mathfrak{g},\mathfrak{g}]\big)=0,\qquad\widetilde{B}|_{\mathfrak{z}(\mathfrak{g})}=\langle\cdot,\cdot\rangle_{\mathfrak{z}},\qquad\widetilde{B}|_{[\mathfrak{g},\mathfrak{g}]}=B.$ (32)
-Proposition B.2 (Non–degeneracy).
+Proposition C.5 (Non–degeneracy).
 $\widetilde{B}$ is nondegenerate on $\mathfrak{g}$.
@@
 ∎
-Proposition B.3 ($\mathrm{Ad}$–invariance on the identity component).
+Proposition C.6 ($\mathrm{Ad}$–invariance on the identity component).
 $\widetilde{B}$ is $\mathrm{ad}$–invariant:
@@
 If $Z\in\mathfrak{z}(\mathfrak{g})$ then $[X,Z]=0$ for all $X$, so any bilinear form on $\mathfrak{z}(\mathfrak{g})$ is automatically $\mathrm{ad}$–invariant.
-Using Proposition B.1 and bilinearity gives the displayed identity.
+Using Proposition C.4 and bilinearity gives the displayed identity.
 Equivalence with $\mathrm{Ad}$–invariance on $G^{\circ}$ follows by integrating the infinitesimal relation along paths in $G^{\circ}$.
 ∎
-Remark 2 (Invariance for nonconnected groups).
+Remark C.7 (Invariance for nonconnected groups).
 In case the group is nonconnected, and one desires invariance under the full group $G$ (not just $G^{\circ}$).
@@
 $\langle Z_{1},Z_{2}\rangle_{\mathfrak{z}}^{\mathrm{avg}}=\frac{1}{|\Gamma|}\sum_{\gamma\in\Gamma}\big\langle\mathrm{Ad}_{\gamma}Z_{1},\,\mathrm{Ad}_{\gamma}Z_{2}\big\rangle_{\mathfrak{z}}.$ (34)
-Replacing $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ by $\langle\cdot,\cdot\rangle_{\mathfrak{z}}^{\mathrm{avg}}$ in 18 makes $\widetilde{B}$ invariant under all of $G$.
-Appendix C Connections to existing bilinear forms
-We demonstrate how the trace-based form $\widetilde{B}$ (Eq.
-2) unifies and recovers prior bilinear constructions in the regimes used by Lie Neurons (Lin et al., 2024a) and Vector Neurons (Deng et al., 2021) .
-The discussion below states precise conditions under which $\widetilde{B}$ (i) equals the Killing-form contractions on semisimple inputs and (ii) is proportional to the Vector Neuron inner product under the $\mathfrak{so}(3)\simeq\mathbb{R}^{3}$ isomorphism.
-Our single form applies on the full reductive algebra $\mathfrak{gl}(n)$ including these specialized approaches, which encompasses general $n\times n$ matrix-valued inputs.
-C.1 Reduction to the Killing form on the semisimple ideal
-Write $X=X_{0}+\tfrac{1}{n}\mathrm{tr}(X)I$ with $X_{0}\in\mathfrak{sl}(n)$. Using Eq. 2 we obtain
-$\widetilde{B}(X,Y)=2n\,\mathrm{tr}(X_{0}Y_{0})+n^{2}\!\left(\tfrac{1}{n}\mathrm{tr}X\right)\!\left(\tfrac{1}{n}\mathrm{tr}Y\right).$ (22)
-Hence, when inputs are restricted to the semisimple ideal $\mathfrak{g},\mathfrak{g}]=\mathfrak{sl}(n)$ (so $\mathrm{tr}(X)=\mathrm{tr}(Y)=0$), the center contribution vanishes and
-$\widetilde{B}|_{\mathfrak{sl}(n)}(X,Y)=2n\cdot\mathrm{tr}(XY)=B_{\mathfrak{sl}(n)}(X,Y),$ (23)
-i.e.
-$\widetilde{B}$ coincides with the (scaled) Killing form used in Lie Neurons.
-More generally, for any semisimple subalgebra $\mathfrak{h}\subset[\mathfrak{g},\mathfrak{g}]$, the restriction $\widetilde{B}|_{\mathfrak{h}}$ matches the Killing-form contraction on $\mathfrak{h}$ up to global scaling.
-C.2 $\mathfrak{so}(3)$ example: recovery of the Vector Neuron inner product
-Recall the hat map $\widehat{(\cdot)}:\mathbb{R}^{3}\to\mathfrak{so}(3)$ with
-$\widehat{v}=\begin{bmatrix}0&-v_{3}&v_{2}\\ v_{3}&0&-v_{1}\\ -v_{2}&v_{1}&0\end{bmatrix},\qquad\mathrm{tr}(\widehat{v}\,\widehat{w})=-2\,v^{\top}w.$ (24)
-For $X=\widehat{v}$ and $Y=\widehat{w}$, Eq. 2 gives
-$\widetilde{B}(X,Y)=2n\cdot\mathrm{tr}(XY)=2n\cdot(-2\,v^{\top}w)=(-4n)\,v^{\top}w,$ (25)
-i.e.
-$\widetilde{B}$ is proportional to the Euclidean inner product on $\mathbb{R}^{3}$.
-The proportionality constant depends only on $n$; in practice this constant is absorbed by adjacent learnable linear layers or normalization, yielding behaviour identical to the inner product used in Vector Neurons to harmless scaling.
-These remarks justify using the single, nondegenerate $\widetilde{B}$ across heterogeneous input types while preserving compatibility with prior architectures.
-Appendix D Detailed Layer Formulations
-This section provides the precise mathematical definitions and equivariance proofs for the core components of the ReLN architecture.
-We consider the input to a layer as a tensor $x\in\mathbb{R}^{K\times C}$, representing $C$ feature channels where $K=\dim\mathfrak{g}$.
-Each column $x_{c}\in\mathbb{R}^{K}$ is the vector representation of a feature.
-We use the wedge ($\wedge$) and vee ($\vee$) operators to map between the vector form $x_{c}$ and the matrix form $X_{c}\in\mathfrak{g}$.
-D.1 Equivariant Linear Layer
-The ReLN-Linear layer applies a linear map to the channel dimension of the input tensor $x$:
-$f_{\mathrm{ReLN-Lin}}(x;W)=xW,\quad\text{where }W\in\mathbb{R}^{C\times C^{\prime}}.$ (26)
+Replacing $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ by $\langle\cdot,\cdot\rangle_{\mathfrak{z}}^{\mathrm{avg}}$ in Equation 31 makes $\widetilde{B}$ invariant under all of $G$.
+Appendix D Summary of Key Lie Groups and Algebras
+To provide a comprehensive overview of the geometric structures discussed in this work, Table 7 summarizes the Lie groups $G$, their associated Lie algebras $\mathfrak{g}$, and their application domains.
+We categorize each group by its algebraic properties—specifically distinguishing between reductive, semisimple, and compact types—and detail their specific roles within both the broader literature and our experimental framework.
+This unified taxonomy situates ReLNs as a general-purpose backbone capable of handling a wide array of linear symmetries encountered in scientific machine learning.
+Table 7: Comprehensive survey of Lie groups, algebras, and their applications.
+We highlight the algebraic classifications that dictate the choice of invariant bilinear forms and contrast general literature examples with our specific experimental tasks.
+Group ($G$) Algebra ($\mathfrak{g}$) Algebraic Type General Applications & Refs Relevance to ReLN Framework
+General Linear $\mathfrak{gl}(n)$ Reductive General linear transformations, stress-strain tensors (Basu et al., 2025; Finzi et al., 2021).
+Core Domain.
+Primary target of ReLNs; resolves the Killing form degeneracy on $\mathfrak{gl}(n)$.
+Special Linear $\mathfrak{sl}(3)$ Semisimple Homography classification, 3D vision, SLAM (Lin et al., 2024a; Finzi et al., 2021).
+Algebraic Benchmark.
+Validates generalization on semisimple non-compact subgroups (Section  5.1.1).
+Symplectic $\mathfrak{sp}(4)$ Semisimple Hamiltonian mechanics, phase space dynamics (Lin et al., 2024a; Finzi et al., 2021).
+Algebraic Benchmark.
+Tests the ability to learn highly nonlinear invariants on conserved systems (Section  5.1.2).
+Special Orthogonal $\mathfrak{so}(3)$ Compact 3D Point clouds, state estimation, rigid body rotations (Deng et al., 2021; Son et al., 2024; Thomas et al., 2018).
+Robotics Application.
+Lifts 3D velocity vectors into $\mathfrak{so}(3)\subset\mathfrak{gl}(3)$ (Sections 5.2 and 5.3).
+Lorentz Group $\mathfrak{so}(1,3)^{+}$ Semisimple Particle physics, jet tagging, relativistic collisions (Bogatskiy et al., 2020; Batatia et al., 2023).
+Physics Application.
+4-momenta embedded into $\mathfrak{gl}(5)$ to unify left-action with adjoint action (Section  H.1.
+Special Unitary $\mathfrak{su}(3)$ Compact Quantum Chromodynamics (QCD), lattice gauge theory (Favoni et al., 2022).
+General Example.
+Illustrates framework applicability to complex-valued compact groups.
+SPD Manifold $\mathrm{SPD}(n)$ Riemannian Geometric uncertainty, inertia tensors, diffusion MRI (Ma et al., 2025; Magnus, 1985).
+Geometric Uncertainty.
+Covariances mapped to $\mathfrak{gl}(n)$ (Sections 5.2 and 5.3).
+Note: $\mathrm{SPD}(n)$ is not a group but a manifold representable as the quotient space $\mathrm{GL}(n)/\mathrm{O}(n)$.
+We process it by mapping to the linear subspace $\mathfrak{sym}(n)\subset\mathfrak{gl}(n)$ via the matrix logarithm (Ma et al., 2025).
+Appendix E Detailed Layer Formulations
+This section provides the precise mathematical definitions and equivariance proofs for the ReLN architecture.
+To maintain consistency with the main text, we denote the input tensor as $x\in\mathbb{R}^{B\times K\times C}$, where $B$ is the batch size, $K=\dim\mathfrak{g}$ is the geometric dimension, and $C$ is the number of feature channels.
+Each column $x_{b,:,c}\in\mathbb{R}^{K}$ corresponds to a matrix $X_{b,c}\in\mathfrak{g}$ via the vee/hat isomorphism.
+E.1 Equivariant Linear Layer
+The ReLN-Linear layer applies a linear map to the channel dimension (the last axis).
+For an input $x\in\mathbb{R}^{B\times K\times C}$ and weights $W\in\mathbb{R}^{C\times C^{\prime}}$, the operation is defined as:
+$f_{\mathrm{ReLN-Lin}}(x;W)=xW\in\mathbb{R}^{B\times K\times C^{\prime}}.$ (35)
 We omit any bias term to preserve exact equivariance.
@@
 $\begin{split}f_{\mathrm{ReLN-Lin}}(\mathrm{Ad}_{g}(x);W)&=(\mathrm{Ad}_{g}x)W\\ &=\mathrm{Ad}_{g}(xW)\\ &=\mathrm{Ad}_{g}(f_{\mathrm{ReLN-Lin}}(x;W)).\end{split}$ (36)
-D.2 Equivariant Nonlinearities
+E.2 Equivariant Nonlinearities
 Standard pointwise activations break equivariance under non-orthogonal transforms. We introduce two equivariant alternatives.
@@
 This layer rectifies a feature based on its alignment with a learnable direction.
-Given the input tensor $x$, we first compute per-channel vector directions $d=xU$.
-The nonlinearity for the input $x$ is then defined as:
-$f_{\mathrm{ReLN-ReLU}}(x)=\begin{cases}x,&\text{if }\widetilde{B}(x^{\wedge},d^{\wedge})\leq 0,\\ x+\widetilde{B}(x^{\wedge},d^{\wedge})d,&\text{otherwise}.\end{cases}$ (28)
-Since all operations—the linear map to compute $d_{i}$, the bilinear form $\widetilde{B}$, and vector addition/scaling—are equivariant, the entire function is equivariant.
-The leaky variant $f_{\mathrm{ReLN-LeakyReLU}}(x)=\alpha\,x+(1-\alpha)\,f_{\mathrm{ReLN-ReLU}}(x)$ follows directly.
+Given the input $x\in\mathbb{R}^{B\times K\times C}$, we first compute equivariant directions $d=f_{\mathrm{ReLN-Lin}}(x;U)$ where $U\in\mathbb{R}^{C\times C}$.
+The nonlinearity is then applied channel-wise as a scalar-gated update:
+$f_{\mathrm{ReLN-ReLU}}(x)_{b,:,c}=\begin{cases}x_{b,:,c},&\text{if }\widetilde{B}(x_{b,:,c}^{\wedge},d_{b,:,c}^{\wedge})\leq 0,\\ x_{b,:,c}+\sigma\big(\widetilde{B}(x_{b,:,c}^{\wedge},d_{b,:,c}^{\wedge})\big)d_{b,:,c},&\text{otherwise},\end{cases}$ (37)
+[…]
 ReLN-Bracket (Lie-bracket nonlinearity).
 This layer uses the matrix commutator, a natural $\mathrm{Ad}$-equivariant primitive, to create learnable interactions between channels.
-Let the input be a batch of features represented by their vector coordinates, $x\in\mathbb{R}^{B\times C_{\mathrm{in}}\times K}$.
-First, two independent linear maps (with learnable weights $W_{a},W_{b}\in\mathbb{R}^{C_{\mathrm{in}}\times C_{\mathrm{out}}}$) transform the input features into two intermediate tensors, $u,v\in\mathbb{R}^{B\times C_{\mathrm{out}}\times K}$.
-The Lie bracket is then computed channel-wise between the corresponding feature vectors of $u$ and $v$.
-This produces an update tensor, $\Delta x\in\mathbb{R}^{B\times C_{\mathrm{out}}\times K}$, where each vector is defined as:
-$(\Delta x)_{b,c^{\prime},:}=\left[(u_{b,c^{\prime},:})^{\wedge},(v_{b,c^{\prime},:})^{\wedge}\right]^{\vee}.$ (29)
-This update is added to the input for a residual connection (requiring $C_{\mathrm{in}}=C_{\mathrm{out}}$ for the shapes to match):
+Let the input be a batch of features represented by their vector coordinates, $x\in\mathbb{R}^{B\times K\times C_{\mathrm{in}}}$.
+First, two independent linear maps with weights $W_{a},W_{b}\in\mathbb{R}^{C_{\mathrm{in}}\times C_{\mathrm{out}}}$ produce intermediate tensors $u=xW_{a}$ and $v=xW_{b}$.
+The Lie bracket is then computed channel-wise between the corresponding feature vectors of $u$ and $v$ to produce an update tensor $\Delta x\in\mathbb{R}^{B\times K\times C_{\mathrm{out}}}$:
+$(\Delta x)_{b,:,c^{\prime}}=\left[(u_{b,:,c^{\prime}})^{\wedge},(v_{b,:,c^{\prime}})^{\wedge}\right]^{\vee}.$ (38)
+This update is added to the input via a residual connection (requiring $C_{\mathrm{in}}=C_{\mathrm{out}}$):
 $f_{\mathrm{ReLN-Bracket}}(x)=x+\Delta x.$ (39)
-Each step in this process (linear map, the Lie bracket, and vee/hat operations) is equivariant under the adjoint action, making the entire block equivariant $f_{\mathrm{ReLN-Bracket}}(\mathrm{Ad}_{g}x)=\mathrm{Ad}_{g}(f_{\mathrm{ReLN-Bracket}}(x))$ for all $g\in G$.
-Appendix E Experimental Details
+Each step in this process is equivariant under the adjoint action, making the entire block exactly $\mathrm{Ad}_{g}$-equivariant: $f_{\mathrm{ReLN-Bracket}}(\mathrm{Ad}_{g}x)=\mathrm{Ad}_{g}(f_{\mathrm{ReLN-Bracket}}(x))$ for all $g\in G$.
+Appendix F Theoretical Complexity and Scaling Analysis
+To address the operational properties of our architecture, we provide an explicit Big-$\mathcal{O}$ complexity analysis connecting the feature channel width $C$ and the Lie algebra matrix dimension $n$.
+We further validate this analysis with empirical benchmarks.
+1. Theoretical Scaling with Matrix Dimension $n$.
+We fix the total number of scalars per token to $C$ to ensure a fair comparison. The effective channel width $M$ is scaled such that $M\times K=C$.
+  • Standard MLP ($K=1$): Uses full width $M=C$.
+  • ReLN ($K=n^{2}$): Operates on $\mathfrak{gl}(n)$ matrices. Effective width $M=C/n^{2}$.
+Let $n$ be the dimension of the matrices (e.g., $n=3$ for $\mathfrak{gl}(3)$).
+  • Linear Layer: The layer performs $K$ independent matrix multiplications of size $M\times M$.
+    $\text{Cost}_{\text{linear}}=K\cdot M^{2}=n^{2}\cdot\left(\frac{C}{n^{2}}\right)^{2}=\frac{C^{2}}{n^{2}}.$ (40)
+    The dominant quadratic cost decreases as the matrix dimension $n$ increases. For $n=3$, ReLN reduces the Linear Layer FLOPs by a factor of $1/9$ compared to an MLP ($n=1$).
+  • Lie Bracket Nonlinearity: This operation computes [X,Y]=XYYX[X,Y]=XY-YX, involving $n\times n$ matrix multiplications.
+    $\text{Cost}_{\text{bracket}}=M\cdot\mathcal{O}(n^{3})=\frac{C}{n^{2}}\cdot\mathcal{O}(n^{3})=\mathcal{O}(C\cdot n).$ (41)
+    In deep learning contexts where $C$ is large (e.g., $C\geq 128$), the quadratic term $\mathcal{O}(C^{2}/n^{2})$ dominates the linear term $\mathcal{O}(Cn)$.
+2. Layer-wise Complexity Comparison.
+Table 8 presents the theoretical breakdown. ReLN-Bracket reduces the total block cost to $\approx 12.5\%$ of the baseline.
+Table 8: Theoretical Complexity Analysis.
+Comparison under a fixed feature budget $C$.
+By utilizing a higher-dimensional geometric space ($n=3,K=9$), ReLN reduces the effective channel width to $C/9$, resulting in a quadratic reduction in the dominant Linear Layer cost.
+Layer Type Model Eff. Channels ($M$) Parameters FLOPs Relative Cost
+Linear Layer MLP ($n=1$) $C$ $C^{2}$ $C^{2}$ $1.00\times$
+VN ($n\approx 1.7$) $C/3$ $0.33C^{2}$ $C^{2} 0.33×0.33\times$
+ReLN ($n=3$) $\mathbf{C/9}$ $\mathbf{0.11C^{2}}$ $\mathbf{0.11C^{2}}$ $\mathbf{0.11\times}$
+Nonlinearity ReLU (MLP) $C$ 0 $\mathcal{O}(C)$ $\approx 0.00\times$
+VN-ReLU $C/3$ $0.11C^{2}$ $C^{2} 0.11×0.11\times$
+ReLN-ReLU $C/9$ $01C^{2} 0.0$$1C^{2} 0.0$$01\times$
+ReLN-Bracket $\mathbf{C/9}$ $\mathbf{0.025C^{2}}$ $\mathbf{0.025C^{2}}$ $\mathbf{0.03\times}$
+Appendix G Experimental Details
 Training and evaluation for all presented experiments, Platonic solid classification, invariant function regression, top tagging, and drone state estimation, were conducted on a single NVIDIA GeForce RTX 4090 GPU.
-E.1 Model Architectures and Implementation Details
+G.1 Model Architectures and Implementation Details
 Across all experiments, our proposed ReLN models are constructed by stacking ReLN-Linear, ReLN-ReLU, and ReLN-Bracket layers.
@@
 We adapt its Lorentz Group Equivariant Blocks (LGEBs) by replacing the invariant feature computation with our proposed bilinear form.
-A detailed description of the architecture, our modifications, and training protocol is provided in Appendix F.
+A detailed description of the architecture, our modifications, and training protocol is provided in Appendix H.
 Drone State Estimation.
 In this task, we compare our ReLN model against two baseline families: a non-equivariant 1D ResNet and an equivariant Vector Neurons (VN) model.
-The specific implementation details and architectural choices for each model are provided next in Appendix G.
-E.2 Platonic Solid Classification on $\mathfrak{sl}(3)$
+The specific implementation details and architectural choices for each model are provided next in Appendix I.
+G.2 Platonic Solid Classification on $\mathfrak{sl}(3)$
 Overview.
@@
 Learning rate (ReLNs/Lie Neurons models) $3\times 10^{-6}$ Lower LR chosen for stable training
-Appendix F Top Tagging Experiment: Framework, Proof, and Implementation
-This appendix provides the complete details for our jet tagging experiment.
-We first present the geometric framework and the mathematical proof of our Lorentz-equivariant embedding, and then describe the model architecture and training protocol.
-F.1 Group Action Equivariance via Embedding Map
-To process four-momenta within our Lie-algebraic framework, we require an embedding that translates the action of the Lorentz group into an adjoint action on a matrix space.
-This is achieved by lifting the four-vector into $\mathfrak{gl}(5)$.
-Definition F.1 (Lorentz-Compatible Embedding).
-Given a four-vector $p\in\mathbb{R}^{4}$ and the Minkowski metric $\eta=\operatorname{diag}(1,-1,-1,-1)$, we define its embedding $\varphi(p)$ into $\mathfrak{gl}(5)$ as:
+Appendix H Top Tagging Experiment: Framework, Proof, and Implementation
+This appendix provides details for our Lorentz-equivariant jet tagging study.
+We first summarize the task and results, then present the Lorentz-compatible embedding and its equivariance proof, and finally describe the model and training protocol.
+H.1 Particle Physics with Lorentz Group $\mathrm{SO}^{+}(1,3)$ Equivariance
+We evaluate on the Top-Tagging benchmark (Kasieczka et al., 2019), which classifies particle jets originating from top quarks against a Quantum Chromodynamics background.
+Since relativistic collisions respect spacetime symmetries, the model should be equivariant under the proper orthochronous Lorentz group $\mathrm{SO}^{+}(1,3)$.
+To express this as an adjoint action within our Lie-algebraic framework, we embed each four-momentum $p\in\mathbb{R}^{1,3}$ into $\mathfrak{gl}(5)$ via
+$\varphi(p)\;=\;\begin{bmatrix}0_{4\times 4}&p\\ p^{\top}\eta&0\end{bmatrix},\qquad\eta=\mathrm{diag}(-1,1,1,1),$ (42)
+and show in Appendix H.2 that $\varphi(\Lambda p)=\mathrm{Ad}_{\mathrm{diag}(\Lambda,1)}\big(\varphi(p)\big)$ for $\Lambda\in\mathrm{SO}^{+}(1,3)$.
+For comparison, we modify LorentzNet by replacing its invariant feature computation with our bilinear-form construction and additionally report a parameter-matched LorentzNet baseline.
+Table 11 shows that ReLN achieves performance comparable to the parameter-matched LorentzNet across Acc./AUC and Rej@30%, with differences within the reported uncertainty (i.e., overlapping standard deviations).
+Table 11: Comparison of performance on the Top-Tagging dataset.
+ReLN results include both the parameter-efficient version (84k) and the parameter-matched version (224k).
+Rej@30% denotes background rejection at 30% signal efficiency (higher is better).
+Benchmark scores are as reported in the original publications.
+Architecture #Params Acc. AUC Rej@30% Reference
+PELICAN 45k 0.943 0.987 $2289\pm 204$ Bogatskiy et al. (2022)
+LorentzNet (orig.) 224k 0.942 0.987 $2195\pm 173$ Gong et al. (2022)
+LorentzNet (reprod.) 84k 0.942 0.987 $1821\pm 94$ Our reprod.
+LGN 4.5k 0.929 0.964 $435\pm 95$ Bogatskiy et al. (2020)
+BIP 4k 0.931 0.981 $853\pm 68$ Munoz et al. (2022)
+partT 2.14M 0.940 0.986 $1602\pm 81$ Qu et al. (2022)
+ParticleNet 498k 0.938 0.985 $1298\pm 46$ Qu and Gouskos (2020)
+EFN 82k 0.927 0.979 $633\pm 31$ Komiske et al. (2019)
+TopoDNN 59k 0.916 0.972 $295\pm 5$ Pearkes et al. (2017)
+LorentzMACE 228k 0.942 0.987 $1935\pm 85$ Batatia et al. (2023)
+CGENN 321K 0.942 0.987 $2172$ Ruhe et al. (2023)
+Lie Neurons 224k 0.941 0.985 $1655\pm 73$ Our reprod.
+ReLN (Ours) 224k 0.942 0.987 $2201\pm 101$
+ReLN (Ours) 84k 0.942 0.987 $1979\pm 87$
+H.2 Group Action Equivariance via Embedding Map
+To process four-momenta within our Lie-algebraic framework, we require an embedding that translates the Lorentz action into an adjoint action on a matrix space.
+Definition H.1 (Lorentz-Compatible Embedding).
+Given $p\in\mathbb{R}^{4}$ and $\eta=\operatorname{diag}(-1,1,1,1)$, define
 $\varphi(p)=\begin{bmatrix}0_{4\times 4}&p\\ p^{\top}\eta&0\end{bmatrix}$. (43)
-Theorem F.1 (Adjoint Equivariance).
-The embedding $\varphi$ correctly models the Lorentz group action.
-For any $p\in\mathbb{R}^{4}$ and Lorentz transformation $\Lambda\in\operatorname{SO}(1,3)$, let $G=\operatorname{diag}(\Lambda,1)\in\operatorname{GL}(5)$.
-The map is equivariant in the sense that the standard action on $p$ corresponds to the adjoint action on $\varphi(p)$:
+Theorem H.2 (Adjoint Equivariance).
+For any $p\in\mathbb{R}^{4}$ and $\Lambda\in\mathrm{SO}^{+}(1,3)$, let $G=\operatorname{diag}(\Lambda,1)\in\operatorname{GL}(5)$. Then
 $\operatorname{Ad}_{G}(\varphi(p))=G\varphi(p)G^{-1}=\varphi(\Lambda p).$ (44)
@@
 Since the condition holds, the proof is complete. ∎
-Remark 3 (Generalization to Orthogonal Groups).
-This embeddinging technique is not limited to the Lorentz group and can be readily generalized to any orthogonal group $\operatorname{O}(n)$ or special orthogonal group $\operatorname{SO}(n)$.
+Remark H.3 (Generalization to Orthogonal Groups).
+This embedding technique is not limited to the Lorentz group and can be readily generalized to any orthogonal group $\operatorname{O}(n)$ or special orthogonal group $\operatorname{SO}(n)$.
 For instance, in applications involving 3D point clouds where the symmetry is $\operatorname{SO}(3)$, a vector $p\in\mathbb{R}^{3}$ would be embedded into the Lie algebra $\mathfrak{gl}(4)$ as:
@@
 This highlights the broad applicability of our embedding strategy to any benchmark involving norm-preserving group transformations.
-F.2 Experimental Implementation
-Dataset
+H.3 Experimental Implementation
+Dataset.
 The experiment uses the Top-Tagging dataset (Kasieczka et al., 2019), which contains 2 million simulated proton-proton collision events.
@@
 Each jet is represented as a set of constituent particles, each with four-momentum $p=(E,p_{x},p_{y},p_{z})$.
-Model
-Our model leverages the established architecture of LorentzNet  (Gong et al., 2022), utilizing its stack of Lorentz Group Equivariant Blocks (LGEBs) for message passing on the jet’s particle cloud.
-While the original LorentzNet computes these features directly from the 4-momenta using the Minkowski inner product, our approach introduces a modified bilinear form based feature extraction.
-We first embed each pair of 4-momenta, $p_{i}$ and $p_{j}$, from the Minkowski space $\mathbb{R}^{1,3}$ into the Lie algebra $\mathfrak{gl}(5)$ via the map $p\mapsto X(p)$.
-The invariant features for the message passing are then derived from the bilinear form, $B(\cdot,\cdot)$, on this Lie algebraic space.
-The edge message $m_{ij}$ is thus constructed as:
-$m_{ij}=\phi_{e}\Big(h_{i},h_{j},\psi\big(\widetilde{B}(X(p_{i}),X(p_{i}))\big),\psi\big(\widetilde{B}(X(p_{i}),X(p_{j}))\big)\Big)$ (38)
+Model.
+[…]
+$m_{ij}=\phi_{e}\Big(h_{i},h_{j},\psi\big(\widetilde{B}(\varphi(p_{i}),\varphi(p_{i}))\big),\psi\big(\widetilde{B}(\varphi(p_{j}),\varphi(p_{j}))\big),\psi\big(\widetilde{B}(\varphi(p_{i}),\varphi(p_{j}))\big)\Big)$ (50)
 where $h_{i},h_{j}$ are scalar features, $\phi_{e}$ is an MLP, and $\psi$ is a stabilizing nonlinearity.
@@
 Edge feature computation Minkowski inner prod. Minkowski inner prod. Bilinear invariant form
-Training Setup
+Training Setup.
 For a fair comparison, our training procedure closely follows the protocol established in the LorentzNet (Gong et al., 2022).
-The model was trained for a total of 35 epochs on a single NVIDIA RTX 4090 GPU.
+The model was trained for a total of 35 epochs on a NVIDIA RTX 4090 GPU.
 We used the AdamW optimizer with a weight decay of 0.01 and a batch size of 128, matching the total effective batch size from the reference work.
@@
 After each epoch, the model with the highest validation accuracy was saved for final evaluation on the test set.
-Appendix G Drone Experiment Details
+Appendix I Drone Experiment Details
 This appendix provides the technical details for the drone state estimation experiment, including the theoretical framework, dataset generation, model implementations, and formal proofs.
-G.1 Geometric Framework for Equivariant Covariance Processing
+I.1 Geometric Framework for Equivariant Covariance Processing
 Our approach leverages the geometry of symmetric positive-definite matrices.
@@
 The following theorem proves that the congruence transformation on $C\in\mathrm{SPD}(n)$ becomes an adjoint action on its image $\log C\in\mathrm{Sym}(n)$, thus preserving the equivariant structure required by our model.
-Theorem G.1 (Equivariance of the Logarithmic Map).
+Theorem I.1 (Equivariance of the Logarithmic Map).
 For any $C\in\mathrm{SPD}(n)$ and any rotation matrix $R\in\mathrm{SO}(n)$, the congruence transformation on $C$ corresponds to an adjoint action on its logarithm:
@@
 Consequently, ReLNs realize $\mathrm{SO}(3)$-equivariance by construction, avoiding the need for the model to learn these symmetries from data.
-G.2 Dataset Generation.
+I.2 Dataset Generation.
 We use the PyBullet engine to simulate 200 aggressive trajectories for a Crazyflie-like nano-quadrotor.
@@
 The steepness $\lambda$ is set to 0.8, and the midpoint velocity $v_{\text{mid}}$ is dynamically adjusted based on the estimated average speed of each trajectory to ensure a realistic noise profile.
-G.3 Baseline and Model Implementation Details
+I.3 Baseline and Model Implementation Details
 We compare ReLN against two baseline classes chosen to isolate the effect of geometric priors.
@@
 The ResNet (velocity + covariance) model receives the flattened $3\times 3$ covariance matrix concatenated to the velocity vector.
-Eigendecomposition-based $\mathrm{SO}(3)$-Equivariant Baseline.
+Eigendecomposition-based $\mathbf{SO}\textbf{(3)}$-Equivariant Baseline.
 This model adapts the 1D ResNet backbone for $\mathrm{SO}(3)$ equivariance using VN layers.
@@
 The two latent features from both streams are fused at the final output layer. Eigenvector ambiguities (sign or multiplicities) are resolved via a deterministic, rotation-equivariant canonicalization.
+Spherical Harmonic (SH)-based Equivariant Baselines (TFN and SE(3)-Transformer).
+We evaluate two representative $\mathrm{SO}(3)$-equivariant graph architectures built from irrep-valued features and steerable bases expanded in spherical harmonics (SH): Tensor Field Networks (TFN) (Thomas et al., 2018) and the SE(3)-Transformer (Fuchs et al., 2020).
+Both enforce equivariance by restricting linear maps to $\mathrm{SO}(3)$ intertwiners, implemented through SH steerable kernels and tensor-product (Clebsch–Gordan) couplings.
+Transformation Rules and Irrep Decomposition.
+Under a rotation $R\in\mathrm{SO}(3)$, the velocity transforms as $\mathbf{v}\mapsto R\mathbf{v}$, while a covariance transforms by congruence $C\mapsto RCR^{\top}$.
+We represent the symmetric rank-2 covariance via the standard irrep decomposition $\mathrm{Sym}^{2}(\mathbb{R}^{3})\cong L{=}0\oplus L{=}2$, i.e., $C=\tfrac{\mathrm{tr}(C)}{3}I+C_{0}$ with $\mathrm{tr}(C)$ as a type-0 scalar and $C_{0}:=C-\tfrac{\mathrm{tr}(C)}{3}I$ as a type-$2$ traceless-symmetric tensor.
+Feature construction and temporal graph.
+Each timestamp is encoded as an $\mathrm{SO}(3)$ fiber with degrees $L\in\{0,1,2\}$: $\mathrm{tr}(C)$ for L=0L{=}0, $\mathbf{v}$ for L=1L{=}1, and $C_{0}$ for L=2L{=}2.
+We form a temporal chain graph connecting $t$ to $t\pm 1$ and drive message passing using edge attributes derived from relative offsets.
+TFN aggregates messages via steerable kernel filtering (with learned radial profiles), whereas SE(3)-Transformer replaces fixed kernel aggregation with multi-head, content-dependent equivariant self-attention (Thomas et al., 2018; Fuchs et al., 2020).
+In both cases, uncertainty is stored in separate type-0 and type-$2$ blocks, and interactions between them are mediated by learned equivariant tensor-product couplings rather than by treating $C$ as a single matrix feature throughout the network.
+Relation to our formulation.
+In contrast, ReLN embeds $(\mathbf{v},C)$ into a unified matrix space $\mathfrak{gl}(3)$ and processes both under a single adjoint action, directly matching the congruence rule $C\mapsto RCR^{\top}$ in a shared Lie-algebraic backbone.
+Empirical observation.
+In our experiments, TFN consistently outperformed the SE(3)-Transformer.
+We attribute this to the nature of drone state estimation as a path integration task.
+The SH-based convolutional kernels in TFN act as learned local integrators, which are better suited for capturing high-frequency temporal dynamics than the global dependency focus of attention mechanisms.
+While attention provides flexibility, it can introduce susceptibility to high-frequency noise in aggressive flight sequences, whereas the local inductive bias of SH-convolutions provides a more stable prior for trajectory reconstruction.
 Reductive Lie Neurons (ReLNs).
-The ReLN model shares the same base architecture as the VN model but incorporates ReLN-Bracket layer as an additional source of nonlinearity after the initial feature extraction block.
-In contrast to the VN backbone, ReLNs provide a unified framework for velocity and covariance processing.
-Velocities $\mathbf{v}\in\mathbb{R}^{3}$ are lifted into the Lie algebra $\mathfrak{so}(3)$ via $K=\mathbf{v}^{\wedge}$, while covariance $C$ (or $\log C$) is treated as a structured geometric input.
-Both transform under the same adjoint action: $K^{\prime}=RKR^{\top}$ and $C^{\prime}=RCR^{\top}$, enabling joint equivariant processing.
-Although $\mathrm{SPD}(n)$ is not a Lie algebra or group, it is a subset of $\mathrm{GL}(n)$.
-By taking the matrix logarithm, $\log C\in\mathrm{Sym}(n)\subset\mathfrak{gl}(n)$, covariances are embedded into a linear subspace compatible with Lie-algebra processing.
-The network $\Phi$ fuses these inputs into a single matrix $A\in\mathbb{R}^{3\times 3}$, from which we extract the velocity estimate:
-$A_{\mathrm{skew}}\coloneqq\tfrac{1}{2}(A-A^{\top})\in\mathfrak{so}(3),\qquad\hat{\mathbf{v}}=\text{Vee}(A_{\mathrm{skew}}).$ (46)
-The extracted velocity is provably $\mathrm{SO}(3)$-equivariant by construction. See Appendix H for the full statement and proof.
-G.4 Training and Evaluation Protocol
-Problem Formulation
+The ReLN model shares a similar backbone but incorporates the ReLN-Bracket layer.
+In contrast to the VN and graph-based baselines, ReLNs provide a unified framework for velocity and covariance processing within the Lie algebra $\mathfrak{gl}(3)$.
+Velocities $\mathbf{v}\in\mathbb{R}^{3}$ are lifted into $\mathfrak{so}(3)\subset\mathfrak{gl}(3)$ via $K=\mathbf{v}^{\wedge}$, while covariance $C$ (or $\log C$) is treated as a structured geometric input.
+Both transform under the same adjoint action: $K^{\prime}=RKR^{\top}$ and $C^{\prime}=RCR^{\top}$.
+By utilizing the matrix logarithm, $\log C\in\mathrm{Sym}(3)\subset\mathfrak{gl}(3)$, we embed the manifold-valued uncertainty into a linear subspace compatible with Lie-algebraic processing.
+The final velocity estimate is equivariantly extracted via the projection $\tilde{\mathbf{v}}=(\tfrac{1}{2}(A-A^{\top}))^{\vee}$ from the network’s matrix output $A\in\mathbb{R}^{3\times 3}$.
+I.4 Training and Evaluation Protocol
+Problem Formulation.
 The network is trained to predict the drone’s 3D position $\mathbf{p}_{t}\in\mathbb{R}^{3}$ at the end of a given time window, based on a sequence of noisy velocity measurements and their corresponding covariances within that window (e.g., a 1-second history).
@@
 We report the following pose-regression metrics over the test set:
-  • Absolute Trajectory Error (ATE): The root-mean-square error between the ground-truth and predicted 3D positions over the entire trajectory, measured in meters.
-  • ATE%: The ATE normalized by the total trajectory length and expressed as a percentage ($100\times\text{ATE}/\text{length}$).
+  • Absolute Trajectory Error (ATE): The root-mean-square error (RMSE) between the ground-truth positions $p_{k}$ and predicted 3D positions $\hat{p}_{k}$ over the entire trajectory, calculated as $\sqrt{\frac{1}{N}\sum_{k=1}^{N}\|\hat{p}_{k}-p_{k}\|^{2}}$, measured in meters.
+  • ATE%: The ATE normalized by the total trajectory length $L$, expressed as a percentage ($100\times\text{ATE}/L$).
 This metric provides a scale-invariant measure of error, which is crucial for fairly comparing performance across our aggressive flight trajectories of varying lengths.
-  • Relative Pose Error (RPE): The error measured over fixed-length sub-trajectories, capturing local drift.
+  • Relative Translation Error (RTE): This metric evaluates local consistency by measuring the relative translational error over fixed-length sub-trajectories.
+For each window of duration $\Delta t=2.0\,\text{s}$ ($n=400$ samples), we compute the RMSE of the positional difference between estimated and ground-truth relative displacements: $\|(\hat{p}_{k+n}-\hat{p}_{k})-(p_{k+n}-p_{k})\|$.
+We report the mean error by sliding this window across the trajectory with a stride of $0.5\,\text{s}$ ($100$ samples) to ensure a robust assessment of short-term stability.
 To explicitly validate equivariance, we also evaluate all models on the test set after applying a set of random $\mathrm{SO}(3)$ rotations to the entire input sequence.
-G.5 Eigenvector Canonicalization for the VN Baseline
-To resolve ambiguities in the eigendecomposition $C=V\Lambda V^{\top}$ for the VN baseline, we canonicalize the eigenvector matrix $V=[\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}]$ as follows:
+I.5 Eigenvector Canonicalization for the VN Baseline
+To resolve ambiguities in the eigendecomposition $C=V\Lambda V^{\top}$ for the VN baseline while preserving $SO(3)$ geometry, we canonicalize the eigenvector matrix $V=[\mathbf{e}_{1},\mathbf{e}_{2},\mathbf{e}_{3}]$ using the following sequence.
+To avoid producing reflection matrices (where $\det V=-1$), sign disambiguation must precede the enforcement of a right-handed frame.
   1. 1.
-    Right-handed Frame: If $\det V<0$, we set $\mathbf{e}_{3}\leftarrow-\mathbf{e}_{3}$ to ensure $\det V=+1$.
+    Sign Disambiguation: For each eigenvector $\mathbf{e}_{i}$ associated with a distinct eigenvalue, we enforce a consistent orientation relative to the velocity vector $\mathbf{v}$ by ensuring $\mathbf{v}^{\top}\mathbf{e}_{i}\geq 0$.
+If $\mathbf{v}^{\top}\mathbf{e}_{i}<0$, we set $\mathbf{e}_{i}\leftarrow-\mathbf{e}_{i}$.
   2. 2.
-    Sign Disambiguation: For each eigenvector $\mathbf{e}_{i}$ with a distinct eigenvalue, we enforce a consistent sign by ensuring $\mathbf{v}^{\top}\mathbf{e}_{i}\geq 0$.
-If not, we set $\mathbf{e}_{i}\leftarrow-\mathbf{e}_{i}$.
+    Right-handed Frame: After the individual signs are fixed, we ensure the matrix $V$ represents a proper rotation.
+If $\det V<0$, we flip the sign of the third eigenvector, $\mathbf{e}_{3}\leftarrow-\mathbf{e}_{3}$, to ensure $\det V=+1$.
+This step is performed last to ensure that any sign changes in the previous step do not inadvertently result in a reflection frame.
   3. 3.
-    Multiplicity Handling: In the rare case of repeated eigenvalues, we use the projection of the velocity vector $\mathbf{v}$ onto the corresponding eigenspace to deterministically define the first basis vector, then complete the basis via Gram-Schmidt.
-All steps use only equivariant operations, preserving the overall symmetry of the baseline.
-Appendix H Proof of $\mathrm{SO}(3)$-Equivariance for ReLN Velocity Extract with Covariance Inputs
+    Multiplicity Handling: In cases of repeated eigenvalues, we project the velocity vector $\mathbf{v}$ onto the corresponding eigenspace to uniquely define the first basis vector of that subspace, then complete the orthonormal basis via the Gram-Schmidt process, followed by the determinant check in Step 2.
+By applying the determinant constraint after sign alignment, we guarantee that the baseline operates on valid rotation matrices, ensuring a fair and geometrically consistent comparison with our $SO(3)$-equivariant model.
+I.6 Full Experimental Results and Ablation Study for Drone State Estimation
+Table 13 provides the exhaustive results for all models across three input modalities: $v,(v,C),(v,\log C)$.
+This comprehensive comparison highlights that while the $\log$-covariance interface generally improves performance for most equivariant models, our ReLN architecture remains the superior backbone due to its ability to jointly model the full reductive structure.
+Table 13: Complete performance report on the drone dataset including all input variants and ablations.
+We report ATE (meters), ATE% (scaled), and RTE (meters).
+SO(3) denotes evaluation under random test-time rotations.
+Model Input ID (In-Distribution) SO(3) (Rotated)
+ATE $\downarrow$ ATE% $\downarrow$ RTE $\downarrow$ ATE $\downarrow$ ATE% $\downarrow$ RTE $\downarrow$
+ResNet $v$ 208.07 95.06 107.60 217.02 100.39 111.29
+ResNet $(v,C)$ 205.11 94.94 106.07 213.26 98.90 109.37
+Vector Neurons $v$ 17.36 7.52 13.51 17.36 7.52 13.51
+Vector Neurons $(v,C)$ 191.78 88.66 98.39 190.22 88.47 98.26
+Tensor Field Network $v$ 24.59 10.95 18.23 24.59 10.95 18.23
+Tensor Field Network $(v,C)$ 17.56 7.60 14.40 17.56 7.60 14.40
+Tensor Field Network $(v,\log C)$ 16.83 7.56 13.34 16.83 7.56 13.34
+SE(3)-Transformer $v$ 22.22 9.85 17.63 22.22 9.85 17.63
+SE(3)-Transformer $(v,C)$ 21.67 9.37 16.77 21.67 9.37 16.77
+SE(3)-Transformer $(v,\log C)$ 20.12 8.84 15.36 20.12 8.84 15.36
+Ablations (Ours)
+Lie Neurons $(v,C)$ 16.86 7.43 13.65 16.86 7.43 13.65
+Lie Neurons $(v,\log C)$ 15.65 6.76 12.04 15.65 6.76 12.04
+ReLN (no semisimple) $(v,\log C)$ 16.27 7.00 12.65 16.27 7.00 12.65
+Our Equivariant Model
+ReLN $v$ 16.85 7.31 12.70 16.85 7.31 12.70
+ReLN $(v,C)$ 16.49 7.21 13.02 16.49 7.21 13.02
+ReLN $(v,\log C)$ 13.92 5.99 11.04 13.92 5.99 11.04
+The Lie Neurons architecture (Lin et al., 2024a) is mathematically equivalent to the “no-center” variant of ReLN, i.e., ReLN with the bilinear form restricted to the semisimple ideal.
+Figure 7: Qualitative comparison of reconstructed trajectories ReLN variants maintain superior tracking fidelity and stability compared to non-equivariant and standard equivariant baselines.
+Qualitative Analysis of Trajectory Fidelity.
+Figure 7 validates the qualitative robustness of ReLNs under high-dynamic flight conditions.
+While standard MLP-based ResNets and eigendecomposition-based Vector Neurons suffer from rapid error accumulation and structural divergence, ReLN maintains high fidelity to the ground truth even in velocity-only configurations.
+The integration of covariance, $(v,C)$, and specifically the $\log$-covariance provides effective regularization via uncertainty-aware weighting.
+Notably, while steerable baselines such as TFN and SE(3)-Transformers show improved stability with $\log$-covariance, they still exhibit observable deviations during sharp maneuvers where geometric consistency is critical.
+This confirms our finding that the reductive Lie-algebraic backbone offers a more stable and geometrically principled space for fusing heterogeneous sensor signals and their associated uncertainties.
+Appendix J Proof of $\mathrm{SO}(3)$-Equivariance for ReLN Velocity Extract with Covariance Inputs
 This section provides a formal proof for the $\mathrm{SO}(3)$-equivariance of our Reductive Lie Neuron (ReLN) architecture when processing a velocity vector and a covariance matrix.
 We first establish the foundations for processing covariance matrices within a Lie-algebraic framework and then present the main proof.
-H.1 $\mathrm{SO}(3)$-Equivariant Vector Extraction via Skew-Symmetric Projection
+J.1 $\mathrm{SO}(3)$-Equivariant Vector Extraction via Skew-Symmetric Projection
 Our network, $\Phi$, is designed to be adjoint-equivariant.
-It maps geometric inputs—such as a embedded velocity $K\in\mathfrak{so}(3)$ and a covariance matrix $S\in\mathrm{SPD}(3)$—to a matrix feature $A\in\mathbb{R}^{3\times 3}$.
+It maps geometric inputs—such as an embedded velocity $K\in\mathfrak{so}(3)$ and a covariance matrix $S\in\mathrm{SPD}(3)$—to a matrix feature $A\in\mathbb{R}^{3\times 3}$.
 The inputs transform under the adjoint action of any rotation $R\in\mathrm{SO}(3)$:
@@
 The following proposition formalizes the equivariance of this extraction mechanism.
-Proposition H.1 (Equivariance of Skew-Symmetric Extraction).
-Let a network $\Phi$ and its inputs transform according to Eqs. 47 and 48. If a vector $\hat{\mathbf{v}}\in\mathbb{R}^{3}$ is extracted from the output matrix $A=\Phi(K,S)$ via the projection
-$A_{\mathrm{skew}}=\tfrac{1}{2}(A-A^{\top}),\qquad\hat{\mathbf{v}}=(A_{\mathrm{skew}})^{\vee},$ (49)
-then the vector $\hat{\mathbf{v}}^{\prime}$ extracted from the transformed output $A^{\prime}=\Phi(K^{\prime},S^{\prime})$ transforms covariantly as $\hat{\mathbf{v}}^{\prime}=R\hat{\mathbf{v}}$.
+Proposition J.1 (Equivariance of Skew-Symmetric Extraction).
+Let a network $\Phi$ and its inputs transform according to Eqs. 58 and 59. If a vector $\tilde{\mathbf{v}}\in\mathbb{R}^{3}$ is extracted from the output matrix $A=\Phi(K,S)$ via the projection
+$A_{\mathrm{skew}}=\tfrac{1}{2}(A-A^{\top}),\qquad\tilde{\mathbf{v}}=(A_{\mathrm{skew}})^{\vee},$ (60)
+then the vector $\tilde{\mathbf{v}}^{\prime}$ extracted from the transformed output $A^{\prime}=\Phi(K^{\prime},S^{\prime})$ transforms covariantly as $\tilde{\mathbf{v}}^{\prime}=R\tilde{\mathbf{v}}$.
 Proof.
 By the adjoint-equivariance property in Eq.
-48, the network satisfies $\Phi(RKR^{\top},RCR^{\top})=R\,\Phi(K,C)\,R^{\top}=RAR^{\top}$.
+59, the network satisfies $\Phi(RKR^{\top},RSR^{\top})=R\,\Phi(K,S)\,R^{\top}=RAR^{\top}$.
 Let $A^{\prime}=RAR^{\top}$.
@@
 Applying this property yields the desired result:
-$\hat{\mathbf{v}}^{\prime}=(A^{\prime}_{\mathrm{skew}})^{\vee}=(\mathrm{Ad}_{R}(A_{\mathrm{skew}}))^{\vee}=R\,(A_{\mathrm{skew}})^{\vee}=R\hat{\mathbf{v}}$. (51)
-Remark 4.
+$\tilde{\mathbf{v}}^{\prime}=(A^{\prime}_{\mathrm{skew}})^{\vee}=(\mathrm{Ad}_{R}(A_{\mathrm{skew}}))^{\vee}=R\,(A_{\mathrm{skew}})^{\vee}=R\tilde{\mathbf{v}}$. (62)
+Remark J.2.
 The proof relies on three properties: (i) both inputs transform under the adjoint action $X\mapsto RXR^{\top}$; (ii) the network $\Phi$ is equivariant to this action; and (iii) the output is projected onto $\mathfrak{so}(3)$ before the vee operator is applied.
 As established previously, these conditions hold whether the network ingests the raw covariance $S$ or its logarithm $\log S$.
-Large Language Model (LLM) Usage
-We used a large language model (LLM) to aid in polishing the writing and improving grammatical clarity of the manuscript.
-The LLM did not contribute to the research ideation, experiments, or technical content; all scientific claims and results were generated solely by the authors.
+Appendix K Details on 3D Gaussian Splatting Experiments
+In this section, we detail the experimental protocols and architectural modifications for the 3D Gaussian Splatting (3DGS) experiments.
+Figure 5 provides a comprehensive overview of the ReLN-integrated Gaussian-MAE architecture, highlighting the equivariant data flow from geometric lifting to projection.
+We provide a rigorous theoretical justification for our architectural design to achieve end-to-end equivariance by jointly learning geometric and photometric features within a unified Lie-algebraic framework.
+K.1 Datasets and Protocol
+We adopt a standard two-stage protocol common in self-supervised learning for 3D representation (Ma et al., 2025), separating the pre-training domain from the downstream evaluation task to assess generalization.
+Pre-training: ShapeSplat (ShapeNet-derived).
+We utilize ShapeSplat (Ma et al., 2025), a large-scale collection of object-level 3D Gaussian primitives derived from ShapeNet.
+Each object is represented by a set of anisotropic Gaussians $\Theta_{i}=\{\mu_{i},\Sigma_{i},\alpha_{i},c_{i}\}_{i=1}^{N}$.
+Following the baseline protocol, we sample $N=1,024$ Gaussians via furthest point sampling (FPS) and apply a $60\%$ masking ratio for the masked autoencoding objective.
+Downstream: ModelNet10 Classification.
+To evaluate rotational robustness, we fine-tune the pre-trained encoder on ModelNet10.
+We report accuracy on: (i) Aligned: The canonical test set.
+(ii) Rotated: The test set under random rotations $R\sim SO(3)$, testing whether the learned representations remain stable under arbitrary orientations.
+K.2 Holistic Equivariance via Joint Feature Learning
+A critical limitation of the baseline architecture (Ma et al., 2025) is the naive concatenation of heterogeneous Gaussian attributes.
+A 3D Gaussian combines equivariant quantities ($\mu,\Sigma$) with invariant quantities ($c,\alpha$).
+Our goal in ReLN is to construct a holistic equivariant architecture where these distinct attribute types are learned jointly without violating the symmetry of the feature space.
+1. The Equivariant Component (Active Geometry).
+The mean position $\mu\in\mathbb{R}^{3}$ and covariance $\Sigma\in\mathbb{R}^{3\times 3}$ dictate the spatial structure.
+Under a global rotation $R\in SO(3)$, these transform via $\mu\mapsto R\mu$ and $\Sigma\mapsto R\Sigma R^{\top}$.
+Recognizing that the covariance update corresponds to the adjoint action $\mathrm{Ad}_{R}(\Sigma)=R\Sigma R^{-1}$, we embed $\Sigma$ into the symmetric subspace of $\mathfrak{gl}(3)$.
+This ensures that the primary geometric features are processed within a consistent $\mathrm{Ad}$-equivariant vector space throughout the network.
+2. The Invariant Component (Photometry and Geometric Norms).
+To support robust representation learning, we augment the state space with a set of $O(3)$-invariant scalars (Type-0 features). This set includes:
+[…]
+  • Derived Geometric Invariants: Crucially, we also explicitly compute rotation-invariant geometric descriptors, such as the norms of position vectors ($\|\mu\|$) and rotation/scale magnitudes derived from $\Sigma$.
+In the ReLN framework, treating these as Type-0 features serves to facilitate its feature integration.
+These scalars participate in the learning process by modulating the equivariant features without violating the underlying group structure and equivariance.
+3. Joint Learning in ReLN Attention.
+The explicit stratification of these streams is a structural prerequisite for joint equivariant learning.
+By assigning each attribute to its correct algebraic type (Type-1/Type-2 for geometry, Type-0 for scalars), the ReLN attention mechanism can validly integrate information across all modalities.
+The invariant scalars ($c,\alpha,\|\mu\|$) contribute to the computation of attention scores (invariant inner products), thereby modulating how the equivariant geometric features ($\mu,\Sigma$) are aggregated.
+This design ensures that the entire network optimizes a single, cohesive objective function while guaranteeing that the final representation remains equivariant end-to-end.
+K.3 Architectural Modifications
+To enforce this holistic equivariance, we replaced non-robust components of the baseline architecture with ReLN equivalents that respect the defined type system.
+ReLNEncoder and Geometric Aggregation.
+Our ReLNEncoder begins with a Geometric Lifting stage that maps raw Gaussian parameters ($C,S,R$) into $\mathfrak{gl}(3)$.
+The baseline uses a ‘SoftEncoder‘ (MLP on raw coordinates) which breaks equivariance because standard linear weights $W$ do not rotate with the input frame.
+We replaced this with ReLNEncoder, which employs $\mathrm{Ad}$-equivariant linear layers (ReLNLinear) to process local neighborhoods.
+ReLNEncoder incorporates a Geometric Aggregator, which utilizes an equivariant norm-based max pooling mechanism to compress grouped splats into structured $C^{\prime}\times 9$ tokens.
+This ensures that the fundamental geometric tokens are processed within a mathematically consistent vector space.
+Central Positional Embedding.
+To provide spatial context without violating symmetry, we process group centroids ($\mu$) through a dedicated Central Positional Embedding.
+This module lifts 3D spatial coordinates into a structured Lie-algebraic representation, which is then injected into both the encoder and decoder stages (as illustrated in Figure 5).
+This injection strategy ensures that the transformer blocks distinguish relative spatial relations while maintaining rotational equivariance across the entire processing pipeline.
+Invariant Global Pooling (vs. CLS Token).
+For downstream classification tasks, we eschew the conventional learnable [CLS] token, recognizing that a fixed-reference-frame parameter inherently imposes a canonical orientation bias that violates rotational symmetry.
+Instead, we employ a permutation-invariant Global Max Pooling over the final sequence of equivariant tokens.
+This ensures that the aggregated global representation remains stable and rotates consistently with the input geometry, thereby satisfying holistic rotational symmetry for robust 3D representation learning.
+K.4 Training and Evaluation Protocol
+Pre-training Setup.
+We pre-trained both the baseline and ReLN models for 300 epochs on the ShapeSplat (ShapeNet) dataset.
+We utilized the AdamW optimizer with an initial learning rate of $1\times 10^{-3}$, a weight decay of $0.05$, and a cosine annealing scheduler with a 10-epoch warm-up.
+For the ReLN model, we explicitly disabled the soft_knn option to prevent the leakage of non-equivariant features from the legacy encoder implementation.
+Fine-tuning Setup.
+For the downstream ModelNet10 classification task, we fine-tuned the pre-trained encoder for an additional 300 epochs.
+In this phase, we adjusted the learning rate to $5\times 10^{-4}$ while maintaining the same weight decay ($0.05$) and cosine scheduling strategy.
+Evaluation Metrics. To rigorously evaluate geometric stability, we construct two test benchmarks:
+  • Standard (Aligned): The original ModelNet10 test set where objects are canonically aligned.
+  • Rotated (SO(3)): We apply a random rotation $R\in\mathrm{SO}(3)$ to every object. For a Gaussian splat, we transform parameters as $\mu^{\prime}=R\mu$, $\Sigma^{\prime}=R\Sigma R^{\top}$.
+K.5 Additional Convergence Results
+Figure 8 presents the detailed training dynamics.
+The ReLN model consistently achieves lower loss values across all geometric attributes—Rotation, Scale, and Spherical Harmonics (SH)—demonstrating that enforcing equivariance stabilizes the learning of intrinsic 3D shape descriptors.
+Figure 8: Pre-training Dynamics on 3D Gaussian Splats.
+Comparison of training convergence between the Baseline (Ma et al., 2025) and ReLN (Ours) across 300 epochs on ShapeNet.
+The red shaded regions highlight the gap where ReLN achieves lower reconstruction error than the baseline.
+ReLN consistently outperforms the baseline across all tracked metrics (including Rotation, Scale, Density, and Chamfer Distance), demonstrating that equivariance improves every aspect of 3D Gaussian reconstruction.
+Appendix L Experimental Details for the Double-Pendulum Benchmark
+To ensure a rigorous and fair comparison, we adhere to the experimental protocol and system parameters of the EMLP benchmark (Finzi et al., 2021).
+System dynamics and dataset generation.
+We learn Hamiltonian dynamics of a 3D double spring pendulum exhibiting $O(2)$ symmetry about the $z$-axis.
+All physical constants are set to unity ($m_{i}=k_{i}=l_{i}=1.0$) with gravity $g=[0,0,1]$.
+Ground-truth trajectories are generated by Hamiltonian integration over a horizon of $T=30$s.
+We construct $1500$ trajectory chunks of duration $1$s (sampled at five $0.2$s intervals) and split them into $500/500/500$ for training/validation/testing.
+Models and training.
+[…]
+EMLP-HNN baseline (equivariant layer construction).
+Let $G=O(2)$ be the symmetry group acting on the $(x,y)$ plane (with $z$ unchanged), e.g., $R(\phi)=\mathrm{diag}(R_{2\times 2}(\phi),1)$ (and similarly for reflections), and let $\rho(\cdot)$ denote the induced representation on the state space.
+Equivariance of a linear map $W:V_{\mathrm{in}}\!\to\!V_{\mathrm{out}}$ requires
+$\rho_{\mathrm{out}}(g)\,W\;=\;W\,\rho_{\mathrm{in}}(g),\qquad\forall g\in G.$ (63)
+Following (Finzi et al., 2021), EMLP enforces this by constraining the vectorized weights $v=\mathrm{vec}(W)$ under the product representation $\rho=\rho_{\mathrm{out}}\otimes\rho_{\mathrm{in}}^{*}$:
+$\rho(g)\,v\;=\;v,\qquad\forall g\in G.$ (64)
+To avoid infinitely many constraints, the condition is reduced to a finite generator set consisting of Lie algebra generators $\{A_{i}\}$ and discrete generators $\{h_{j}\}$:
+$d\rho(A_{i})\,v=0\quad\text{and}\quad(\rho(h_{j})-I)\,v=0,$ (65)
+which are assembled into a linear system $Cv=0$.
+An equivariant basis $Q$ is obtained by computing a nullspace of $C$ (e.g., via SVD/Krylov methods as in the reference implementation), and the layer is parameterized as $v=Q\beta$ during training (no further constraint solving during SGD).
+This basis computation is group- and representation-specific (e.g., $O(2)$ vs.
+$SO(2)$).
+EMLP nonlinearity.
+We use the standard EMLP gated nonlinearity (Finzi et al., 2021), where scalar channels produce gates that multiplicatively modulate non-scalar feature blocks, preserving equivariance while enabling expressive nonlinear mixing.
+ReLN-HNN (ours).
+In contrast, ReLN-HNN bypasses numerical constraint solving and constructs equivariant/invariant primitives in closed form within a reductive Lie-algebraic space (here $\mathfrak{gl}(3)$), using bilinear/Killing-form-based operations to obtain symmetry-consistent features without group-specific basis recomputation.
+The same architecture is used across different symmetry settings (e.g., $O(2)$, $SO(2)$, $D_{6}$) without re-solving linear constraints.
+Computational efficiency measurement.
+Per-step FLOPs in Table 6 are measured using JAX HLO cost analysis, which counts device-executed floating-point operations including internal tensor contractions and projections.